### v4

# On the Fontaine-Mazur conjecture for $p=3$

Abstract.
In this article, we prove the remaining open cases of the Fontaine-Mazur conjecture on two-dimensional regular Galois representations over $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ when $p=3$, hence concluding the conjecture in the regular case for all odd primes. Our result is a sequel to Pan’s work, based on some recent progress on $p$-adic Langlands correspondence, Galois deformation theory and a potential pro-modularity result.
1. Introduction
In a recent marvelous work [Pan22], Pan has proved the following theorem:
Theorem 1.0.1.
[Pan22, Theorem 1.0.4] Let $p$ be an odd prime number and
[…]
be a continuous, irreducible representation such that
  • $\rho$ is only ramified at finitely many places,
  • $\rho|_{G_{\mathbb{Q}_{p}}}$ is de Rham of distinct Hodge-Tate weights,
  • $\rho$ is odd, i.e. $\det\rho(c)=-1$ for any complex conjugation $c\in\operatorname{Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$.
  • If $p=3$, the semi-simplification of $\bar{\rho}|_{G_{\mathbb{Q}_{p}}}$ is not of the form $\eta\oplus\eta\omega$ for some character $\eta$, where $\omega$ denotes the $\mod p$ cyclotomic character.
Then $\rho$ arises from a cuspidal eigenform up to twist.
In this paper, we will remove the last assumption. More precisely, combining the work [SW01], [HT15], [Kis09], [Pan22] and [KW09a], we have
Theorem 1.0.2 (Theorem 5.1.1, Theorem 6.1.1).
Let $p$ be an odd prime number and
[…]
be a continuous, irreducible representation such that
  • $\rho$ is only ramified at finitely many places,
  • $\rho|_{G_{\mathbb{Q}_{p}}}$ is de Rham of distinct Hodge-Tate weights,
  • $\rho$ is odd,
Then $\rho$ arises from a cuspidal eigenform up to twist.
The main difficulty of the Fontaine-Mazur conjecture at $p=3$ is the residually reducible case. To outline the proof of Theorem 1.0.2 in this case, we suppose the semisimplification of $\bar{\rho}$ is (up to twist) isomorphic to $1\oplus\bar{\chi}$ for some continuous odd almost unramified character $\bar{\chi}:G_{\mathbb{Q}}\to\mathbb{F}^{\times}$, where $\mathbb{F}$ is a finite field of characteristic $p$.
Our strategy is a refinement of Pan’s. Precisely, we prove a potential pro-modularity argument, rather than just identify some irreducible component of the universal deformation ring as the in the big Hecke algebra, so that our strategy works for all odd primes.
To illustrate the obstructions in the case $p=3$ and $\bar{\chi}|_{G_{\mathbb{Q}_{p}}}=\omega^{-1}$ (we call it “exceptional” afterwards) in Pan’s method, we first briefly recall the main steps of Theorem 1.0.1 in the residually reducible case.
1) Local-global compatibilty results. There are two main purposes in this part. The first one is to justify whether the Galois representation defined by a map from the big Hecke algebra to $\overline{\mathbb{Q}_{p}}$ comes from a cuspidal automorphic representation. The second one is to collect some necessary results for the patching argument (see [Pan22, page 1033] for details). Most results in this part are applications of Paškūnas theory [Paš13], which excludes our exceptional case.
2) Patching at nice primes. In this part, Pan proves an $R_{\mathfrak{q}}=\mathbb{T}_{\mathfrak{q}}$ theorem following the strategy of Skinner-Wiles [SW99], where $\mathfrak{q}$ is a one-dimensional prime (called “nice”) of the big Hecke algebra satisfying some special properties. In other words, we can identify some irreducible components of the universal pseudo-deformation ring as in the big Hecke algebra, if those contain a “nice” prime.
3) Ordinary Fontaine-Mazur conjecture in the residually reducible case. In this part, Pan generalizes the main theorem of [SW99] and proves a finiteness result between the Iwasawa algebra and the universal ordinary pseudo-deformation ring. This allows us to find enough pro-modular primes, i.e. primes coming from the big Hecke algebra, in the ordinary locus of the universal pseudo-deformation ring.
4) Galois deformation theory for pseudo-deformation rings. From the previous three steps, we only need to show that there exists an irreducible component of $\operatorname{Spec}R$ containing both a “nice” prime and the one defined by our Galois representation $\rho$. To produce enough “nice” primes in the ordinary locus of $\operatorname{Spec}R$ (write as $\operatorname{Spec}R^{\operatorname{ord}}$), we first need to know the kernel from the local pseudo-deformation at $p$ to its ordinary locus (write the map as $R_{p}\to R_{p}^{\operatorname{ord}}$), and obtain an exact bound of the Krull dimension of $\operatorname{Spec}R^{\operatorname{ord}}$. In this step, Pan only finds “potentially nice” primes, which is enough for his purpose.
Here we point out some differences and difficulties in the exceptional case and describe the improvements in this paper:
In step 1), as we mentioned before, some results in the exceptional case are missing in [Paš13]. The reason is that the behavior of the block in this case is specialzed and more complicated (see the list in A.2). Recently, this case has been studied in [PT21] (even for $p=2$). Based on Paškūnas theory for small primes, Tung studied the Fontaine-Mazur conjecture for $p=2$ in [Tun21b] and $p=3$ in [Tun21a] under Talylor-Wiles hypothesis. In this article, we also use these results to obtain the same local-global compatibility arguments (see Section 3.3) and further prove some $R=\mathbb{T}$ theorems (see Theorem 4.1.4 and Theorem 4.3.3) without Pan’s assumption.
In step 2), some crucial properties of the local framed deformation ring at $p$, such as its normality, for proving the patching argument were missing at the exceptional case. Combining recent progress [BIP23] on local framed deformation ring at $p$, we are able to fill in such gaps in Pan’s strategy.
In step 4), there are several obstructions, and here for instance, we explain the most subtle case, $p=3$ and $\bar{\chi}=\omega=\omega^{-1}$.
  • When $p\geq 5$ and $\bar{\chi}|_{G_{\mathbb{Q}_{p}}}\neq\mathbf{1},\omega^{-1}$ (after a twist, we may assume this as $\omega$ is not quadratic), the kernel of $R_{p}\to R_{p}^{\operatorname{ord}}$ has been studied in [Paš13, Appendix B]. Note that in this case, we have $\dim_{\mathbb{F}}H^{1}(G_{\mathbb{Q}_{p}},\mathbb{F}(\bar{\chi}^{-1}))=1$. Then by [Kis09, Corollary 1.4.4], there is a natural surjection (hence isomorphism) from the universal pseudo-deformation ring to the universal (unframed) deformation ring. Thus, we can calculate the pseudo-deformation ring explicitly. See [Pan22, Lemma 7.4.4 & 7.4.8]. However, such a calculation is unknown when $p=3$ and $\bar{\chi}=\omega^{-1}$ as $\dim_{\mathbb{F}}H^{1}(G_{\mathbb{Q}_{p}},\mathbb{F}(\bar{\chi}^{-1}))=2$. Despite the difficulties, combining some geometric properties of the local pseudo-deformation ring studied in the recent work [BIP23] and [BJ23], we are able to develop analogous results to tackle this issue without knowing the structure of the local pseudo-deformation ring.
  • When $p\geq 5$ and $\bar{\chi}\neq\omega^{-1}$ (again, we may assume this as $\omega$ is not quadratic), then we have $H^{0}(G_{\mathbb{Q}},\operatorname{ad}^{0}(\bar{\rho}_{b}(1)))=0$. Then by the Greenberg-Wiles formula, the universal unframed ordinary deformation ring $R_{b}^{\mathrm{ord}}$ is of dimension at least $2$, and hence of the same dimension as the Iwasawa algebra. By Pan’s finiteness result (in step 3)), we can find enough pro-modular primes in $\operatorname{Spec}R^{\operatorname{ord}}$, and also “potentially nice” primes. In our case, the dimension of $H^{0}$ is $1$, leading the lower bound of $R_{b}^{\mathrm{ord}}$ to be $1$. Hence, the arguments above cannot work. To overcome this difficulty, we prove a generalized Greenberg-Wiles formula (Proposition 2.4.3) to study the dimension of some irreducible component of $R_{b}^{\mathrm{ord}}$.
  • In the definition of a “nice” prime (see Definition 4.1.3), we need to know the corresponding representation is not dihedral, i.e. induced from a quadratic character. When $p\geq 5$ and $\bar{\chi}=\omega$, this condition is trivially satisfied, whereas in our case, we must not ignore it. Note that by Leopoldt’s conjecture for the abelian CM field $F(\mu_{3})$, the upper bound of the dihedral locus of $\operatorname{Spec}R$ is $[F:\mathbb{Q}]+2$, where $F$ is an appropriate abelian base change of $\mathbb{Q}$. Thus, to find (potentially) “nice” primes, we first need to find a pro-modular prime with larger dimension to avoid the dihedral locus. However, in the non-generic reducible case (see Section 5.4), the lower bound of the dimenson of the pro-modular prime we can find there is $[F:\mathbb{Q}]+2$, which is not large enough to find a (potentially) “nice” prime when we are in the exceptional case. Therefore, we need a new strategy of this part here. See Remark 5.4.8.
    In this paper, we use the method developed in the author’s previous work [Zha24] so that we only need to find an irreducible one-dimensional pro-modular prime instead of a “nice” prime in each irreducible component of $\operatorname{Spec}R^{\mathrm{ps}}$ of large dimension, which is much more convenient to find out. Furthermore, it allows us to prove a potential big $R=\mathbb{T}$ theorem in the residually reducible case in fact. See Remark 5.1.3.
As we mentioned in the previous discussions, we actually prove the following potential pro-modularity result in the residually reducible case, which is the main innovation in this paper. To state it more clearly, we give the following settings.
Let $p$ be an odd prime and let $F$ be an abelian totally real field in which $p$ splits completely. Write $\Sigma_{p}$ as the set of places of $F$ above $p$ and let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$. Let $\bar{\chi}:G_{F,\Sigma}\to\mathbb{F}^{\times}$ be a continuous, odd, almost unramified character which can be extended to a character of $G_{\mathbb{Q}}$. Then $T=1+\bar{\chi}$ defines a pseudo-representation. Further assume that for any $v|p$, $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1}$. Let $\chi$ be a de Rham lifting of $\bar{\chi}$.
Theorem 1.0.3 (Theorem 5.5.1, Proposition 5.5.2).
There exists an abelian totally real extension $F_{1}/F$ of even degree such that $p$ splits completely in $F_{1}$ and for any irreducible component of the universal pseudo-deformation ring $R_{F_{1}}^{\mathrm{ps}}$ (parametrizing all pseudo-deformations of $T|_{G_{F_{1}}}$ unramified out places lying above $\Sigma$ with fixed determinant $\chi$) of dimension at least $1+2[F_{1}:\mathbb{Q}]$, it is pro-modular of dimension $1+2[F_{1}:\mathbb{Q}]$. The corresponding big Hecke algebra over $F_{1}$ is equidimensional of dimension $1+2[F_{1}:\mathbb{Q}]$.
Remark 1.0.4.
1) Actually, by Proposition 2.5.2, we know that for any irreducible component of $R_{F_{1}}^{\mathrm{ps}}$, it is either of dimension at least $1+2[F_{1}:\mathbb{Q}]$ or contained in its reducible locus, hence of dimension at most $2$ by Leopoldt’s conjecture. Thus, we may view this result as a potential big $R=\mathbb{T}$ theorem, identifying $\mathbb{T}$ as the part of $R$ with large dimension.
2) In a future work, we will discuss more about applications of this potential pro-modularity result.
This paper is organized as follows. In Section 2, we recall some classical results in Galois deformation theory and prepare some results used for the proof of the main theorem. We refer readers to Böckle’s paper in [BBD+13] for classical theory and [BIP23] and [BJ23] for the recent progress. In Section 3, we give the local-global compatibility results without Pan’s assumption. Most of the proof in this section is the same as in [Pan22, Section 3]. In Section 4, we give some $R=\mathbb{T}$ theorems in both residually reducible and irreducible cases. Also, we recall Pan’s finiteness result in the ordinary case, which is devoted to finding pro-modular primes. In Section 5, we prove the main theorem in the residually reducible case, following the strategy in [Pan22, Section 7.4]. In Section 6, we prove the main theorem in the residually irreducible case. For convenience of readers, in Appendix, we present most of the results of Paškūnas theory in [Paš13] and [PT21] used in Section 3.
1.1. Notations
[…]
We use $\mathfrak{U}_{\mathcal{O}}$ to denote the category of Artinian local $\mathcal{O}$-algebras with residue field $\mathbb{F}$. For an $\mathcal{O}$-module $M$, we use $M^{\vee}$ to denote its Pontryagin dual $\operatorname{Hom}_{\mathcal{O}}^{\mathrm{cont}}(M,K/\mathcal{O})$, and we use $M^{d}$ to denote its Schikhof dual $\operatorname{Hom}_{\mathcal{O}}^{\mathrm{cont}}(M,\mathcal{O})$. We use $M_{\operatorname{tf}}$ to denote its maximal $\mathcal{O}$-torsion free quotient.
For a prime ideal $\mathfrak{p}$ of a commutative ring $R$, we denote its residue field by $k(\mathfrak{p})$. Let $R_{\mathfrak{p}}$ be the localization at $\mathfrak{p}$. We write $\widehat{R_{\mathfrak{p}}}$ as its $\mathfrak{p}$-adic completion. We say a commutative ring $R$ is a CNL ring if it is a complete noetherian local ring, and we usually denote its maximal ideal by $\mathfrak{m}_{R}$.
Suppose $F$ is a number field with maximal order $O_{F}$. We write $F_{+}$ for the set of totally positive elements in $F$. For any finite place $v$, we write $F_{v}$ (resp. $O_{F_{v}}$) for the completion of $F$ (resp. $O_{F}$) at $v$, $\varpi_{v}$ for a uniformizer of $F_{v}$, $k(v)$ for the residue field, $\operatorname{Nm}(v)$ for the norm of $v$ (in $\mathbb{Q}$), $G_{F_{v}}$ for the decomposition group above $v$, $I_{F_{v}}$ (or just $I_{v}$ if there is no confusion for the number field $F$) for its inertia group and $\operatorname{Frob}_{v}$ for a geometric Frobenius element in $G_{k(v)}:=G_{F_{v}}/I_{F_{v}}$. If $l$ is a rational prime, then we denote $O_{F}\otimes_{\mathbb{Z}}\mathbb{Z}_{l}$ by $O_{F,l}$. The adele ring of $F$ will be denoted by $\mathbb{A}_{F}$ Suppose $\Sigma$ is a finite set of places of $F$. We use $G_{F,\Sigma}$ for the Galois group of the maximal extension of $F$ unramified outside $\Sigma$ and all infinite places. The absolute Galois group of $F$ is denoted by $G_{F}=\operatorname{Gal}(\mkern 1.5mu\overline{\mkern-1.5muF\mkern-1.5mu}\mkern 1.5mu/F)$.
We use $\varepsilon$ to denote the $p$-adic cyclotomic character and $\omega$ to denote the mod $p$ cyclotomic character. Our convention for the Hodge-Tate weight of $\varepsilon$ is $-1$. We use $\mu_{n}$ to denote a primitive $n$-th root of unity.
1.2. Acknowledgements
The author would like to thank his supervisor Professor Takeshi Saito for his constant help and encouragement. The author would also like to thank Kojiro Matsumoto for many helpful discussions.
2. Galois deformation theory
In this section, we study some properties about universal Galois deformation rings.
2.1. Local framed deformation rings
In this subsection, we collect some helpful results about the local deformation rings. We recommend a recent remarkable work [BIP23] to readers about this topic.
Let $p$ be a prime. Let $F_{v}$ be a finite extension of $\mathbb{Q}_{p}$. Let $K$ be an another finite extension of $\mathbb{Q}_{p}$ with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$.
Let $\chi:G_{F_{v}}\to\mathcal{O}^{\times}$ be a continuous character. Let $\bar{\rho}:G_{F_{v}}\to\operatorname{GL}_{2}(\mathbb{F})$ be a continuous representation with determinant $\bar{\chi}$. Denote by $D_{\bar{\rho}}^{\square,\chi}:\mathfrak{U}_{\mathcal{O}}\to\operatorname{Sets}$ the functor such that for $(A,\mathfrak{m}_{A})\in\mathfrak{U}_{\mathcal{O}}$, $D_{\bar{\rho}}^{\square,\chi}(A)$ is the set of continuous representations $\rho_{A}:G_{F_{v}}\to\operatorname{GL}_{2}(A)$ satisfying $\rho_{A}\mod\mathfrak{m}_{A}=\bar{\rho}$ and $\operatorname{det}\rho_{A}=\chi$. The functor $D_{\bar{\rho}}^{\square,\chi}$ of liftings (framed deformations) of $\bar{\rho}$ is pro-represented by a CNL $\mathcal{O}$-algebra $R_{v}^{\square}$.
Proposition 2.1.1.
1) The ring $R_{v}^{\square}$ is a local complete intersection, flat over $\mathcal{O}$ and of relative dimension $3([F_{v}:\mathbb{Q}_{p}]+1)$.
2) The ring $R_{v}^{\square}$ is a normal integral domain.
Proof.
The first result is [BIP23, Corollary 5.4]. The second one is [BIP23, Theorem 5.6]. ∎
For the rest of this subsection, we assume $p=3$, $F_{v}=\mathbb{Q}_{p}$ and $\bar{\rho}(\sigma)=\begin{pmatrix}\omega(\sigma)&*\\ 0&1\end{pmatrix}$, where * is possible to be zero. To decribe the framed deformation ring $R_{v}^{\square}$, we recall the main work in [Böc10].
Let $E=\mathbb{Q}_{3}(\mu_{3})$, and $L$ be the splitting field of $\bar{\rho}$. Then the group $U:=\operatorname{Gal}(L/E)$ is of order $1,3$ or $9$ (depending on *). If it is non-trivial, we denote by $u\in U$ a non-trivial element. Up to conjugation, we may assume that $\bar{\rho}(u)=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$.
Let $P_{E}$ be the pro-$3$ completion of $G_{E}$, which is actually a quotient of $G_{E}$ by a closed normal subgroup. We denote the fixed field of this group by $E(3)$. Define $U_{2}(\mathbb{F})$ as the upper triangular unipotent subgroup of $\operatorname{GL}_{2}(\mathbb{F})$. For any CNL $\mathcal{O}$-algebra $R$, define $\tilde{\Gamma}(R)$ as the inverse image of $U_{2}(\mathbb{F})$ via the natural surjection $\operatorname{GL}_{2}(R)\to\operatorname{GL}_{2}(\mathbb{F})$. Viewing $u$ as an element of $P_{E}$, we define the functor $\operatorname{EH}:\mathfrak{U}_{\mathcal{O}}\to\operatorname{Sets}$ of equivariant homomorphisms corresponding to $\bar{\rho}$ as follows:
$\operatorname{EH}:=\{\alpha\in\operatorname{Hom}_{\operatorname{Gal}(E/\mathbb{Q}_{3})}(P_{E},\tilde{\Gamma}(R)):\alpha\mod\mathfrak{m}_{R}=\bar{\rho}|_{G_{E}},\alpha(u)=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\}.$
Proposition 2.1.2.
The functor $\operatorname{EH}$ is representable. Let $(\tilde{R},\tilde{\alpha})$ denote the universal pair. Then $\tilde{R}$ is a versal hull of $\bar{\rho}$. If $\bar{\rho}$ is non-split, then $\tilde{R}$ is universal.
Proof.
For the first part, it follows from [Böc10, Proposition 2.2]. For the second part, see [Böc10, page 531]. ∎
Let $G:=\operatorname{Gal}(E/\mathbb{Q}_{3})=\{1,\sigma\}$. Then the versal deformation $\tilde{\rho}:G_{\mathbb{Q}_{3}}\to\operatorname{GL}_{2}(\tilde{R})$ factors through the natural surjection $G_{\mathbb{Q}_{3}}\twoheadrightarrow\operatorname{Gal}(E(3)/\mathbb{Q}_{3})=P_{E}\rtimes G$.
Let $F_{4}$ be the free pro-$3$ group with topological generators $x_{1},x_{2},x_{3},x_{4}$ with the following $G$-action:
$\sigma(x_{1})=x_{1}^{-1},\sigma(x_{2})=x_{2},\sigma(x_{3})=x_{3}^{-1},\sigma(x_{4})=x_{4}.$
Define $r_{0}:=x_{1}^{3}[x_{1},x_{2}][x_{3},x_{4}]$, where $[g,h]=ghg^{-1}h^{-1}$, and define
$r:=r_{0}\sigma(r_{0})^{-1}=x_{1}^{3}[x_{1},x_{2}][x_{3},x_{4}][x_{4},x_{3}^{-1}][x_{2},x_{1}^{-1}]x_{1}^{3}.$
Let $N_{4}\subset F_{4}$ be the closed normal subgroup generated by $r$. By [Böc10, Lemma 3.2], we know that the pro-$3$ group $F_{4}/N_{4}$ is as a group with $G$-action isomorphic to $P_{E}$.
Denote by $A_{i}\in\tilde{\Gamma}(R)\subset\operatorname{GL}_{2}(R)$ the image of $x_{i}$. Then we can deduce that
$A_{1}=\begin{pmatrix}\sqrt{1+bc}&b\\ c&\sqrt{1+bc}\end{pmatrix},A_{2}=\sqrt{1+a}\begin{pmatrix}\sqrt{1+d}&0\\ 0&\sqrt{1+d}^{-1}\end{pmatrix},$
$A_{3}=\begin{pmatrix}\sqrt{1+b^{\prime}c^{\prime}}&b^{\prime}\\ c^{\prime}&\sqrt{1+b^{\prime}c^{\prime}}\end{pmatrix},A_{4}=\sqrt{1+a^{\prime}}\begin{pmatrix}\sqrt{1+d^{\prime}}&0\\ 0&\sqrt{1+d^{\prime}}^{-1}\end{pmatrix},$
where $a,a^{\prime},c,c^{\prime},d,d^{\prime}\in\mathfrak{m}_{R}$ and $b,b^{\prime}\in R$ (whether $b$ or $b^{\prime}$ lie in $\mathfrak{m}_{R}$ depending on $\bar{\rho}$).
Let $B_{1}:=[A_{2},A_{1}]A_{1}^{-6}[A_{1}^{-1},A_{2}]$, $B_{2}:=[A_{3},A_{4}][A_{4},A_{3}^{-1}]$. Then $B=1$ is equivalent to $B_{1}=B_{2}$, where $B=A_{1}^{3}[A_{1},A_{2}][A_{3},A_{4}][A_{4},A_{3}^{-1}][A_{2},A_{1}^{-1}]A_{1}^{3}$. Let $B_{k}(i,j)$ be the $(i,j)$-entry of the matrix $B_{k}$. Then $B_{1}=B_{2}$ is equivalent to the equalities $B_{1}(1,2)=B_{2}(1,2)$ and $B_{1}(2,1)=B_{2}(2,1)$.
Consider the $1$-cocycle $\beta$ defined by $\bar{\rho}$. We distinguish the following cases:
1) $\beta(x_{1})\neq 0$. We choose $u=x_{1}$, so that $b=1$, and we write $b^{\prime}=\tilde{\bar{b^{\prime}}}+\delta_{b}^{\prime}$, where $\tilde{\bar{b^{\prime}}}$ is the Teichmüller lifting and $\delta_{b}^{\prime}\in\mathfrak{m}_{R}$.
2) $\beta(x_{3})\neq 0=\beta(x_{1})$. We choose $u=x_{3}$, so that $b^{\prime}=1$ and $b\in\mathfrak{m}_{R}$.
3) $\beta=0$. Then $U$ is trivial and $b,b^{\prime}\in\mathfrak{m}_{R}$.
Proposition 2.1.3.
According to the above three cases, we have
1) $\tilde{R}=\mathcal{O}[[a,a^{\prime},\delta_{b}^{\prime},c,c^{\prime},d,d^{\prime}]]/(B_{1}(1,2)-B_{2}(1,2),B_{1}(2,1)-B_{2}(2,1))$, where we replace $b$ by $1$ and $b^{\prime}$ by $\tilde{\bar{b^{\prime}}}+\delta_{b}^{\prime}$ in the expressions of the entries.
2) $\tilde{R}=\mathcal{O}[[a,a^{\prime},b,c,c^{\prime},d,d^{\prime}]]/(B_{1}(1,2)-B_{2}(1,2),B_{1}(2,1)-B_{2}(2,1))$, where we replace $b^{\prime}$ by $1$ in the expressions of the entries.
3) $\tilde{R}=\mathcal{O}[[a,a^{\prime},b,b^{\prime},c,c^{\prime},d,d^{\prime}]]/(B_{1}(1,2)-B_{2}(1,2),B_{1}(2,1)-B_{2}(2,1))$.
Proof.
See [Böc10, Theorem 4.1]. ∎
Using this, we can study the ordinary deformation ring. Suppose $R_{v}^{\triangle}$ pro-represents the functor from $\mathfrak{U}_{\mathcal{O}}$ to the category of sets sending $R$ to the set of pairs $(\rho_{R},\psi_{R})$ such that
  • $\rho_{R}:G_{F_{v}}\to\mathrm{GL}_{2}(R)$ is a lifting of $\bar{\rho}$ with determinant $\chi$.
  • $\psi_{R}:G_{F_{v}}\to R^{\times}$ is a lifting of $\omega$ such that $\rho_{R}$ has a unique $G_{F_{v}}$-stable rank one $R$-submodule of $R^{2}$ with $G_{F_{v}}$-action via $\psi_{R}$.
Proposition 2.1.4.
According to the above three cases, we have
1) $R_{v}^{\triangle}\cong\mathcal{O}[[\delta_{b}^{\prime},d^{\prime},x,y,z]]$ if $\beta(x_{1})\neq 0$.
2) $R_{v}^{\triangle}\cong\mathcal{O}[[b,d,x,y,z]]$ if $\beta(x_{3})\neq 0=\beta(x_{1})$.
3) $R_{v}^{\triangle}\cong\mathcal{O}[[b,b^{\prime},d,d^{\prime},x,y]]/(3b-bd-b^{\prime}d^{\prime})$ if $\beta=0$.
In each case, the ring $R_{v}^{\triangle}$ is a normal domain of dimension $6$.
Proof.
We only prove the first case, and the other two cases are similar.
Note that $R_{v}^{\square}$ is the universal framed deformation (lifting) ring. By [KW09b, Proposition 2.1], we have $R_{v}^{\square}\cong\mathcal{O}[[\delta_{b}^{\prime},c,c^{\prime},d,d^{\prime},x,y,z]]/(B_{1}(1,2)-B_{2}(1,2),B_{1}(2,1)-B_{2}(2,1))$ as $a,a^{\prime}$ are uniquely determined by the determinant $\chi$. Therefore, we have $R_{v}^{\triangle}\cong R_{v}^{\square}/(c,c^{\prime})$. Then our result follows from an easy calculation. ∎
Remark 2.1.5.
This proposition is not used in the proof of our main theorem. See Remark 5.3.5.
2.2. Pseudo-representations
In this subsection, we will give some results on 2-dimensional pseudo-representations. Our main references are [SW99] and [Pan22].
Definition 2.2.1.
For a profinite group $G$ and a topological commutative ring $R$ in which 2 is invertible, a 2-dimensional pseudo-representation is a continuous function $T:G\to R$ such that
1) $T(1)=2$,
2) $T(\sigma\tau)=T(\tau\sigma)$ for all $\sigma,\tau\in G$,
3) $T(\gamma\delta\eta)+T(\gamma\eta\delta)-T(\gamma\eta)T(\delta)-T(\eta\delta)T(\gamma)-T(\delta\gamma)T(\eta)+T(\gamma)T(\delta)T(\eta)=0,$ for any $\delta,\gamma,\eta\in G$.
The determinant $\operatorname{det}(T)$ of a pseudo-representation $T$ is a character (using 3)) defined by
$\operatorname{det}(T):G\to R^{\times},~{}\operatorname{det}(T)(\sigma)=\frac{1}{2}(T(\sigma)^{2}-T(\sigma^{2})),\sigma\in G.$
If there exists an order $2$ element $c\in G$ such that $T(c)=0$, then for a pseudo-representation $T$ and $\sigma,\tau\in G$, we define:
1) $a(\sigma)=\frac{1}{2}(T(c\sigma)+T(\sigma))$,
2) $d(\sigma)=\frac{1}{2}(-T(c\sigma)+T(\sigma))$,
3) $y(\sigma,\tau)=a(\sigma\tau)-a(\sigma)a(\tau).$
If $\rho:G\to\operatorname{GL}_{2}(R)$ is a continuous representation with an element $c\in G$ satisfying $\rho(c)=\begin{pmatrix}1&~{}\\ ~{}&-1\end{pmatrix}$, then $\operatorname{tr}(\rho)$ is a pseudo-representation. More explicitly, if $\rho(\sigma)=\begin{pmatrix}a_{\sigma}&b_{\sigma}\\ c_{\sigma}&d_{\sigma}\end{pmatrix}$, then $a(\sigma)=a_{\sigma}$, $d(\sigma)=d_{\sigma}$ are as defined above and $y(\sigma,\tau)=b_{\sigma}c_{\tau}$.
Pan [Pan22, 2.1.4] verifies that the continuous functions $\{a,d,y\}$ satisfy the following equations.
(1) $y(\sigma,\tau)=d(\tau\sigma)-d(\tau)d(\sigma)$.
(2) $y(\sigma\tau,\delta)=a(\sigma)y(\tau,\delta)+y(\sigma,\delta)d(\tau)$.
(3) $y(\sigma,\tau\delta)=a(\delta)y(\sigma,\tau)+y(\sigma,\delta)d(\tau)$.
(4) $y(\alpha,\beta)y(\sigma,\tau)=y(\alpha,\tau)y(\sigma,\beta)$.
The following result is clear from the last equation.
Proposition 2.2.2.
[Zha24, Proposition 2.1.2] Let $R$ be an integral domain, and $T:G\to R$ be a pseudo-representation for a profinite group $G$. Let $\alpha,\beta$ be two elements in $G$.
1) If $y(\alpha,\beta)=0$ in $R$, then either $y(\alpha,\tau)=0$ or $y(\tau,\beta)=0$ for any $\tau\in G$.
2) Assume that the set $J_{\beta}=\{\tau\in G:y(\beta,\tau)\neq 0\}$ (resp. $J^{\prime}_{\beta}=\{\tau\in G:y(\tau,\beta)\neq 0\}$) is non-empty. For $\tau\in J_{\beta}$ (resp. $J^{\prime}_{\beta}$), let $u(\alpha,\beta)(\tau)=\frac{y(\alpha,\tau)}{y(\beta,\tau)}$ (resp. $u^{\prime}(\alpha,\beta)(\tau)=\frac{y(\tau,\alpha)}{y(\tau,\beta)}$) be an element in the fraction field of $R$. Then we have $u(\alpha,\beta)(\tau)=u(\alpha,\beta)(\tau^{\prime})$ (resp. $u^{\prime}(\alpha,\beta)(\tau)=u^{\prime}(\alpha,\beta)(\tau^{\prime})$) for any $\tau,\tau^{\prime}\in J_{\beta}$ (resp. $J^{\prime}_{\beta}$). (Afterwards, we will use notations $u(\alpha,\beta)$ and $u^{\prime}(\alpha,\beta)$ for simplicity if they are well-defined.)
Definition 2.2.3.
[Pan22, 2.1.5] Assume that $R$ is either a field or a DVR. For a pseudo-representation $T:G\to R$, a representation $\rho$ associated to $T$ is in the following form.
1) $\rho(\sigma)=\begin{pmatrix}a(\sigma)&~{}\\ ~{}&d(\sigma)\end{pmatrix}$, if all $y(\sigma,\tau)=0$. We call this case reducible.
2) Choose $\sigma_{0},\tau_{0}$ such that $\frac{y(\sigma,\tau)}{y(\sigma_{0},\tau_{0})}\in R$ for any $\sigma,\tau$, if $y(\sigma,\tau)\neq 0$ for some $\sigma,\tau$. Define $\rho(\sigma)=\begin{pmatrix}a(\sigma)&\frac{y(\sigma,\tau_{0})}{y(\sigma_{0},\tau_{0})}\\ y(\sigma_{0},\sigma)&d(\sigma)\end{pmatrix}$. We call this case irreducible.
The next result is helpful to study the representation constructed in the previous definition.
Proposition 2.2.4.
Let $R$ be a CNL domain with maximal ideal $\mathfrak{m}_{R}$ and residue field $\mathbb{F}$ satisfying $\operatorname{char}\mathbb{F}=p>0$. Let $G$ be a profinite group and $T:G\to R$ be a pseudo-representation of $G$. Assume that $T~{}\operatorname{mod}~{}\mathfrak{m}_{R}$ is reducible. Let $S$ be a finite subset (not necessarily a group) of $G$. Then there exist a partition of $S=S_{1}\amalg S_{2}$, a positive integer $n>1$ satisfying $(n,p)=1$ and a CNL domain $R^{\prime}$ satisfying the following conditions.
a) $R^{\prime}$ is a quotient of $R$.
b) For any $\theta\in S_{1}$, we have $y(\theta,\alpha)=0$ for any $\alpha\in G$ in $R^{\prime}$. For any $\theta^{\prime},\theta^{\prime\prime}\in S_{2}$, we have $y(\theta^{\prime},\alpha)^{n}=y(\theta^{\prime\prime},\alpha)^{n}$ for any $\alpha\in G$ in $R^{\prime}$.
c) For any $\alpha\in G,~{}\theta\in S_{2}$, either $y(\theta,\beta)=0$ for any $\beta\in G$ in $R^{\prime}$ or $u(\alpha,\theta)$ is well-defined and integral over $R^{\prime}$.
d) We have $\operatorname{dim}R^{\prime}\geq\operatorname{dim}R-|S|$. If $S_{2}$ is not empty, then further $\operatorname{dim}R^{\prime}\geq\operatorname{dim}R-|S|+1$.
Proof.
This is [Zha24, Corollary 2.1.7]. ∎
Now suppose the group $G$ satisfies Mazur’s finiteness condition $\Phi_{p}$ (see [Pan22, 2.1.2 Assumption 1]).
Let $\mathcal{O}$ be a complete DVR of characteristic 00 with residue field $\mathbb{F}$. Let $T_{\mathbb{F}}:G\to\mathbb{F}$ be a $2$-dimensional pseudo-representation. In our case, the functor sending each object $R$ of $\mathfrak{U}_{\mathcal{O}}$ to the set of pseudo-representations $T:G\to R$ which lift $T_{\mathbb{F}}$ is pro-represented by a CNL $\mathcal{O}$-algebra $R^{\mathrm{ps}}_{T_{\mathbb{F}}}$.
Suppose $\mathfrak{p}$ is a one-dimensional prime of $R^{\mathrm{ps}}_{T_{\mathbb{F}}}$. Let $T(\mathfrak{p})$ be the pseudo-deformation associated to the prime $\mathfrak{p}$. Consider the functor $D_{\mathfrak{p}}$ from the category of Artinian local $k(\mathfrak{p})$-algebras with residue field $k(\mathfrak{p})$ to the category of sets which sends $A$ to the set of $2$-dimensional pseudo-deformations over $A$ lifting $T(\mathfrak{p})$. Using Definition 2.2.3, we can construct a representation $\rho(\mathfrak{p}):G\to\mathrm{GL}_{2}(k(\mathfrak{p}))$ associated to $T(\mathfrak{p})\otimes k(\mathfrak{p})$.
Proposition 2.2.5.
1) The deformation problem $D_{\mathfrak{p}}$ is pro-represented by $\widehat{(R^{\mathrm{ps}}_{T_{\mathbb{F}}})_{\mathfrak{p}}}$ with universal pseudo-representation $G\to R^{\mathrm{ps}}_{T_{\mathbb{F}}}\to\widehat{(R^{\mathrm{ps}}_{T_{\mathbb{F}}})_{\mathfrak{p}}}$.
2) Suppose that $\rho(\mathfrak{p})$ is absolutely irreducible. Let $D_{\rho(\mathfrak{p})}$ be the functor from the category of Artinian local $k(\mathfrak{p})$-algebras with residue field $k(\mathfrak{p})$ to the category of sets which sends $A$ to the set of deformations of $\rho(\mathfrak{p})$ to $A$. Then $D_{\rho(\mathfrak{p})}$ is pro-represented by $\widehat{(R^{\mathrm{ps}}_{T_{\mathbb{F}}})_{\mathfrak{p}}}$.
Proof.
See [Pan22, Proposition 2.2.1 & Corollary 2.2.2]. ∎
2.3. Local pseudo-deformation rings
In this subsection, we collect some results about local universal pseudo-deformation rings.
Let $p$ be a prime. Let $F_{v}$ be a finite extension of $\mathbb{Q}_{p}$. Let $K$ be an another finite extension of $\mathbb{Q}_{p}$ containing all the roots of unity of $p$-power order in $F_{v}$, with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$. Let $\chi:G_{F_{v}}\to\mathcal{O}^{\times}$ be a continuous character.
Let $\bar{D}=1+\bar{\chi}$ be a pseudo-representation. Let $D^{\operatorname{ps}}:\mathfrak{U}_{\mathcal{O}}\to\operatorname{Sets}$ be the functor such that for $(A,\mathfrak{m}_{A})\in\mathfrak{U}_{\mathcal{O}}$, $D^{\operatorname{ps}}(A)$ is the set of continuous $A$-valued $2$-dimensional pseudo-representation $D:G_{F_{v}}\to A$ satisfying $(D\mod\mathfrak{m}_{A})=\bar{D}$. The functor is pro-represented by a CNL $\mathcal{O}$-algebra $(R_{v}^{\operatorname{ps}},\mathfrak{m}_{R_{v}^{\operatorname{ps}}})$. Note that in our definition of $R_{v}^{\operatorname{ps}}$, we do not fix the determinant of the pseudo-representation.
Proposition 2.3.1.
Assume that $\bar{D}$ is multiplicity free, i.e., $\bar{\chi}$ is non-trivial.
1) The ring $R_{v}^{\operatorname{ps}}$ is reduced and $\mathcal{O}$-torsion free.
2) The ring $R_{v}^{\operatorname{ps}}/(\varpi)$ is equidimensional of dimension $4[F_{v}:\mathbb{Q}_{p}]+1$.
3) Let $R_{v}^{\mathrm{ps},\chi}$ be the universal pseudo-deformation ring parametrizing all pseudo-deformations of $\bar{D}$ with determinant $\chi$. Then $R_{v}^{\mathrm{ps},\chi}$ is an $\mathcal{O}$-torsion free integral domain.
Proof.
For 1), see [BIP23, Corollary 4.28]. For 2), see [BJ23, Theorem 5.5.1]. For 3), see [BIP23, Corollary 5.11]. ∎
For the rest of this subsection, we assume $p=3$, $F_{v}=\mathbb{Q}_{p}$ and $\bar{\chi}=\omega=\omega^{-1}$.
Let $R_{v}^{\operatorname{ps,ord}}$ be the quotient of $R_{v}^{\operatorname{ps}}$ parametrizing all reducible liftings of $1+\bar{\chi}$, i.e., $y(\sigma,\tau)=0$ for any $\sigma,\tau\in G_{F_{v}}$.
Proposition 2.3.2.
1) The ring $R_{v}^{\operatorname{ps}}$ is of dimension $6$.
2) The ring $R_{v}^{\operatorname{ps,ord}}$ is a power series ring over $\mathcal{O}$ of relative dimension $4$.
Proof.
1) By Krull’s principal ideal theorem and 2) of Proposition 2.3.1, $R_{v}^{\operatorname{ps}}$ is of dimension $5$ or $6$. If $\operatorname{dim}R_{v}^{\operatorname{ps}}=5$, then $\varpi$ is a zero divisor in $R_{v}^{\operatorname{ps}}$, which is against to 1) of Proposition 2.3.1.
2) Let $R_{1}$ and $R_{\bar{\chi}}$ be the universal deformation ring of the one-dimensional representation $1$ and $\bar{\chi}$, respectively. By [BIP23, page 6], each of them is a power series ring over $\mathcal{O}$ of relative dimension $2$. Let $R=R_{1}\hat{\otimes}_{\mathcal{O}}R_{\bar{\chi}}$ and it is a power series ring over $\mathcal{O}$ of relative dimension $4$. By the universal property, we get a natural surjection $R\twoheadrightarrow R_{v}^{\operatorname{ps,ord}}$.
Recall the universal framed deformation ring $\tilde{R}$ in the third case of Proposition 2.1.3, i.e., the cocycle $\beta=0$. Then we get natural maps $R_{v}^{\operatorname{ps}}\to\tilde{R}\twoheadrightarrow\tilde{R}/(b,b^{\prime},c,c^{\prime})\cong\mathcal{O}[[a,a^{\prime},d,d^{\prime}]]$. From the description in the previous subsection, the last ring defines a natural reducible pseudo-representation. Hence, we have natural maps $R_{v}^{\operatorname{ps}}\twoheadrightarrow R_{v}^{\operatorname{ps,ord}}\to\mathcal{O}[[a,a^{\prime},d,d^{\prime}]]$.
By the universal property, we get a natural surjection $R\to\mathcal{O}[[a,a^{\prime},d,d^{\prime}]]$. Since each of them is a power series ring over $\mathcal{O}$ of relative dimension $4$, the surjection is actually an isomorphism. Now we consider the natural maps $R\twoheadrightarrow R_{v}^{\operatorname{ps,ord}}\to\mathcal{O}[[a,a^{\prime},d,d^{\prime}]]$. The composite map is an isomorphism, so the natural surjection $R\twoheadrightarrow R_{v}^{\operatorname{ps,ord}}$ is an isomorphism. ∎
The following result is an analogue to [Pan22, Lemma 7.4.4].
Proposition 2.3.3.
The kernel of the natural surjection $R_{v}^{\operatorname{ps}}\twoheadrightarrow R_{v}^{\operatorname{ps,ord}}$ can be generated by two elements.
Proof.
For a Noetherian local ring $R$, we write $\mathfrak{t}_{R}$ as the $(R/\mathfrak{m}_{R})$-vector space $\mathfrak{m}_{R}/\mathfrak{m}_{R}^{2}$.
Using [BJ23, (36) page 73] or [Bel12, Theorem 2], we have $\operatorname{dim}_{\mathbb{F}}\mathfrak{t}_{R_{v}^{\operatorname{ps}}/(\varpi)}=6$. Then we get natural surjections $R_{6}:=\mathcal{O}[[x_{1},...,x_{6}]]\twoheadrightarrow R_{v}^{\operatorname{ps}}\twoheadrightarrow R_{v}^{\operatorname{ps,ord}}$. We just need to show that the kernel of the composite surjection can be generated by $2$ elements.
Using 2) of Proposition 2.3.2, we find that both $R_{6}$ and $R_{v}^{\operatorname{ps,ord}}$ are regular local rings. Then by [Sta24, Lemma 10.106.4], we know that the kernel can be generated by $\dim R_{6}-\dim R_{v}^{\operatorname{ps,ord}}=2$ elements. This shows the result.
Actually, the universal pseudo-deformation ring $R_{v}^{\operatorname{ps}}$ may not be regular in a more general setting (see [BJ23] for more details). For example, in our case, as $\varpi$ is non-zero in $\mathfrak{t}_{R_{v}^{\operatorname{ps,ord}}}$, we know that $\operatorname{dim}_{\mathbb{F}}\mathfrak{t}_{R_{v}^{\operatorname{ps}}}=7$, but $\dim R_{v}^{\operatorname{ps}}=6$. However, we may use the following result to study its singular locus.
Proposition 2.3.4.
Let $l$ be an odd prime. Let $\kappa$ be a local field with residue field of characteristic $l$ or a finite field of characteristic $l$. Let $D_{i}:G_{\mathbb{Q}_{l}}\to\kappa^{\times},i=1,2$ be continuous characters of dimension $1$. Let $D=D_{1}\oplus D_{2}$, and let $R_{D}^{\operatorname{ps}}$ be the universal pseudo-deformation ring. If $D_{1}\neq D_{2}(m)$ for $m\in\{0,\pm 1\}$, then $R_{D}^{\operatorname{ps}}$ is regular.
Proof.
For the case $\kappa$ a finite field or a local field of characteristic $l$, see [BJ23, Lemma 5.5.4]. For the case $\kappa$ a local field of characteristic 00, see [Che11, Théorème]. ∎
The following result is an analogue to [Pan22, Lemma 7.4.8].
Proposition 2.3.5.
Let $\mathfrak{q}$ be prime of $R_{v}^{\operatorname{ps,ord}}$ of dimension $1$, and let $\mathfrak{q}^{\prime}$ be the inverse image of $\mathfrak{q}$ in $R_{v}^{\operatorname{ps}}$. Suppose that the pseudo-deformation $D_{\mathfrak{q}}=\psi_{1}+\psi_{2}$ associated to $\mathfrak{q}$ satisfies $\psi_{1}\neq\psi_{2}(\pm 1)$. Then the kernel of the natural surjection $(R_{v}^{\operatorname{ps}})_{\mathfrak{q}^{\prime}}\twoheadrightarrow(R_{v}^{\operatorname{ps,ord}})_{\mathfrak{q}}$ is principal.
Proof.
Note that in our case, the residual pseudo-representation is multiplicity-free, so $\psi_{1}\neq\psi_{2}$. Using Proposition 2.2.5 and Proposition 2.3.4, we know that $\widehat{(R_{v}^{\mathrm{ps}})_{\mathfrak{q}^{\prime}}}$ is a regular local ring. Combining Proposition 2.3.2, $(R_{v}^{\operatorname{ps}})_{\mathfrak{q}^{\prime}}$ is a regular local ring (see [Mat80, Corollary 28.2 & Proposition 28.2]) of dimension $5$. Note that $R_{v}^{\operatorname{ps,ord}}$ is a power series ring. By Proposition 2.3.2 and [Mat80, Corollary 18.3], $(R_{v}^{\operatorname{ps,ord}})_{\mathfrak{q}}$ is a regular local ring of dimension $4$. Using [Sta24, Lemma 10.106.4], we know that the kernel can be generated by one element. This shows the result.
2.4. Galois cohomology
In this subsection, we give some results about Galois cohomology. More precisely, we will give the local and global characteristic formulae and the Greenberg-Wiles formula in the case that the residue field is a $p$-adic local field.
Let $p$ be a prime. Let $K$ be a $p$-adic local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$. Let $G$ be a profinite group, and $V$ a finite dimensional $K$-vector space with continuous $G$-action. At first, we give the Euler-Poincaré characteristic formulae in the version for $V$. In the case that $V$ is a finite $G$-module, one can see [NSW08, Theorem 7.3.1] and [NSW08, 8.7.4].
Theorem 2.4.1 (Local Characteristic Formula).
Let $l$ be a prime. Let $E$ be an $l$-adic field, and suppose $G=G_{E}$.
1) For any $i\geq 0$, we have $\operatorname{dim}_{K}H^{i}(G_{E},V)<\infty$. If further $i\geq 3$, then we have $\operatorname{dim}_{K}H^{i}(G_{E},V)=0$.
2) If $l=p$, we have $\sum_{i=0}^{2}(-1)^{i}\operatorname{dim}_{K}H^{i}(G_{E},V)=-[E:\mathbb{Q}_{p}]\cdot\operatorname{dim}_{K}V$. If $l\neq p$, we have $\sum_{i=0}^{2}(-1)^{i}\operatorname{dim}_{K}H^{i}(G_{E},V)=0.$
Proof.
See [BBD+13, page 78, Theorem 3.9.1]. For the case $l=p$ and $K$ a local field of characteristic $p$, this result is also true and one can see [BJ23, Theorem 3.4.1]. ∎
Theorem 2.4.2 (Global Characteristic Formula).
Let $E$ be a number field, and $E_{S}$ the maximal extension of $E$ unramified outside a finite set of places $S$, containing the set of archimendean places $S_{\infty}$, and the finite places over $p$. Suppose that $K$ is a finite extension of $\mathbb{Q}_{p}$ and $G=G_{E,S}=\operatorname{Gal}(E_{S}/E)$.
1) For any $i\geq 0$, we have $\operatorname{dim}_{K}H^{i}(G_{E,S},V)<\infty$. If further $i\geq 3$, then we have $\operatorname{dim}_{K}H^{i}(G_{E,S},V)=0$.
2) We have $\sum_{i=0}^{2}(-1)^{i}\operatorname{dim}_{K}H^{i}(G_{E,S},V)=\sum_{v\in S_{\infty}}\operatorname{dim}_{K}V^{G_{E_{v}}}-[E:\mathbb{Q}]\cdot\operatorname{dim}_{K}V$.
Proof.
See [Kis03, Lemma 9.7]. ∎
Now let $F$ be a number field, and $T$ a finite set of places of $F$ containing the finite places over $p$. Suppose $G=G_{F,T}$. Let $M$ be a $G$-stable $\mathcal{O}$-lattice of $V$, and $\rho:G\to\operatorname{GL}_{n}(\mathcal{O})$ the corresponding Galois representation, where $n=\operatorname{dim}_{K}V$. For simplicity, we write $\operatorname{ad}^{0}_{m}:=\operatorname{ad}^{0}\rho\otimes\mathcal{O}/\varpi^{m}$ and $\operatorname{ad}_{m}:=\operatorname{ad}\rho\otimes\mathcal{O}/\varpi^{m}$.
We define the following complex as an analogue to [Gee22, Section 3.23] (or [CHT08, page 21]). Let $\partial:C^{i}(G,\operatorname{ad}^{0}_{m})\to C^{i+1}(G,\operatorname{ad}^{0}_{m})$ be the usual coboundary map. Define $C^{i}_{T,\mathrm{loc}}(G,\operatorname{ad}^{0}_{m})$ by
$C^{0}_{T,\mathrm{loc}}(G,\operatorname{ad}^{0}_{m})=\oplus_{v\in T}C^{0}(G_{F_{v}},\operatorname{ad}_{m}),~{}~{}C^{1}_{T,\mathrm{loc}}(G,\operatorname{ad}^{0}_{m})=\oplus_{v\in T}C^{1}(G_{F_{v}},\operatorname{ad}^{0}_{m}),$
$C^{i}_{T,\mathrm{loc}}(G,\operatorname{ad}^{0}_{m})=\oplus_{v\in T}C^{i}(G_{F_{v}},\operatorname{ad}^{0}_{m}),i\geq 2.$
Let $C^{0}_{0}(G,\operatorname{ad}^{0}_{m}):=C^{0}(G,\operatorname{ad}_{m})$, and let $C^{i}_{0}(G,\operatorname{ad}^{0}_{m})=C^{i}(G,\operatorname{ad}^{0}_{m})$ for $i>0$. Then we let $H^{i}_{T}(G,\operatorname{ad}^{0}_{m})$ denote the cohomology of the complex
$C^{i}_{T}(G,\operatorname{ad}^{0}_{m}):=C_{0}^{i}(G,\operatorname{ad}^{0}_{m})\oplus C^{i-1}_{T,\mathrm{loc}}(G,\operatorname{ad}^{0}_{m}),$
where the coboundary map is given by $(\phi,(\psi_{v}))\mapsto(\partial\phi,(\phi|_{G_{F_{v}}}-\partial\psi_{v}))$. Now we get an exact sequence of complexes
$0\to C^{i-1}_{T,\mathrm{loc}}(G,\operatorname{ad}^{0}_{m})\to C^{i}_{T}(G,\operatorname{ad}^{0}_{m})\to C^{i}_{0}(G,\operatorname{ad}^{0}_{m})\to 0,$
and the corresponding long exact sequence in cohomology is
$0\to H^{0}_{T}(G,\operatorname{ad}^{0}_{m})\to H^{0}(G,\operatorname{ad}_{m})\to\bigoplus_{v\in T}H^{0}(G_{F_{v}},\operatorname{ad}_{m})\to$
$H^{1}_{T}(G,\operatorname{ad}^{0}_{m})\to H^{1}(G,\operatorname{ad}^{0}_{m})\to\bigoplus_{v\in T}H^{1}(G_{F_{v}},\operatorname{ad}^{0}_{m})\to$
$H^{2}_{T}(G,\operatorname{ad}^{0}_{m})\to H^{2}(G,\operatorname{ad}^{0}_{m})\to\bigoplus_{v\in T}H^{2}(G_{F_{v}},\operatorname{ad}^{0}_{m})\to H^{3}_{T}(G,\operatorname{ad}^{0}_{m})\to\cdots.$
We can also replace $\operatorname{ad}^{0}_{m}$ by $\operatorname{ad}^{0}\rho\otimes K$ and obtain exact sequences similarly.
From Theorem 2.4.1 and Theorem 2.4.2, we have
$H^{i}(G,\operatorname{ad}^{0}\rho\otimes K)=0,~{}~{}~{}H^{i}(G_{F_{v}},\operatorname{ad}^{0}\rho\otimes K)=0,~{}~{}i\geq 3,v\in T.$
Hence, we get an exact sequence
$H^{1}(G,\operatorname{ad}^{0}\rho\otimes K)\to\bigoplus_{v\in T}H^{1}(G_{F_{v}},\operatorname{ad}^{0}\rho\otimes K)\to H^{2}_{T}(G,\operatorname{ad}^{0}\rho\otimes K)\to$
$H^{2}(G,\operatorname{ad}^{0}\rho\otimes K)\to\bigoplus_{v\in T}H^{2}(G_{F_{v}},\operatorname{ad}^{0}\rho\otimes K)\to H^{3}_{T}(G,\operatorname{ad}^{0}\rho\otimes K)\to 0.$
Using the Poitou-Tate duality ([NSW08, Theorem 8.6.14]) and the Five Lemma, we have $H^{2}_{T}(G,\operatorname{ad}^{0}\rho)\cong H^{1}(G,(\operatorname{ad}^{0}\rho)^{\prime})^{\vee}$, where $(\operatorname{ad}^{0}\rho)^{\prime}=\operatorname{Hom}(\operatorname{ad}^{0}\rho,\mu_{p^{\infty}})$. Hence, we have
$\operatorname{dim}_{K}H^{2}_{T}(G,\operatorname{ad}^{0}\rho\otimes K)=\operatorname{rank}_{\mathcal{O}}H^{1}(G,(\operatorname{ad}^{0}\rho)^{\prime})=\operatorname{dim}_{K}H^{1}(G,(\operatorname{ad}^{0}\rho)^{\prime}\otimes K).$
[…]
Now we can conclude a generalized Greenberg-Wiles formula. For the usual version, one can see [CHT08, Lemma 2.3.4].
Proposition 2.4.3.
Suppose that $\rho\otimes K$ is irreducible, and $H^{0}(G,\operatorname{ad}^{0}_{m}(1))=0$ holds for any sufficiently large $m$. Then we have
$\operatorname{dim}_{K}H^{1}_{T}(G,\operatorname{ad}^{0}\rho\otimes K)-\operatorname{dim}_{K}H^{1}(G,(\operatorname{ad}^{0}\rho)^{\prime}\otimes K)=-1+|T|-\sum_{v\mid\infty}\operatorname{dim}_{K}H^{0}(G_{F_{v}},\operatorname{ad}^{0}\rho\otimes K).$
Proof.
As $T$ is not empty, we have $H^{0}_{T}(G,\operatorname{ad}^{0}\rho\otimes K)=0$. Note that we have
$\dim_{K}H^{0}(G,\operatorname{ad}\rho\otimes K)-\dim_{K}H^{0}(G,\operatorname{ad}^{0}\rho\otimes K)=1,$
$\dim_{K}H^{0}(G_{F_{v}},\operatorname{ad}\rho\otimes K)-\dim_{K}H^{0}(G_{F_{v}},\operatorname{ad}^{0}\rho\otimes K)=1,~{}~{}~{}v\in T.$
Then our result follows from Theorem 2.4.1, Theorem 2.4.2 and the previous discussions. ∎
2.5. Global Galois deformation theory
In this subsection, we study some global deformation rings. Our aim is to construct the Galois deformation theory in the case that the residue field is a $p$-adic local field instead of a finite field. For the classical theory, we recommend Böckle’s paper in [BBD+13] to readers for reference.
Let $p$ be an odd prime. Let $F$ be a totally real field such that $p$ splits completely, and $T$ a finite set of places of $F$ containing the finite places over $p$. Let $K$ be a $p$-adic local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$. Let $\chi:G_{F,T}\to\mathcal{O}^{\times}$ be a continuous character, and we suppose that $\bar{\chi}$ is totally odd.
Let $R^{\mathrm{ps}}$ be the universal pseudo-deformation ring parametrizing all liftings of the pseudo-representation $1+\bar{\chi}$ with determinant $\chi$. Let $\mathfrak{p}\in\operatorname{Spec}R^{\mathrm{ps}}$ (we do not require $\varpi\notin\mathfrak{p}$ here) and suppose that it defines an irreducible pseudo-representation. Then by Definition 2.2.3, we can construct an absolutely irreducible Galois representation $\rho(\mathfrak{p})^{\circ}:G_{F,T}\to\operatorname{GL}_{2}(A)$, where $A$ is the normal closure of $R^{\mathrm{ps}}/\mathfrak{p}$. Write $\bar{\rho}_{b}=\rho(\mathfrak{p})^{\circ}\mod\mathfrak{m}_{A}$, where $0\neq b\in H^{1}(G_{F,T},\mathbb{F}(\bar{\chi}^{-1}))$. Then by [SW99, Lemma 2.13], up to some non-zero scalars of $\mathbb{F}$, $b$ is independent of the choice of $\rho(\mathfrak{p})^{\circ}$ given in Definition 2.2.3.
In general, it is difficult to study some geometric properties of $\operatorname{Spec}R^{\mathrm{ps}}$. However, we can use Proposition 2.2.5 to study $\operatorname{Spec}R^{\mathrm{ps}}_{\mathfrak{p}}$ in the classical way.
Definition 2.5.1.
[Pan22, Definition 5.4.3] Let $R$ be a CNL ring. The connectedness dimension of $R$ is defined to be
$c(R):=\min_{Z_{1},Z_{2}}\{\dim(Z_{1}\cap Z_{2})\},$
where $Z_{1},Z_{2}$ are non-empty unions of irreducible components of $\operatorname{Spec}R$ such that $Z_{1}\cup Z_{2}=\operatorname{Spec}R$.
Proposition 2.5.2.
1) $\operatorname{dim}R^{\mathrm{ps}}_{\mathfrak{p}}\geq 2[F:\mathbb{Q}]$.
2) The connectedness dimension $c(R^{\mathrm{ps}}_{\mathfrak{p}})$ is at least $2[F:\mathbb{Q}]-1$.
Proof.
For the first part, see [Pan22, Lemma 7.1.2]. For the second part, see [Pan22, Lemma 7.4.6]. ∎
In the rest of this subsection, we study the universal deformation ring of the Galois representation $\rho(\mathfrak{p}):=\rho(\mathfrak{p})^{\circ}\otimes k(\mathfrak{p})$. Suppose that $\rho(\mathfrak{p})$ is ordinary at each $v|p$ and $\varpi\notin\mathfrak{p}$.
For any finite place $v\in T$, denote by $D_{v}^{\square,\chi}:\mathfrak{U}_{k(\mathfrak{p})}\to\operatorname{Sets}$ the functor such that for $(B,\mathfrak{m}_{B})\in\mathfrak{U}_{k(\mathfrak{p})}$, $D_{v}^{\square,\chi}(B)$ is the set of continuous representations $\rho_{B}:G_{F_{v}}\to\operatorname{GL}_{2}(B)$ satisfying $\rho_{B}\mod\mathfrak{m}_{B}=\rho(\mathfrak{p})|_{G_{F_{v}}}$ and $\operatorname{det}\rho_{B}=\chi|_{G_{F_{v}}}$. Using the same proof as [BBD+13, page 27, Proposition 1.3.1], we can show that this functor is pro-representable by some $R_{v,k(\mathfrak{p})}^{\square,\chi}$, which is a CNL $k(\mathfrak{p})$-algebra.
Similarly, for any $v|p$, we can also define the universal ordinary (framed) deformation ring $R_{v,k(\mathfrak{p})}^{\triangle,\chi}$ parametrizing all ordinary liftings of $\rho(\mathfrak{p})|_{G_{F_{v}}}$.
The following result is standard using local characteristic formula Theorem 2.4.1.
Proposition 2.5.3.
Let $v$ be a finite place in the set $T$.
1) If $v\nmid p$, then we have $\operatorname{dim}R_{v,k(\mathfrak{p})}^{\square,\chi}\geq 3$.
2) If $v|p$, then we have $\operatorname{dim}R_{v,k(\mathfrak{p})}^{\triangle,\chi}\geq 5$.
Proof.
The proof is almost the same as the case that the residue field is finite. In that case, one can see [BBD+13, page 64, Theorem 3.3.1] for $v\nmid p$ and [SW99, Lemma 2.2] for $v\mid p$. ∎
Let $S\subset T$ be a finite set of primes of $F$. Define the functor $D_{k(\mathfrak{p})}^{\square_{S}}$ from $\mathfrak{U}_{k(\mathfrak{p})}$ to the category of sets sending $R$ to the set of tuples $(\rho_{R},\alpha_{v})_{v\in S}$ modulo the equivalence relation $\sim_{S}$ where
  • $\rho_{R}:G_{F,T}\to\mathrm{GL}_{2}(R)$ is a lifting of $\rho(\mathfrak{p})$ to $R$ with determinant $\chi$ such that $(\rho_{R})|_{G_{F_{v}}}$ is the ordinary lifting of $\rho(\mathfrak{p})|_{G_{F_{v}}}$ for any $v|p$.
  • $\alpha_{v}\in 1+M_{2}(\mathfrak{m}_{R}),v\in S$. Here $\mathfrak{m}_{R}$ is the maximal ideal of $R$.
  • $(\rho_{R};\alpha_{v})_{v\in S}\sim_{S}(\rho^{\prime}_{R};\alpha^{\prime}_{v})_{v\in S}$ if there exists an element $\beta\in 1+M_{2}(\mathfrak{m}_{R})$ with $\rho^{\prime}_{R}=\beta\rho_{R}\beta^{-1},\alpha^{\prime}_{v}=\beta\alpha_{v}$ for any $v\in M$.
Then the functor $D_{k(\mathfrak{p})}^{\square_{S}}$ is pro-representable by a similar argument of [BBD+13, page 102, Proposition 5.1.1]. Note that in the case $S=\varnothing$, we need to use the fact that $\rho(\mathfrak{p})$ is absolutely irreducible. We write $R_{k(\mathfrak{p})}^{\square_{S},\chi}$ for the universal $S$-framed deformation ring and $R_{k(\mathfrak{p})}^{\operatorname{univ}}$ in the case $S=\varnothing$. Then the following result is also standard.
Proposition 2.5.4.
Write $R_{\operatorname{loc}}=(\hat{\bigotimes}_{v|p}R_{v,k(\mathfrak{p})}^{\triangle,\chi})\hat{\otimes}_{k(\mathfrak{p})}(\hat{\bigotimes}_{v\nmid p,v\in T}R_{v,k(\mathfrak{p})}^{\square,\chi})$.
1) $R_{k(\mathfrak{p})}^{\square_{T},\chi}\cong R_{k(\mathfrak{p})}^{\operatorname{univ}}[[x_{1},...,x_{4|T|-1}]]$.
2) There exists an isomorphism $R_{\operatorname{loc}}[[y_{1},...,y_{r}]]/(f_{1},...,f_{t})\cong R_{k(\mathfrak{p})}^{\square_{T},\chi}$, where
$r=\operatorname{dim}_{k(\mathfrak{p})}H^{1}_{T}(G,\operatorname{ad}^{0}\rho\otimes k(\mathfrak{p})),~{}~{}~{}~{}t\leq\operatorname{dim}_{k(\mathfrak{p})}H^{1}(G,(\operatorname{ad}^{0}\rho)^{\prime}\otimes k(\mathfrak{p})).$
Proof.
The proof is also the same as the case that the residue field is finite. For 1), one can see [BBD+13, page 103, Corollary 5.2.1]. For 2), one can see [KW09b, Lemma 4.4 & 4.6]. ∎
Corollary 2.5.5.
Assume $H^{0}(G_{F,T},\operatorname{ad}^{0}\rho(\mathfrak{p})(1))=0$. We have $\operatorname{dim}R_{k(\mathfrak{p})}^{\operatorname{univ}}\geq[F:\mathbb{Q}]$.
Proof.
At first, by Proposition 2.5.3 and Krull’s principal ideal theorem, we have $\dim R_{\mathrm{loc}}\geq 3|T|+2[F:\mathbb{Q}]$. Then combining the second part of the previous proposition and Proposition 2.4.3, we have
$\dim R_{k(\mathfrak{p})}^{\square_{T},\chi}\geq 3|T|+2[F:\mathbb{Q}]+(-1+|T|-\sum_{v\mid\infty}\operatorname{dim}_{K}H^{0}(G_{F_{v}},\operatorname{ad}^{0}\rho\otimes K))=4|T|-1+[F:\mathbb{Q}].$
Here the last equality is because $\bar{\chi}$ is totally odd. Using the first part of the previous proposition, we have $\operatorname{dim}R_{k(\mathfrak{p})}^{\operatorname{univ}}\geq[F:\mathbb{Q}]$. ∎
Similar to $R_{k(\mathfrak{p})}^{\operatorname{univ}}$, we can also define the universal deformation ring $R_{b}^{\operatorname{univ}}$ parametrizing all liftings of $\bar{\rho}_{b}$ which are ordinary at each $v|p$, and with determinant $\chi$. Recall that $A$ is the normal closure of $R^{\mathrm{ps}}/\mathfrak{p}$, and hence its fraction field is $k(\mathfrak{p})$. By the universal property, we get a homomorphism $R_{b}^{\operatorname{univ}}\to A$. Let $\mathfrak{q}$ be the kernel of this map, and we assume the residue field of $\widehat{(R_{b}^{\operatorname{univ}})_{\mathfrak{q}}}$ is $K^{\prime}$, which is contained in $k(\mathfrak{p})$. Let $\rho_{k(\mathfrak{p})}^{\operatorname{univ}}$ (resp. $\rho_{b}^{\operatorname{univ}}$) be the universal lifting. Then $\hat{\rho}=\rho_{b}^{\operatorname{univ}}\otimes_{R_{b}^{\operatorname{univ}}}(\widehat{(R_{b}^{\operatorname{univ}})_{\mathfrak{q}}}\otimes_{K^{\prime}}k(\mathfrak{p}))$ defines an ordinary lifting of $\rho(\mathfrak{p})$. By the universal property, we get a natural map $R_{k(\mathfrak{p})}^{\operatorname{univ}}\to\widehat{(R_{b}^{\operatorname{univ}})_{\mathfrak{q}}}\otimes_{K^{\prime}}k(\mathfrak{p})$.
Proposition 2.5.6.
The map $f:R_{k(\mathfrak{p})}^{\operatorname{univ}}\to\widehat{(R_{b}^{\operatorname{univ}})_{\mathfrak{q}}}\otimes_{K^{\prime}}k(\mathfrak{p})$ is actually an isomorphism.
Proof.
Using [BJ23, Proposition 3.1.5], we just need to show that $f$ is formally étale. The proof is nearly the same as [Kis03, Proposition 9.5] (also see [BJ23, Theorem 3.3.1]), and we briefly sketch it here. For simplicity for notations, we may suppose $K=K^{\prime}=k(\mathfrak{p})$ (using [BJ23, Lemma 3.2.6]) and write $\widehat{R}=\widehat{(R_{b}^{\operatorname{univ}})_{\mathfrak{q}}}\otimes_{K^{\prime}}k(\mathfrak{p})$.
We consider a commutative diagram with $C$ an Artinian local $K$-algebra and $I$ a square zero ideal. We need to show that there exists a unique dashed arrow $g$ such that the two triangular subdiagrams commute.
[…]
By possibly conjugating $\hat{\rho}$ by some matrix $X\in\mathrm{GL}_{2}(\widehat{R})$ satisfying $X\equiv\operatorname{id}\mod\mathfrak{m}_{\widehat{R}}$, we can suppose that $\rho_{K}^{\operatorname{univ}}\otimes_{R_{K}^{\operatorname{univ}}}\widehat{R}=\hat{\rho}$.
Choose a finitely generated free $\mathcal{O}$-subalgebra $C^{\circ}\subset C$ surjecting onto $\mathcal{O}$ under the natural surjection $c:C\twoheadrightarrow K$. Set $\mathfrak{n}=\operatorname{ker}c$ and $\mathfrak{n}^{\circ}=\mathfrak{n}\cap C^{\circ}$. For any $n\geq 0$, define $C_{n}^{\circ}=\sum_{j>0}\varpi^{-nj}(\mathfrak{n}^{\circ})^{j}+C^{\circ}\subset C$. Note that the sum is finite since $\mathfrak{n}^{\circ}$ is nilpotent. Then we have $\cup_{n>0}C_{n}^{\circ}=c^{-1}(\mathcal{O})$. We write $(C/I)_{n}^{\circ}=C_{n}^{\circ}/(I\cap C_{n}^{\circ})$. As $G_{F,T}$ is compact, we know that $\rho_{K}^{\operatorname{univ}}\otimes_{R_{K}^{\operatorname{univ}}}C$ factors through $\operatorname{GL}_{2}(C_{n}^{\circ})$ for $n$ sufficiently large.
Note that $R_{b}^{\operatorname{univ}}$ is topologically finitely generated, the composite $R_{b}^{\operatorname{univ}}\to\widehat{R}\to C/I$ factors through $(C/I)_{n}^{\circ}$ for $n$ sufficiently large.
From our discussions above, we know that for sufficiently large $n$, $\rho_{K}^{\operatorname{univ}}\otimes_{R_{K}^{\operatorname{univ}}}C$ lies in $\operatorname{GL}_{2}(C_{n}^{\circ})$, and it reduces to $\rho_{b}^{\operatorname{univ}}\otimes_{R_{b}^{\operatorname{univ}}}(C/I)_{n}^{\circ}$ modulo $(I\cap C_{n}^{\circ})$. By the universal property, there exists a unique map $g_{n}^{\circ}:R_{b}^{\operatorname{univ}}\to C_{n}^{\circ}$ such that $\rho_{b}^{\operatorname{univ}}\otimes_{R_{b}^{\operatorname{univ}}}C_{n}^{\circ}$ is equivalent to $\rho_{K}^{\operatorname{univ}}\otimes_{R_{K}^{\operatorname{univ}}}C$. Taking the completion, we obtain a map $g:\widehat{R}\to C$, and we can check that the diagram commuts with such $g$. The uniqueness of $g$ follows from the uniqueness of $g_{n}^{\circ}$. ∎
Combining Corollary 2.5.5 and Proposition 2.5.6, we obtain the following result.
Corollary 2.5.7.
Assume $H^{0}(G_{F,T},\operatorname{ad}^{0}\rho(\mathfrak{p})(1))=0$. Then there exists an irreducible component of $\operatorname{Spec}R_{b}^{\operatorname{univ}}$ containing $\mathfrak{q}$ of dimension at least $1+[F:\mathbb{Q}]$.
3. Local-global compatibility
In this section, our goal is to generalize Pan’s local-global compatibility results (see [Pan22, Theorem 3.5.3, Corollary 3.5.8 & 3.5.10]) to cover the case $p=3$. It is based on the recent work [PT21].
[…]
3.1. Completed cohomology
In this subsection, we recall the definition of completed cohomology and the relation between it and classical automorphic forms. For more details, one can see [Pan22, Section 3.1 & 3.2].
Let $D$ be a quaternion algebra with centre $F$ which is ramified at all infinite places of $F$ and unramified at all places above $p$, and we fix isomorphisms $D\otimes F_{v}\cong M_{2}(F_{v})$ for any $v$ where $D$ is unramified. We view $K_{p}=\prod_{v\mid p}\operatorname{GL}_{2}(O_{F_{v}})$ and $D_{p}^{\times}=\prod_{v\mid p}\operatorname{GL}_{2}(F_{v})$ as subgroups of $(D\otimes_{F}\mathbb{A}_{F})^{\times}$. We also write $N_{D/F}:(D\otimes_{F}\mathbb{A}_{F})^{\times}\to\mathbb{A}_{F}^{\times}$ as the reduced norm.
Let $A$ be a topological $\mathcal{O}$-algebra and $U=\prod_{v\nmid\infty}U_{v}$ be an open compact subgroup of $(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}$ such that $U_{v}\subseteq\operatorname{GL}_{2}(O_{F_{v}})$ for $v\mid p$. We write $U^{p}=\prod_{v\nmid p}U_{v}$ (tame level) and $U_{p}=\prod_{v\mid p}U_{v}$. Let $\psi:(\mathbb{A}_{F}^{\infty})^{\times}/F_{+}^{\times}\to A^{\times}$ be a continuous character, where $F_{+}$ is the set of totally positive elements in $F$. Let $\tau:\prod_{v\mid p}U_{v}\to\operatorname{Aut}(W_{\tau})$ be a continuous representation on a finite $A$-module $W_{\tau}$, and we also view $\tau$ as a representation of $U$ by projecting to $\prod_{v\mid p}U_{v}$.
Let $S_{\tau,\psi}(U,A)$ be the space of continuous functions:
$f:D^{\times}\setminus(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}\to W_{\tau}$
such that for any $g\in(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times},u\in U,z\in(\mathbb{A}_{F}^{\infty})^{\times}$, we have
  • $f(gu)=\tau(u^{-1})(f(g))$,
  • $f(gz)=\psi(z)f(g)$.
Write $(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}=\bigsqcup_{i\in I}D^{\times}t_{i}U(\mathbb{A}_{F}^{\infty})^{\times}$ for some finite set $I$ and $t_{i}\in(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}$. If $\tau^{-1}|_{U\cap(\mathbb{A}_{F}^{\infty})^{\times}}$ acts as $\psi|_{U\cap(\mathbb{A}_{F}^{\infty})^{\times}}$, then there is an isomorphism:
$S_{\tau,\psi}(U,A)\simeq\bigoplus_{i\in I}W_{\tau}^{(t_{i}^{-1}D^{\times}t_{i}\cap U(\mathbb{A}_{F}^{\infty})^{\times})/F^{\times}},$
by sending $f$ to $(f(t_{i}))_{i}$. We say $U$ is sufficiently small if $(t_{i}^{-1}D^{\times}t_{i}\cap U(\mathbb{A}_{F}^{\infty})^{\times})/F^{\times}$ is trivial for all $i\in I$. This can be achieved by shrinking $U_{v}$ for some $v$.
Suppose that $\xi:U^{p}\to A^{\times}$ is a continuous smooth character such that $\psi\mid_{\prod_{v\nmid p}(O_{F_{v}}^{\times})\cap U_{v}}=\xi\mid_{\prod_{v\nmid p}(O_{F_{v}}^{\times})\cap U_{v}}$. We define $S_{\tau,\psi,\xi}(U,A)$ as the space of continuous functions: $f:D^{\times}\setminus(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}\to W_{\tau}$, such that for any $g\in(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}$, $z\in(\mathbb{A}_{F}^{\infty})^{\times}$, $u=u^{p}u_{p}\in U$, we have
$f(guz)=\psi(z)\xi(u^{p})\tau(u_{p}^{-1})(f(g)).$
Similarly, if $\psi\mid_{U_{p}\cap\mathcal{O}_{F,p}^{\times}}=\tau^{-1}\mid_{U_{p}\cap\mathcal{O}_{F,p}^{\times}}$, then we also have
$S_{\tau,\psi,\xi}(U,A)\simeq\bigoplus_{i\in I}W_{\tau}^{(t_{i}^{-1}D^{\times}t_{i}\cap U(\mathbb{A}_{F}^{\infty})^{\times})/F^{\times}},$
where $I=D^{\times}\setminus(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}/U(\mathbb{A}_{F}^{\infty})^{\times}$ and $\{t_{i}\}_{i\in I}$ is a set of representatives.
We will ignore $\tau$ in the notation if it is the trivial action on $A$.
[…]
Note that the definition of $S_{\psi}(U,A)$ makes sense for any topological $\mathcal{O}$-module $A$. We will also use $S_{\psi}(U,A)$ for any topological $\mathcal{O}$-module $A$ by abuse of notation.
Definition 3.1.1.
The completed cohomology of tame level $U^{p}$ is defined to be
$S_{\psi}(U^{p}):=\operatorname{Hom}_{\mathcal{O}}(K/\mathcal{O},S_{\psi}(U^{p},K/\mathcal{O}))$
equipped with $p$-adic topology, and the completed homology is defined to be
$M_{\psi}(U^{p}):=S_{\psi}(U^{p},K/\mathcal{O})^{\vee}=\operatorname{Hom}_{\mathcal{O}}(S_{\psi}(U^{p},K/\mathcal{O}),K/\mathcal{O})$
equipped with compact-open topology.
We can also define $S_{\psi,\xi}(U^{p})$ and $M_{\psi,\xi}(U^{p})$ in a similar way.
From the definition above, there is a natural action of $D_{p}^{\times}=\prod_{v|p}\mathrm{GL}_{2}(F_{v})$ on all spaces defined above by right translation. It is almost by definition that for any open compact subgroup $U_{p}=\prod_{v|p}U_{v}\subseteq K_{p}=\prod_{v|p}\mathrm{GL}_{2}(O_{F_{v}})$,
$S_{\psi}(U^{p},A)^{U_{p}}=S_{\psi}(U^{p}U_{p},A).$
Hence $S_{\psi}(U^{p},K/\mathcal{O})$ is a smooth admissible $\mathcal{O}$-representation of $D_{p}^{\times}$. It is also clear that $S_{\psi}(U^{p})$ is $\varpi$-torsion free and
[…]
Remark 3.1.2.
[…]
$W_{(\vec{k},\vec{w}),K}=\bigotimes_{\sigma:F\to K}(\operatorname{Sym}^{k_{\sigma}-2}(K^{2})\otimes\det{}^{w_{\sigma}}),$
where $\operatorname{Sym}^{k_{\sigma}-2}$ denotes the space of homogeneous polynomials of degree $k_{\sigma}-2$ in two variables with an action of $\mathrm{GL}_{2}(F_{v(\sigma)})$ given by
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}(f)(X,Y)=f(\sigma(a)X+\sigma(c)Y,\sigma(b)X+\sigma(d)Y).$
Here $v(\sigma)$ is the place above $p$ given by $\sigma$. Let $\psi:(\mathbb{A}^{\infty}_{F})^{\times}/F^{\times}_{+}\to K^{\times}$ be a continuous character such that $\tau_{(\vec{k},\vec{w})}^{-1}|_{U\cap(\mathbb{A}_{F}^{\infty})^{\times}}=\psi|_{U\cap(\mathbb{A}_{F}^{\infty})^{\times}}$ and write $\tau=\tau_{(\vec{k},\vec{w})}$. For simplicity, we will write $S_{(\vec{k},\vec{w}),\psi}(U,K)$ for $S_{\tau_{(\vec{k},\vec{w})},\psi}(U,K)$ from now on. Then there is an isomorphism:
$S_{(\vec{k},\vec{w}),\psi}(U,K)\otimes_{K,\iota_{p}}\mathbb{C}$ $\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}$ $\operatorname{Hom}_{D_{\infty}^{\times}}(W_{\iota_{p}(\vec{k},\vec{w}),\mathbb{C}}^{*},C^{\infty}(D^{\times}\setminus(D\otimes\mathbb{A}_{F})^{\times}/U,\psi_{\mathbb{C}})),$
[…]
where
  • $D_{\infty}^{\times}=(D\otimes_{\mathbb{Q}}\mathbb{R})^{\times}$,
  • $W_{\iota_{p}(\vec{k},\vec{w}),\mathbb{C}}=W_{(\vec{k},\vec{w}),K}\otimes_{K,\iota_{p}}\mathbb{C}$ is viewed as an algebraic representation $\tau_{\mathbb{C}}$ of $D_{\infty}^{\times}$ (induced by $\iota_{p}$), and $W_{\iota_{p}(\vec{k},\vec{w}),\mathbb{C}}^{*}$ is its $\mathbb{C}$-linear dual.
  • $\psi_{\mathbb{C}}:\mathbb{A}_{F}^{\times}/F^{\times}\to\mathbb{C}^{\times}$ sends $g$ to $N_{F/\mathbb{Q}}(g_{\infty})^{1-w}\iota_{p}(N_{F/\mathbb{Q}}(g_{p})^{w-1}\psi(g^{\infty}))$.
  • $C^{\infty}(D^{\times}\setminus(D\otimes\mathbb{A}_{F})^{\times}/U,\psi_{\mathbb{C}})$ is the space of smooth $\mathbb{C}$-valued functions on $D^{\times}\setminus(D\otimes\mathbb{A}_{F})^{\times}$, right invariant by $U$ and with central character $\psi_{\mathbb{C}}$.
Note that the right hand side is a subspace of automorphic forms on $(D\otimes\mathbb{A}_{F})^{\times}$.
Furthermore, write $S_{\psi}(U^{p})_{K}$ for $S_{\psi}(U^{p})\otimes_{\mathcal{O}}K$, which is a Banach space with unit ball $S_{\psi}(U^{p})$. Then we have an isomorphism
$S_{(\vec{k},\vec{w}),\psi}(U^{p}U_{p},K)\simeq\operatorname{Hom}_{K[U_{p}]}(W^{*}_{(\vec{k},\vec{w}),K},S_{\psi}(U^{p})_{K}).$
See [Pan22, 3.1.2 & 3.2.4].
Let $v|p$ be a finite place of $F$. We assume $\psi|_{N_{D/F}(U^{p})}$ is trivial and $\psi|_{O_{F_{v}}^{\times}}$ is an algebraic character. In particular, there exists an integer $m$ such that, $\psi(a_{v})=\sigma_{v}(a_{v})^{m},~{}a_{v}\in O_{F_{v}}^{\times}$, where $\sigma_{v}:F\to K$ is the embedding induced by $v$. Let $U^{v}$ be an open compact subgroup of $\prod_{w\neq v,w|p}\mathrm{GL}_{2}(O_{F_{w}})$.
Definition 3.1.3.
The subspace of $\mathrm{GL}_{2}(O_{F_{v}})U^{v}$-algebraic vectors of $S_{\psi}(U^{p})_{K}$ is defined to be the image of the evaluation map:
$\bigoplus_{(\vec{k},\vec{w})}\operatorname{Hom}_{K[\mathrm{GL}_{2}(O_{F_{v}})U^{v}]}(W_{(\vec{k},\vec{w}),K},S_{\psi}(U^{p})_{K})\otimes_{K}W_{(\vec{k},\vec{w}),K}\to S_{\psi}(U^{p})_{K},$
where the sum is taken over all $(\vec{k},\vec{w})$ with $k_{\sigma}+2w_{\sigma}=m+2$ for any $\sigma$. The subspace $S_{\psi}(U^{p})_{K}^{v-\mathrm{a},v^{\prime}-\mathrm{la}}$ of $\mathrm{GL}_{2}(O_{F_{v}})$-algebraic, $\prod_{w\neq v,w|p}\mathrm{GL}_{2}(O_{F_{w}})$-locally algebraic vectors of $S_{\psi}(U^{p})_{K}$ is defined to be the union of all $\mathrm{GL}_{2}(O_{F_{v}})U^{v}$-algebraic vectors where $U^{v}$ runs through all open compact subgroups of $\prod_{w\neq v,w|p}\mathrm{GL}_{2}(O_{F_{w}})$.
The following result is helpful for our local-global compatibility argument.
Proposition 3.1.4.
The subspace $S_{\psi}(U^{p})_{K}^{v-\mathrm{a},v^{\prime}-\mathrm{la}}$ is dense in $S_{\psi}(U^{p})_{K}$.
Proof.
See [Pan22, Proposition 3.2.6]. ∎
3.2. Hecke algebras
In this subsection, we recall the definition of the $p$-adic big Hecke algebra. Our reference is [Pan22, Section 3.3].
Let $A$ be a Noetherian topological $\mathbb{Z}_{p}$-algebra, and consider a level $U$, a character $\psi$ and a representation $\tau:U_{p}\to\operatorname{Aut}(W_{\tau})$. Let $\Sigma$ be a finite set of primes of $F$ containing all places above $p$ and places $v$ where either $D$ is ramified or $U_{v}$ is not a maximal open subgroup. For any $v\notin\Sigma$, we define the Hecke operator $T_{v}\in\operatorname{End}(S_{\tau,\psi}(U,A))$ to be the double coset action $[U_{v}\begin{pmatrix}\varpi_{v}&~{}\\ ~{}&1\end{pmatrix}U_{v}]$. Precisely, if we write $U_{v}\begin{pmatrix}\varpi_{v}&~{}\\ ~{}&1\end{pmatrix}U_{v}=\coprod_{i}\gamma_{i}U_{v}$, then
$(T_{v}\cdot f)(g)=\sum_{i}f(g\gamma_{i}),~{}~{}f\in S_{\tau,\psi}(U,A).$
We define the Hecke algebra $\mathbb{T}_{\tau,\psi}^{\Sigma}(U,A)\subseteq\operatorname{End}(S_{\tau,\psi}(U,A))$ to be the $A$-subalgebra generated by all $T_{v},v\notin\Sigma$. This is a finite commutative $A$-algebra. We will write $\mathbb{T}_{\psi}^{\Sigma}(U,A)$ for the Hecke algebra if $\tau$ is the trivial action on $A$. Suppose that $A=\mathcal{O}/\varpi^{n}$, $U$ is sufficiently small and $\psi\mid_{U\cap(\mathbb{A}_{F}^{\infty})^{\times}}$ is trivial modulo $\varpi^{n}$. Then we can check that the Hecke algebra $\mathbb{T}_{\psi}^{\Sigma}(U,\mathcal{O}/\varpi^{n})$ is independent of $\Sigma$, and we may drop $\Sigma$ in this situation.
Definition 3.2.1.
(Big Hecke algebra) Let $U^{p}$ be a tame level and $\psi:(\mathbb{A}_{F}^{\infty})^{\times}/(U^{p}\cap(\mathbb{A}_{F}^{\infty})^{\times})F_{+}^{\times}\to\mathcal{O}^{\times}$ be a continuous character. Then we define the big Hecke algebra
[…]
where $\mathcal{I}$ is the set of pairs $(n,U_{p})$ with $U_{p}\subseteq K_{p}$ and $n$ a positive integer such that $\psi\mid_{U_{p}\cap\mathcal{O}_{F,p}^{\times}}\equiv 1~{}\operatorname{mod}~{}\varpi^{n}$.
We conclude some basic properties of the big Hecke algebra.
Proposition 3.2.2.
Let $U^{p}$ be a tame level and $\psi:(\mathbb{A}_{F}^{\infty})^{\times}/(U^{p}\cap(\mathbb{A}_{F}^{\infty})^{\times})F_{+}^{\times}\to\mathcal{O}^{\times}$ be a continuous character.
1) $\mathbb{T}_{\psi}(U^{p})$ is semi-local. (Write $\mathfrak{m}_{1},\cdots,\mathfrak{m}_{r}$ for all the maximal ideals of $\mathbb{T}_{\psi}(U^{p})$.)
2) $\mathbb{T}_{\psi}(U^{p})\cong\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}_{1}}\times\cdots\times\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}_{r}}$ and each $\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}_{i}}$ is $\mathfrak{m}_{i}$-adically complete and separated.
3) $S_{\psi}(U^{p})\cong S_{\psi}(U^{p})_{\mathfrak{m}_{1}}\oplus\cdots\oplus S_{\psi}(U^{p})_{\mathfrak{m}_{r}}$ and each $S_{\psi}(U^{p})_{\mathfrak{m}_{i}}$ is $\mathfrak{m}_{i}$-adically complete and separated.
4) $\mathbb{T}_{\psi}(U^{p})$ commutes with base change. In other words, if $K^{\prime}$ is a finite extension of $K$ with ring of integers $\mathcal{O}^{\prime}$ and replace $\mathcal{O}$ by $\mathcal{O}^{\prime}$ in all the definitions, then we have $\mathbb{T}^{\prime}_{\psi}(U^{p})\cong\mathbb{T}_{\psi}(U^{p})\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$.
Proof.
See [Pan22, Proposition 3.3.2, Corollary 3.3.3, 3.3.5]. ∎
Suppose $\mathbb{T}_{\psi}(U^{p})$ is non-zero and $\mathfrak{m}\in\operatorname{Spec}\mathbb{T}_{\psi}(U^{p})$ is a maximal ideal. Enlarging $\mathcal{O}$ if necessary, we may assume $\mathfrak{m}$ has residue field $\mathbb{F}$. Denote by $T_{\mathfrak{m}}$ the two-dimensional pseudo-representation over $\mathbb{F}$ (see [Pan22, 3.3.3 & 3.3.4] for the existence):
$G_{F,\Sigma}\longrightarrow\mathbb{T}^{\psi}(U^{p})\to\mathbb{T}^{\psi}(U^{p})/\mathfrak{m}=\mathbb{F}.$
For any $v|p$, let $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ be the universal deformation ring which parametrizes all two-dimensional pseudo-representations of $G_{F_{v}}$ lifting $T_{\mathfrak{m}}|_{G_{F_{v}}}$ with determinant $\psi\varepsilon^{-1}|_{G_{F_{v}}}$. By the universal property, we have a natural map $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}\to\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}$. Taking tensor products over all $v|p$, we get
$R^{\mathrm{ps},\psi\varepsilon^{-1}}_{p}:=\widehat{\bigotimes}_{v|p}R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}\to\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}.$
Therefore, we have defined an action of $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{p}$ on the completed cohomology $S_{\psi}(U^{p})_{\mathfrak{m}}$ via the map above.
Suppose that $\xi:U^{p}\to A^{\times}$ is a continuous smooth character such that $\psi\mid_{\prod_{v\nmid p}(O_{F_{v}}^{\times})\cap U_{v}}=\xi\mid_{\prod_{v\nmid p}(O_{F_{v}}^{\times})\cap U_{v}}$. Then we can also define the big Hecke algebra $\mathbb{T}_{\psi,\xi}(U^{p})$ and obtain the properties in the same way.
3.3. Local-global compatibility
In this subsection, we generalize Pan’s local-global compatibility result and give some applications. We will frequently use Paškūnas’ results and keep the notations given in the appendix.
Let $G=\prod_{i=1}^{f}\mathrm{GL}_{2}(\mathbb{Q}_{p})$ and $Z(G)\simeq\prod_{i=1}^{f}\mathbb{Q}_{p}^{\times}$ be its centre. (In our case, we may take $f=[F:\mathbb{Q}]$.) Let $\zeta:Z(G)\to\mathcal{O}^{\times}$ be a continuous character. We first recall Paškūnas theory for the $p$-adic analytic group $G$.
Let $\mathrm{Mod}_{G,\zeta}^{\mathrm{l.adm}}(\mathcal{O})$ be the full subcategory of $\mathrm{Mod}_{G,\zeta}^{\mathrm{sm}}(\mathcal{O})$ consisting of locally admissible representations. Here an object $V$ is locally admissible if for every $v\in V$, the smallest $\mathcal{O}[G]$-submodule of $V$ containing $v$ is admissible. Then by [Pan22, Lemma 3.4.1], the categories $\mathrm{Mod}_{G,\zeta}^{\mathrm{l.adm}}(\mathcal{O})$ and $\mathrm{Mod}_{G,\zeta}^{\mathrm{l.fin}}(\mathcal{O})$ are actually the same. Note that the categories $\mathrm{Mod}_{G,\zeta}^{\mathrm{sm}}(\mathcal{O})$ and $\mathrm{Mod}_{G,\zeta}^{\mathrm{pro\,aug}}(\mathcal{O})$ are anti-equivalent via the Pontryagin duality. We write $\mathfrak{C}_{G,\zeta}(\mathcal{O})$ as the full subcategory of $\mathrm{Mod}_{G,\zeta}^{\mathrm{pro\,aug}}(\mathcal{O})$ anti-equivalent to $\operatorname{Mod}^{\mathrm{l\,fin}}_{G,\zeta}(\mathcal{O})$.
Let $\mathrm{Irr}_{G,\zeta}$ be the set of irreducible representations in $\mathrm{Mod}_{G,\zeta}^{\mathrm{sm}}(\mathcal{O})$. Recall that a block is an equivalence class of the relation $\sim$, where $\pi\sim\pi^{\prime}$ if there exists $\pi_{1},\cdots,\pi_{n}\in\mathrm{Irr}_{G,\zeta}$ such that $\pi\cong\pi_{1}$, $\pi^{\prime}\cong\pi_{n}$, and for $11\leq i\leq n-$ 1, $\pi_{i}\cong\pi_{i+1}$ or $\operatorname{Ext}^{1}_{G}(\pi_{i},\pi_{i+1})\neq 0$ or $\operatorname{Ext}^{1}_{G}(\pi_{i+1},\pi_{i})\neq 0$. Then there exists a natural decomposition of $\mathrm{Mod}^{\mathrm{l\,adm}}_{G,\zeta}(\mathcal{O})$ with respect to the blocks:
$\mathrm{Mod}^{\mathrm{l\,adm}}_{G,\zeta}(\mathcal{O})\cong\prod_{\mathfrak{B}\in\mathrm{Irr}_{G,\zeta}/\sim}\mathrm{Mod}^{\mathrm{l\,adm}}_{G,\zeta}(\mathcal{O})^{\mathfrak{B}},$
where $\mathrm{Mod}^{\mathrm{l\,adm}}_{G,\zeta}(\mathcal{O})^{\mathfrak{B}}$ is the full subcategory of $\mathrm{Mod}^{\mathrm{l\,adm}}_{G,\zeta}(\mathcal{O})$ consisting of representations with all irreducible subquotients in $\mathfrak{B}$. Taking Pontryagin dual, this gives:
$\mathfrak{C}_{G,\zeta}(\mathcal{O})\cong\prod_{\mathfrak{B}\in\mathrm{Irr}_{G,\zeta}/\sim}\mathfrak{C}_{G,\zeta}(\mathcal{O})^{\mathfrak{B}}.$
For a block $\mathfrak{B}$, write $\pi_{\mathfrak{B}}=\bigoplus_{\pi\in\mathfrak{B}_{i}}\pi$, where $\mathfrak{B}_{i}$ is the set of isomorphism classes of elements of $\mathfrak{B}$. Suppose that it is a finite set. Let $\pi_{\mathfrak{B}}\hookrightarrow J_{\mathfrak{B}}$ be an injective envelope of $\pi_{\mathfrak{B}}$ in $\mathrm{Mod}^{\mathrm{l\,adm}}_{G,\zeta}(\mathcal{O})$. Its Pontryagin dual $P_{\mathfrak{B}}:=J_{\mathfrak{B}}^{\vee}$ is a projective envelope of $\pi^{\vee}\cong\bigoplus_{\pi\in\mathfrak{B}_{i}}\pi^{\vee}$ in $\mathfrak{C}_{G,\zeta}(\mathcal{O})$. Let $E_{\mathfrak{B}}:=\operatorname{End}_{\mathfrak{C}_{G,\zeta}(\mathcal{O})}(P_{\mathfrak{B}})\cong\operatorname{End}_{G}(J_{\mathfrak{B}}).$ This is a pseudo-compact ring. Recall that the functor $M\mapsto\operatorname{Hom}_{\mathfrak{C}_{G,\zeta}(\mathcal{O})}(P_{\mathfrak{B}},M)$ defines an anti-equivalence of categories between $\mathfrak{C}_{G,\zeta}(\mathcal{O})^{\mathfrak{B}}$ and the category of right pseudo-compact $E_{\mathfrak{B}}$-modules. An inverse functor is given by $\mathrm{m}\mapsto(\mathrm{m}\hat{\otimes}_{E_{\mathfrak{B}}}P_{\mathfrak{B}}).$
As the block of $\mathrm{GL}_{2}(\mathbb{Q}_{p})$ has been well-understood (see Appendix A.2), we can use the following results to study the blocks of $G=\prod_{i=1}^{f}\mathrm{GL}_{2}(\mathbb{Q}_{p})$.
Lemma 3.3.1.
1) For any absolutely irreducible representation $\pi$ in $\mathrm{Mod}_{G,\zeta}^{\mathrm{sm}}(\mathcal{O})$, there is a finite extension $\mathbb{F}^{\prime}/\mathbb{F}$ such that $\pi\otimes_{\mathbb{F}}\mathbb{F}^{\prime}$ is isomorphic to some $\bigotimes_{i=1}^{f}\pi_{i}$, where $\pi_{i}$ are absolutely irreducible $\mathrm{GL}_{2}(\mathbb{Q}_{p})$ representations over $\mathbb{F}^{\prime}$.
2) Keep $\pi_{i}$ in 1) and assume that $\mathbb{F}^{\prime}=\mathbb{F}$. Let $P_{r}\to\pi_{r}^{\vee}$ be a projective envelope of $\pi_{r}^{\vee}$ for $r=1,\cdots,f$. Then $P_{1}\hat{\otimes}\cdots\hat{\otimes}P_{f}\to\widehat{\bigotimes}_{i=1}^{f}\pi_{i}^{\vee}$ is a projective envelope of ^$\widehat{\bigotimes}_{i=1}^{f}\pi_{i}^{\vee}\cong(\bigotimes_{i=1}^{f}\pi_{i})^{\vee}$ in $\mathfrak{C}_{G,\zeta}(\mathcal{O})$.
3) Write $G=G_{1}\times G_{2}$, where each $G_{r}=\prod_{i=1}^{f_{r}}\mathrm{GL}_{2}(\mathbb{Q}_{p})$ with centre $Z_{r},~{}r=1,2$. Suppose $M_{r},N_{r}\in\mathfrak{C}_{G_{r},\zeta|_{Z_{r}}}(\mathcal{O}),~{}r=1,2$. Then there exists a natural isomorphism:
$\operatorname{Hom}_{\mathfrak{C}_{G,\zeta}(\mathcal{O})}(M_{1}\hat{\otimes}M_{2},N_{1}\hat{\otimes}N_{2})\cong\operatorname{Hom}_{\mathfrak{C}_{G_{1},\zeta|_{Z_{1}}}(\mathcal{O})}(M_{1},N_{1})\hat{\otimes}\operatorname{Hom}_{\mathfrak{C}_{G_{2},\zeta|_{Z_{2}}}(\mathcal{O})}(M_{2},N_{2}).$
4) Given $f$ blocks $\mathfrak{B}_{1},\cdots,\mathfrak{B}_{f}$ of $\mathrm{GL}_{2}(\mathbb{Q}_{p})$ that contain an absolutely irreducible representation, then
$\mathfrak{B}_{1}\otimes\cdots\otimes\mathfrak{B}_{f}:=\{\pi_{1}\otimes\cdots\otimes\pi_{f}|\pi_{r}\in\mathfrak{B}_{r},r=1,\cdots,f\}$
is a block of $G$.
Proof.
See [Pan22, Lemma 3.4.5, 3.4.6 & 3.4.7]. ∎
Using the previous lemma and the explicit description of the blocks of $\mathrm{GL}_{2}(\mathbb{Q}_{p})$ (see Appendix A.2), we know that for each block $\mathfrak{B}$ of $G$, it consists of a finite number of isomorphism classes of elements of $\mathfrak{B}$.
For the rest of this subsection, we suppose that $p$ is an odd prime.
For each block $\mathfrak{B}_{r}$ of $\mathrm{GL}_{2}(\mathbb{Q}_{p})$, we write $\bar{\rho}_{\mathfrak{B}_{r}}$ for the associated semi-simple $2$-dimensional representation of $G_{\mathbb{Q}_{p}}$ over $\mathbb{F}$, and $R^{\mathrm{ps},\zeta\varepsilon}_{\mathfrak{B}_{r}}$ for the universal deformation ring which parametrizes all $2$-dimensional pseudo-representations of $G_{\mathbb{Q}_{p}}$ lifting $\mathrm{tr}\bar{\rho}_{\mathfrak{B}_{r}}$ with determinant $\zeta\varepsilon$. Using Theorem A.2.1, we know that $R^{\mathrm{ps},\zeta\varepsilon}_{\mathfrak{B}_{r}}$ is exactly the centre of $E_{\mathfrak{B}_{r}}$, and hence we obtain a natural injection $f_{r}:R^{\mathrm{ps},\zeta\varepsilon}_{\mathfrak{B}_{r}}\hookrightarrow E_{\mathfrak{B}_{r}}$.
Proposition 3.3.2.
Let $\mathfrak{B}=\mathfrak{B}_{1}\otimes\cdots\otimes\mathfrak{B}_{f}$ be a block of $G$ such that each $\mathfrak{B}_{i}$ contains an absolutely irreducible representation. Then
  1. (1)
    $E_{\mathfrak{B}}\cong\widehat{\bigotimes}_{r=1}^{f}E_{\mathfrak{B}_{r}}$
  2. (2)
    The natural inclusion $f_{r}:R^{\mathrm{ps},\zeta\varepsilon}_{\mathfrak{B}_{r}}\hookrightarrow E_{\mathfrak{B}_{r}}$ induces a natural finite map
    ^$\widehat{\bigotimes}_{r=1}^{f}R^{\mathrm{ps},\zeta\varepsilon}_{\mathfrak{B}_{r}}\to E_{\mathfrak{B}},$
    which makes $E_{\mathfrak{B}}$ into a finitely generated module over ^$\widehat{\bigotimes}_{r=1}^{f}R^{\mathrm{ps},\zeta\varepsilon}_{\mathfrak{B}_{r}}$.
  3. (3)
    The centre of $E_{\mathfrak{B}}$ is Noetherian and $E_{\mathfrak{B}}$ is a finitely generated module over its centre.
Proof.
The first assertion follows from 2) and 3) of Lemma 3.3.1. The second one follows from Theorem A.2.1. The third one follows from the second one and the fact that the image of ^$\widehat{\bigotimes}_{r=1}^{f}R^{\mathrm{ps},\zeta\varepsilon}_{\mathfrak{B}_{r}}$ is in the centre of $E_{\mathfrak{B}}$. ∎
Keep the level $U$, the set $\Sigma$ and the character $\psi$ as in the previous subsection. Let $\mathfrak{m}$ be a maximal ideal of $\mathbb{T}_{\psi}(U^{p})$. We get a two-dimensional pseudo-representation with determinant $\psi\varepsilon^{-1}$:
$T_{\mathfrak{m}}:G_{F,\Sigma}\to\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}.$
After restricting it to $G_{F_{v}},v|p$, we can attach a two-dimensional semi-simple representation $\bar{\rho}_{\mathfrak{m},v}$ of $G_{F_{v}}$ over $\mathbb{F}$. Enlarging $\mathcal{O}$ if necessary, we assume that $\bar{\rho}_{\mathfrak{m},v}$ is a direct sum of absolutely irreducible representations. Let $\bar{\rho}^{\prime}_{\mathfrak{m},v}=\bar{\rho}_{\mathfrak{m},v}\otimes\varepsilon$ (as we use the normailzed Colmez’s functor $\mathbf{V}$ here). Based on the list given in Appendix A.2, we can define a block $\mathfrak{B}_{\mathfrak{m},v}$ of $\mathrm{GL}_{2}(F_{v})\cong\mathrm{GL}_{2}(\mathbb{Q}_{p})$ from $\bar{\rho}^{\prime}_{\mathfrak{m},v}$ (since $p$ splits completely in $F$). Let $\mathfrak{B}_{\mathfrak{m}}=\otimes_{v|p}\mathfrak{B}_{\mathfrak{m},v}$ be the block of $D_{p}^{\times}=\prod_{v|p}\mathrm{GL}_{2}(F_{v})$ (using Lemma 3.3.1). Note that it has central character $\psi$. We denote its projective generator by $P_{\mathfrak{B}_{\mathfrak{m}}}$.
Our main goal is to study the following objects:
$\mathrm{m}:=\operatorname{Hom}_{\mathfrak{C}_{D_{p}^{\times},\psi}(\mathcal{O})}(P_{\mathfrak{B}_{\mathfrak{m}}},M_{\psi}(U^{p})_{\mathfrak{m}}).$
These objects will be the Hecke modules for the patching arguments to prove some $R=\mathbb{T}$ theorems.
There are two actions of $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{p}:=\widehat{\bigotimes}_{v|p}R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ on (the Hecke module) $\mathrm{m}$:
  1. (1)
    $\tau_{\operatorname{Gal}}$: which comes from the action of $\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}$ on $M_{\psi}(U^{p})_{\mathfrak{m}}$. See the discussions at the end of the previous subsection.
  2. (2)
    $\tau_{\mathrm{Aut}}$: which comes from the action of ^$\widehat{\bigotimes}_{v|p}R^{\mathrm{ps},\psi\varepsilon}_{v}\cong\widehat{\bigotimes}_{v|p}R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ on $P_{\mathfrak{B}_{\mathfrak{m}}}$ via $E_{\mathfrak{B}_{\mathfrak{m}}}$ due to the map in (2) of Proposition 3.3.2. The natural isomorphisms between $R^{\mathrm{ps},\psi\varepsilon}_{v}$ and $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ are given by twisting the inverse of cyclotomic character. We fix this isomorphism from now on.
Theorem 3.3.3 (Local-global compatibility).
Keep notations and assumptions as above.
  1. (1)
    $M_{\psi}(U^{p})_{\mathfrak{m}}\in\mathfrak{C}_{D_{p}^{\times},\psi}(\mathcal{O})^{\mathfrak{B}_{\mathfrak{m}}}$.
  2. (2)
    Both actions $\tau_{\operatorname{Gal}},\tau_{\mathrm{Aut}}$ of $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{p}$ on $\mathrm{m}$ are the same.
Proof.
The proof of this theorem is nearly the same as the proof of [Pan22, Theorem 3.5.3], where we just need to replace Paškūnas’ results in [Paš13] by the ones in [PT21]. We sketch the proof here.
Let $v$ be a place above $p$. It suffices to prove
  1. (1)
    $M_{\psi}(U^{p})_{\mathfrak{m}}\in\mathfrak{C}_{\mathrm{GL}_{2}(F_{v}),\psi|_{F_{v}^{\times}}}(\mathcal{O})^{\mathfrak{B}_{\mathfrak{m},v}}$,
  2. (2)
    both actions $\tau_{\operatorname{Gal}},\tau_{\mathrm{Aut}}$ of $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ on $\operatorname{Hom}_{\mathfrak{C}_{\mathrm{GL}_{2}(F_{v}),\psi|_{F_{v}^{\times}}}(\mathcal{O})}(P_{\mathfrak{B}_{\mathfrak{m},v}},M_{\psi}(U^{p})_{\mathfrak{m}})$ are the same.
After twisting some character, we may assume $\psi$ is crystalline at $v$ of Hodge-Tate weight $w_{\psi}$. We may also shrink $U^{p}$ so that $\psi|_{U^{p}\cap(\mathbb{A}_{F}^{\infty})^{\times}}$ is trivial and $U^{p}K_{p}$ is sufficiently small.
Recall that for any $(\vec{k},\vec{w})\in\mathbb{Z}_{>1}^{\operatorname{Hom}(F,\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu)}\times\mathbb{Z}^{\operatorname{Hom}(F,\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu)}$ such that $k_{\sigma}+2w_{\sigma}=w_{\psi}+2$ independent of $\sigma$ and $U_{p}=\mathrm{GL}_{2}(O_{F_{v}})U^{v}$ with $U^{v}$ an open subgroup of $\prod_{w\neq v,w|p}\mathrm{GL}_{2}(O_{F_{w}})$, we have the following isomorphism (see Remark 3.1.2):
$S_{(\vec{k},\vec{w}),\psi}(U^{p}U_{p},K)\cong\operatorname{Hom}_{K[U_{p}]}(W^{*}_{(\vec{k},\vec{w}),K},S_{\psi}(U^{p})_{K}).$
Then we get a natural surjective map $t_{\psi}:\mathbb{T}_{\psi}(U^{p})[\frac{1}{p}]\to\mathbb{T}_{(\vec{k},\vec{w}),\psi}(U^{p}U_{p},K)$ sending $T_{w}$ to $T_{w}$ for $w\notin\Sigma$. Let $\mathfrak{p}$ be a prime ideal of $\mathbb{T}_{(\vec{k},\vec{w}),\psi}(U^{p}U_{p},K)\otimes_{K}\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu$, and it corresponds to an automorphic representation $\pi_{\mathfrak{p}}=\pi_{\mathfrak{p}}^{\infty}\otimes(\pi_{\mathfrak{p}})_{\infty}$ on $(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}$. Combining the discussions in Remark 3.1.2, we get
[…]
where the direct limit ranges over all open compact subgroups $U_{v}$ of $\mathrm{GL}_{2}(O_{F_{v}})$, and $(\pi_{\mathfrak{p}})_{v}$ is the local representation of $\pi_{\mathfrak{p}}$ at place $v$. Choose a model $\pi^{K(\mathfrak{p})}_{v}$ of $(\pi_{\mathfrak{p}})_{v}$ over a finite extension $K(\mathfrak{p})$ of $K$. Therefore, we get a map
$\Phi_{\mathfrak{p}}:W^{*}_{(\vec{k},\vec{w}),K}\otimes_{K}(\pi^{K(\mathfrak{p})}_{v})^{\oplus d(\mathfrak{p})}\to(S_{\psi}(U^{p})\otimes_{\mathcal{O}}K(\mathfrak{p}))[\mathfrak{p}],$
where we also view $\mathfrak{p}$ as a maximal ideal of $\mathbb{T}_{\psi}(U^{p})\otimes K(\mathfrak{p})$ via the surjection $t_{\psi}$. Then the image of $\Phi_{\mathfrak{p}}$ in $S_{\psi}(U^{p})\otimes_{\mathcal{O}}K(\mathfrak{p})$ contains the $\mathrm{GL}_{2}(O_{F_{v}})U^{v}$-algebraic vectors of $(S_{\psi}(U^{p})_{K}\otimes_{K}K(\mathfrak{p}))[\mathfrak{p}]$ (see Definition 3.1.3), and we denote its closure in $S_{\psi}(U^{p})\otimes_{\mathcal{O}}K(\mathfrak{p})$ by $\Pi(\mathfrak{p})$.
Let $\Pi_{\mathfrak{B}_{\mathfrak{m},v}}:=\operatorname{Hom}^{\mathrm{cont}}_{\mathcal{O}}(P_{\mathfrak{B}_{\mathfrak{m},v}},K)$. This is a Banach space with unit ball $\operatorname{Hom}^{\mathrm{cont}}_{\mathcal{O}}(P_{\mathfrak{B}_{\mathfrak{m},v}},\mathcal{O})$. The following lemma is an application of Proposition 3.1.4, and one can see [Pan22, Lemma 3.5.4] for the proof.
Lemma 3.3.4.
The inclusion map $\Pi(\mathfrak{p})\hookrightarrow S_{\psi}(U^{p})\otimes_{\mathcal{O}}K(\mathfrak{p})$ induces a natural injective map:
$\operatorname{Hom}^{\mathrm{cont}}_{K[\mathrm{GL}_{2}(F_{v})]}(S_{\psi}(U^{p})_{\mathfrak{m}}\otimes_{\mathcal{O}}K,\Pi_{\mathfrak{B}_{\mathfrak{m},v}})\hookrightarrow\prod_{U^{v}}\prod_{(\vec{k},\vec{w})}\prod_{\mathfrak{p}}\operatorname{Hom}^{\mathrm{cont}}_{K(\mathfrak{p})[\mathrm{GL}_{2}(F_{v})]}(\Pi(\mathfrak{p}),\Pi_{\mathfrak{B}_{\mathfrak{m},v}}\otimes K(\mathfrak{p})),$
where $U^{v}$ runs over all open subgroups of $\prod_{w\neq v,w|p}\mathrm{GL}_{2}(O_{F_{w}})$, the pair $(\vec{k},\vec{w})$ runs over all elements in $\mathbb{Z}_{>1}^{\operatorname{Hom}(F,\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu)}\times\mathbb{Z}^{\operatorname{Hom}(F,\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu)}$ such that $k_{\sigma}+2w_{\sigma}=w_{\psi}+2$ for any $\sigma$, and $\mathfrak{p}$ runs over all the maximal ideals of $\mathbb{T}_{(\vec{k},\vec{w}),\psi}(U^{p}\mathrm{GL}_{2}(O_{F_{v}})U^{v})_{\mathfrak{m}}\otimes\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu$.
Note that we have a natural isomorphism
$\operatorname{Hom}^{\mathrm{cont}}_{K[\mathrm{GL}_{2}(F_{v})]}(S_{\psi}(U^{p})_{\mathfrak{m}}\otimes K,\Pi_{\mathfrak{B}_{\mathfrak{m},v}})\cong\operatorname{Hom}_{\mathfrak{C}_{\mathrm{GL}_{2}(F_{v}),\psi|_{F_{v}^{\times}}}(\mathcal{O})}(P_{\mathfrak{B}_{\mathfrak{m},v}},M_{\psi}(U^{p})_{\mathfrak{m}})\otimes K.$
To prove the theorem, we just need to check $\tau_{\operatorname{Gal}}|_{R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}}=\tau_{\mathrm{Aut}}|_{R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}}$ on the module $\operatorname{Hom}^{\mathrm{cont}}_{K(\mathfrak{p})[\mathrm{GL}_{2}(F_{v})]}(\Pi(\mathfrak{p}),\Pi_{\mathfrak{B}_{\mathfrak{m},v}}\otimes K(\mathfrak{p}))$ using the previous lemma and the $p$-torsion freeness of $S_{\psi}(U^{p})$ (see the discussion under Definition 3.1.1). Now we fix such a $\mathfrak{p}$ and suppose that it comes from $\mathbb{T}_{(\vec{k},\vec{w}),\psi}(U^{p}\mathrm{GL}_{2}(O_{F_{v}})U^{v})_{\mathfrak{m}}\otimes\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu$. For simplicity, we just assume $K(\mathfrak{p})=K$ as our settings are compatible with base change.
We divide the proof into two cases.
Case 1: $\pi_{\mathfrak{p}}$ does not factor through the reduced norm $N_{D/F}$.
The action $\tau_{\operatorname{Gal}}$ of $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ on $\operatorname{Hom}^{\mathrm{cont}}_{K[\mathrm{GL}_{2}(F_{v})]}(\Pi(\mathfrak{p}),\Pi_{\mathfrak{B}_{\mathfrak{m},v}}\otimes K)$ is clear by our knowledge on the classical local-global compatibility (Hecke correspondence) at primes above $p$. We give it in our version here.
Lemma 3.3.5.
Let $\mathfrak{p}_{v}=R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}[\frac{1}{p}]\cap\mathfrak{p}$ and $\rho(\mathfrak{p})_{v}:G_{F_{v}}\to\mathrm{GL}_{2}(K)$ be the semi-simple representation given by $\mathfrak{p}_{v}$. Then
  1. (1)
    $\rho(\mathfrak{p})_{v}$ is de Rham of Hodge-Tate weights $(w_{\sigma_{v}},w_{\sigma_{v}}+k_{\sigma_{v}}-1)$, where $\sigma_{v}:F\to K$ is the embedding induced by $v$. More precisely,
    $\operatorname{gr}^{i}(\rho(\mathfrak{p})_{v}\otimes B_{\mathrm{dR}})^{G_{F_{v}}}=0$
    unless $i=w_{\sigma_{v}},w_{\sigma_{v}}+k_{\sigma_{v}}-1$.
  2. (2)
    The semi-simple Weil-Deligne representation $\mathrm{WD}(\varepsilon\otimes\rho(\mathfrak{p})_{v})^{\mathrm{ss}}$ corresponds to $\pi_{v}^{K(\mathfrak{p})}$ under the Hecke correspondence.
Proof.
It is [Pan22, Lemma 3.5.5]. Also see [BLGHT11, Theorem 1.1]. ∎
For the automorphic side, we need to study the object $\Pi(\mathfrak{p})$. Recall that this is the closure of
$W^{*}_{(\vec{k},\vec{w}),K}\otimes(\pi^{K(\mathfrak{p})}_{v})^{\oplus d(\mathfrak{p})}=[(\operatorname{Sym}^{k_{\sigma_{v}-2}}(K^{2})\otimes\det{}^{w_{\sigma_{v}}})^{*}\otimes\pi^{K(\mathfrak{p})}_{v}]^{\oplus d(\mathfrak{p})^{\prime}}$
in $S_{\psi}(U^{p})_{K}$. Here $d(\mathfrak{p})^{\prime}$ is some multiple of $d(\mathfrak{p})$.
Let $\Pi_{v}$ be the universal unitary completion of $(\operatorname{Sym}^{k_{\sigma_{v}-2}}(K^{2})\otimes\det{}^{w_{\sigma_{v}}})^{*}\otimes\pi^{K(\mathfrak{p})}_{v}$ as a $K$-representation of $\mathrm{GL}_{2}(F_{v})$. Note that $\pi^{K(\mathfrak{p})}_{v}$ is an irreducible principal series as the level is $\mathrm{GL}_{2}(O_{F_{v}})$ (hence unramified). Otherwise it is one-dimensional and $\pi_{\mathfrak{p}}$ has to factor through the reduced norm map by the approximation Theorem (for example, see [Gee22, Fact 4.17] or [KST20, Lemma 6.2] in general). Then $\Pi_{v}$ is a topologically irreducible admissible unitary $\mathrm{GL}_{2}(F_{v})$-representation. Here we use [Paš09, Theorem 1.1] and [BB10, Théorème] for the non-ordinary case and [BE10, Proposition 2.2.1] for the ordinary case.
The following lemma describes $\Pi(\mathfrak{p})$ by $\Pi_{v}$. See [Pan22, Lemma 3.5.7] for the proof.
Lemma 3.3.6.
$\Pi(\mathfrak{p})$ is a quotient of $\Pi_{v}^{\oplus d(\mathfrak{p})^{\prime}}$.
Using this lemma, we get an injective map
$\operatorname{Hom}^{\mathrm{cont}}_{K[\mathrm{GL}_{2}(F_{v})]}(\Pi(\mathfrak{p}),\Pi_{\mathfrak{B}_{\mathfrak{m},v}})\hookrightarrow\operatorname{Hom}^{\mathrm{cont}}_{K[\mathrm{GL}_{2}(F_{v})]}(\Pi_{v}^{\oplus d(\mathfrak{p})^{\prime}},\Pi_{\mathfrak{B}_{\mathfrak{m},v}}).$
Let $\Pi_{v}^{0}$ be a $\mathrm{GL}_{2}(F_{v})$-invariant bounded open ball of $\Pi_{v}$ and let $(\Pi_{v}^{0})^{d}=\operatorname{Hom}_{\mathcal{O}}(\Pi_{v}^{0},\mathcal{O})$ be its Schikhof dual. Then we also have
$\operatorname{Hom}^{\mathrm{cont}}_{K[\mathrm{GL}_{2}(F_{v})]}(\Pi_{v}^{\oplus d(\mathfrak{p})^{\prime}},\Pi_{\mathfrak{B}_{\mathfrak{m},v}})\cong K^{\oplus d(\mathfrak{p})^{\prime}}\otimes\operatorname{Hom}_{\mathfrak{C}_{\mathrm{GL}_{2}(F_{v}),\psi|_{F_{v}^{\times}}}(\mathcal{O})}(P_{\mathfrak{B}_{\mathfrak{m},v}},(\Pi_{v}^{0})^{d}).$
Note that $\operatorname{Hom}(P_{\mathfrak{B}_{\mathfrak{m},v}},(\Pi_{v}^{0})^{d})\otimes K$ is a right $E_{\mathfrak{B}_{\mathfrak{m},v}}$-module. Then we just need to prove that the action of the centre $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ of $E_{\mathfrak{B}_{\mathfrak{m},v}}$ on $\operatorname{Hom}(P_{\mathfrak{B}_{\mathfrak{m},v}},(\Pi_{v}^{0})^{d})\otimes K$ also factors through $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}[\frac{1}{p}]/\mathfrak{p}_{v}$. We also view $\mathfrak{p}_{v}$ as a prime ideal of $R^{\mathrm{ps},\psi\varepsilon}_{v}$ (see the isomorphism in the definition of $\tau_{\mathrm{Aut}}$), and it suffices to show that $(\Pi_{v}^{0})^{d}$ appears as a subquotient of $P_{\mathfrak{B}_{\mathfrak{m},v}}/\mathfrak{p}_{v}P_{\mathfrak{B}_{\mathfrak{m},v}}$.
If $\rho(\mathfrak{p})_{v}$ is absolutely irreducible, then using Proposition A.1.8 and Proposition A.2.5, we know that there is only one (up to isomorphism) irreducible Banach representation $\Pi^{\prime}_{v}$ appeared in the subquotient of $\operatorname{Hom}^{\mathrm{cont}}_{\mathcal{O}}(P_{\mathfrak{B}_{\mathfrak{m},v}}/\mathfrak{p}_{v}P_{\mathfrak{B}_{\mathfrak{m},v}},K)$, which is characterized by $\mathbf{V}(\Pi^{\prime}_{v})\cong\rho(\mathfrak{p})_{v}\otimes\varepsilon$. By [CDP14, Theorem 1.3] and Lemma 3.3.5, $\Pi_{v}^{\prime}$ is a unitary completion of
$\operatorname{Sym}^{k_{\sigma_{v}}-2}(K^{2})\otimes\det{}^{-(w_{\sigma_{v}}+k_{\sigma_{v}}-2)}\otimes\pi_{v}^{K(\mathfrak{p})}.$
However, such unitary completion is unique up to isomorphism (see [BB10, Théorème 4.3.1]). We have $\Pi_{v}^{\prime}\cong\Pi_{v}$ and hence $(\Pi_{v}^{0})^{d}\otimes K$ appears as a subquotient of $P_{\mathfrak{B}_{\mathfrak{m},v}}/\mathfrak{p}_{v}P_{\mathfrak{B}_{\mathfrak{m},v}}\otimes K$.
If $\rho(\mathfrak{p})_{v}=\psi_{1}\oplus\psi_{2}$ is reducible, we assume the Hodge-Tate weight of $\psi_{1}$ (resp. $\psi_{2}$) is $w_{\sigma_{v}}$ (resp. $w_{\sigma_{v}}+k_{\sigma_{v}}-1$). Then we know that
$\pi^{K(\mathfrak{p})}_{v}\cong(\mathrm{Ind}^{\mathrm{GL}_{2}(F_{v})}_{B(F_{v})}\psi_{1}\varepsilon^{w_{\sigma_{v}}}|\cdot|^{1-w_{\sigma_{v}}}\otimes\psi_{2}\varepsilon^{w_{\sigma_{v}}+k_{\sigma_{v}}-1}|\cdot|^{-w_{\sigma_{v}}-k_{\sigma_{v}}+1})_{\mathrm{sm}}$
is irreducible by our assumption, and we have $\psi_{1}/\psi_{2}\neq\varepsilon^{\pm 1}$. Using [BE10, Proposition 2.2.1], $\Pi_{v}$ is the unitary parabolic induction $(\mathrm{Ind}^{\mathrm{GL}_{2}(F_{v})}_{B(F_{v})}\psi_{2}\varepsilon\otimes\psi_{1})_{\mathrm{cont}}$. Combining Proposition A.1.8 and Proposition A.2.4, we also conclude that $(\Pi_{v}^{0})^{d}$ appears in the subquotient of $P_{\mathfrak{B}_{\mathfrak{m},v}}/\mathfrak{p}_{v}P_{\mathfrak{B}_{\mathfrak{m},v}}$.
Case 2: $\pi_{\mathfrak{p}}$ factors through the reduced norm $N_{D/F}$.
In this case, $\Pi(\mathfrak{p})$ has the form $\eta\circ\det$ for some continuous character $\eta:F_{v}^{\times}\to\mathcal{O}^{\times}$ and the corresponding pseudo-character of $G_{F_{v}}$ is $\eta+\eta\varepsilon^{-1}$. Our claims follow from Proposition A.1.8 and Proposition A.2.4.
We now finish the proof of the second assertion of this theorem.
For the first part, we recall the injection map given in Lemma 3.3.4. As we have shown that $\Pi(\mathfrak{p})$ belongs to the block $\mathfrak{B}_{\mathfrak{m},v}$, for any block $\mathfrak{B}^{\prime}$ of $\mathrm{GL}_{2}(F_{v})$, we have $\operatorname{Hom}^{\mathrm{cont}}_{K[\mathrm{GL}_{2}(F_{v})]}(S_{\psi}(U^{p})_{\mathfrak{m}}\otimes K,\Pi_{\mathfrak{B}^{\prime}})=0$ unless $\mathfrak{B}^{\prime}=\mathfrak{B}_{\mathfrak{m},v}$. This proves the first claim. ∎
There are some applications of the local-global compatibility result. The first one is the finiteness of the Hecke module $\mathrm{m}$, which contributes to our patching argument.
Corollary 3.3.7.
Keep notations as above.
  1. (1)
    $\mathrm{m}$ is a faithful, finitely generated $\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}$-module.
  2. (2)
    $\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}$ is a finite $R^{\mathrm{ps},\psi\varepsilon^{-1}}_{p}$-algebra.
  3. (3)
    $S_{\psi}(U^{p},\mathbb{F})[\mathfrak{m}]$ is a representation of $D_{p}^{\times}$ of finite length.
Proof.
The faithfulness of the first claim follows from the first part of the Theorem directly. The finiteness assertions in both the first and the second claims follow from Proposition 3.3.2, 1) of Proposition A.1.4 and the previous theorem. The last claim follows from the first one. ∎
The following application describes modular primes of the big Hecke algebra.
Corollary 3.3.8.
Keep notations as above. For a maximal ideal $\mathfrak{p}$ of $\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}[\frac{1}{p}]$, we assume that for any $v|p$,
  • $\rho(\mathfrak{p})|_{G_{F_{v}}}$ is absolutely irreducible and de Rham with distinct Hodge-Tate weights,
where $\rho(\mathfrak{p}):G_{F,\Sigma}\to\mathrm{GL}_{2}(k(\mathfrak{p}))$ is the semi-simple representation associated to $\mathfrak{p}$. Then $\mathfrak{p}$ is a pull-back of a maximal ideal of $\mathbb{T}_{(\vec{k},\vec{w}),\psi}(U^{p}U_{p})[\frac{1}{p}]$ for some weight $(\vec{k},\vec{w})\in\mathbb{Z}_{>1}^{\operatorname{Hom}(F,\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu)}\times\mathbb{Z}^{\operatorname{Hom}(F,\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu)}$ and open compact subgroup $U_{p}\subseteq K_{p}$. By Jacquet-Langlands correspondence, it also arises from a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{2}(\mathbb{A}_{F})$.
Proof.
The proof is nearly the same as [Pan22, Corollary 3.5.10], and the main differences are that we use Theorem 3.3.3 and Paškūnas-Tung’s results (see Appendix A.2) instead. We sketch it here.
Let $\mathfrak{p}_{v}\in\operatorname{Spec}R^{\mathrm{ps},\psi\varepsilon^{-1}}_{v}$ be the pull-back of $\mathfrak{p}$. Using Proposition A.1.8 and Proposition A.2.5, all the irreducible subquotients of $\operatorname{Hom}^{\mathrm{cont}}_{\mathcal{O}}(P_{\mathfrak{B}_{\mathfrak{m},v}}/\mathfrak{p}_{v}P_{\mathfrak{B}_{\mathfrak{m},v}},K)$ are isomorphic, and we denote it by $\Pi_{v}$. Enlarging $K$ if necessary, we may assume $\Pi_{v}$ is absolutely irreducible. We recall the following lemma ([Pan22, Lemma 3.5.11]).
Lemma 3.3.9.
Keep notations as above.
  1. (1)
    The unitary representation $\Pi:=\widehat{\bigotimes}_{v|p}\Pi_{v}$ of $D_{p}^{\times}$ is topologically irreducible.
  2. (2)
    Let $\mathfrak{p}_{p}=\mathfrak{p}\cap R^{\mathrm{ps},\psi\varepsilon^{-1}}_{p}$. Then $E_{\mathfrak{B}_{\mathfrak{m}}}[\frac{1}{p}]/(\mathfrak{p}_{p})$ is a central simple $E$-algebra.
Proof of lemma.
Using Proposition A.2.3, Proposition 3.3.2 and Proposition A.1.5, the conditions of Theorem A.1.6 hold for $G=D_{p}^{\times}$. Therefore, to prove the lemma, we just need to show that $\mathrm{m}(\Pi)$ is a simple $E_{\mathfrak{B}_{\mathfrak{m}}}[\frac{1}{p}]$-module (see Appendix A.1 for the definition of the functor $\mathrm{m}$). Combining Lemma 3.3.1, we have $\mathrm{m}(\Pi)\cong\widehat{\bigotimes}_{v|p}\mathrm{m}(\Pi_{v})$. As $\Pi_{v}$ is absolutely irreducible, using Lemma A.1.1 and Proposition A.1.4, we know that $E_{v}$ is a central simple algebra over $K$ and $\mathrm{m}(\Pi_{v})$ is a simple $E_{v}$-module. Hence $\mathrm{m}(\Pi)$ is a simple module of $\bigotimes E_{v}$. As $\rho(\mathfrak{p}_{v})$ is absolutely irreducible, using Theorem A.1.7 and Proposition A.2.5, we know that $E_{v}$ is in fact $E_{\mathfrak{B}_{\mathfrak{m},v}}[\frac{1}{p}]/(\mathfrak{p}_{v})$, and it implies the lemma. ∎
Now we return to the proof our the corollary.
Let $\tilde{\mathfrak{p}}\in\operatorname{Spec}\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}$ be the pull-back of $\mathfrak{p}$. Note that we have isomorphisms
$\operatorname{Hom}^{\mathrm{cont}}_{\mathcal{O}}(M_{\psi}(U^{p})_{\mathfrak{m}}/\tilde{\mathfrak{p}}M_{\psi}(U^{p})_{\mathfrak{m}},K)\cong S_{\psi}(U^{p})_{\mathfrak{m}}[\tilde{\mathfrak{p}}]\otimes K\cong(S_{\psi}(U^{p})_{\mathfrak{m}}\otimes K)[\mathfrak{p}].$
Using Corollary 3.3.7 and the previous lemma, we obtain an isomorphism $(S_{\psi}(U^{p})_{\mathfrak{m}}\otimes K)[\mathfrak{p}]\cong(\widehat{\bigotimes}_{v|p}\Pi_{v})^{d}$ for some positive integer $d$. By [CDP14, Theorem 1.3], for each $v|p$, $\Pi_{v}$ has non-zero locally algebraic vectors of $\mathrm{GL}_{2}(F_{v})$. Thus $(S_{\psi}(U^{p})_{\mathfrak{m}}\otimes K)[\mathfrak{p}]$ contains non-zero locally algebraic vectors of $D_{p}^{\times}$. Combining the discussion in Remark 3.1.2, we prove the result. ∎
Another application is the lower bound of the dimension of the big Hecke algebra.
Corollary 3.3.10.
Keep notations as above. Each irreducible component of $\mathbb{T}_{\psi}(U^{p})_{\mathfrak{m}}$ is of characteristic zero and of dimension at least $1+2[F:\mathbb{Q}]$.
Proof.
The proof is the same as [Pan22, Theorem 3.6.1]. See [Pan22, Section 3.6] for more details. ∎
Keep the character $\xi$ as in the previous subsection. Then one can obtain similar results for $M_{\psi,\xi}(U^{p})$ and $\mathbb{T}_{\psi,\xi}(U^{p})$, which are to apply Taylor’s Ihara avoidance trick in [Tay08] to prove the patching argument.
4. Pro-modularity
In this section, we give some pro-modularity results.
For notations, we suppose $p$ is an odd prime. We denote by $K$ a $p$-adic local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$.
4.1. The residually reducible case
Let $F$ be a totally real number field of even degree over $\mathbb{Q}$ in which $p$ splits completely. Write $\Sigma_{p}$ as the set of places of $F$ above $p$. Let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$ such that for all $v\in\Sigma\setminus\Sigma_{p}$, we have $p\mid\operatorname{Nm}(v)-1$. Let $\xi_{v}:k(v)^{\times}\to\mathcal{O}^{\times}$ be characters of $p$-power order for $v\in\Sigma\setminus\Sigma_{p}$, and we can view them as characters of $I_{v}$ by local class field theory. Let $\chi:G_{F,\Sigma}\to\mathcal{O}^{\times}$ be a continuous totally odd character such that
1) $\chi$ is unramified outside $\Sigma_{p}$,
2) $\chi(\operatorname{Frob}_{v})\equiv 1~{}\operatorname{mod}~{}\varpi$ for $v\in\Sigma\setminus\Sigma_{p}$.
Let $D$ be a quaternion algebra over $F$ ramified exactly at all infinite places, and we fix an isomorphism between $(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}$ and $\operatorname{GL}_{2}(\mathbb{A}_{F}^{\infty})$. Let $\psi=\chi\varepsilon$ be a character of $(\mathbb{A}_{F}^{\infty})^{\times}/F_{+}^{\times}$ via global class field theory. Let $U^{p}=\prod_{v\nmid p}U_{v}$ be a tame level such that $U_{v}=\operatorname{GL}_{2}(O_{F_{v}})$ if $v\notin\Sigma$ and $U_{v}=\operatorname{Iw}_{v}:=\{g\equiv\begin{pmatrix}*&*\\ 0&*\end{pmatrix}~{}\operatorname{mod}~{}\varpi_{v},g\in\operatorname{GL}_{2}(O_{F_{v}})\}$ otherwise. For any $v\in\Sigma\setminus\Sigma_{p}$, the map $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\to\xi_{v}(\frac{a}{d}~{}\operatorname{mod}~{}\varpi_{v})$ defines a character of $U_{v}$ and the product of $\xi_{v}$ can be viewed as a character $\xi$ of $U^{p}$ by projecting to $\prod_{v\in\Sigma\setminus\Sigma_{p}}U_{v}$. From our setting, we can define a Hecke algebra $\mathbb{T}_{\xi}:=\mathbb{T}_{\psi,\xi}(U^{p})$ as in the previous section. We also make the following assumption:
Assumption 4.1.1.
There exists a maximal ideal $\mathfrak{m}_{\xi}$ of $\mathbb{T}_{\xi}$ generated by $T_{v}-(1+\chi(\operatorname{Frob}_{v})),v\notin\Sigma$ and $\varpi$.
Consider the universal deformation ring $R^{\mathrm{ps}}$ which pro-represents the functor from $\mathfrak{U}_{\mathcal{O}}$ to the category of sets sending $R$ to the set of two-dimensional pseudo-representations $T$ of $G_{F,\Sigma}$ over $R$ such that $T$ is a lifting of $1+\bar{\chi}$ with determinant $\chi$. Let $R^{\mathrm{ps},\{\xi_{v}\}}$ be the natural quotient of $R^{\mathrm{ps}}$ further satisfying $T|_{I_{F_{v}}}=\xi_{v}+\xi_{v}^{-1}$ for any $v\in\Sigma\setminus\Sigma_{p}$.
Remark 4.1.2.
Using [Zha24, Proposition 3.1.5], for any minimal prime $\mathfrak{q}$ of $R^{\mathrm{ps}}$, the natural surjection $R^{\mathrm{ps}}\twoheadrightarrow R^{\mathrm{ps}}/\mathfrak{q}$ factors through $R^{\mathrm{ps},\{\xi_{v}\}}$ for some character $\xi$.
Recall that there is a natural pseudo-deformation $T_{\mathfrak{m}_{\xi}}:G_{F,\Sigma}\to(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ with determinant $\chi$ sending $\operatorname{Frob}_{v}$ to the Hecke operator $T_{v}$ for $v\notin\Sigma$. By the local-global compatibility at $v\in\Sigma\setminus\Sigma_{p}$, we get a natural surjection $R^{\textnormal{ps},\{\xi_{v}\}}\twoheadrightarrow(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$.
For any $\mathfrak{p}\in\operatorname{Spec}(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$, we can define a two-dimensional semi-simple representation $\rho(\mathfrak{p}):G_{F,\Sigma}\to\operatorname{GL}_{2}(k(\mathfrak{p}))$ with trace $T_{\mathfrak{m}_{\xi}}~{}\operatorname{mod}~{}\mathfrak{p}$ in the sense of Definition 2.2.3.
Definition 4.1.3.
(1) We say that a prime of $R^{\textnormal{ps},\{\xi_{v}\}}$ is pro-modular if it comes from a prime of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$. We say that a prime of $R^{\textnormal{ps}}$ is pro-modular if it comes from a pro-modular prime of some $R^{\textnormal{ps},\{\xi_{v}\}}$.
(2) Let $\mathfrak{q}$ be a prime of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ and $A$ be the normal closure of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}/\mathfrak{q}$ in $k(\mathfrak{q})$. We say that $\mathfrak{q}$ is a nice prime if $\mathfrak{q}$ contains $p$ and $\operatorname{dim}(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}/\mathfrak{q}=1$ and there exists a two-dimensional representation $\rho(\mathfrak{q})^{o}:G_{F,\Sigma}\to\operatorname{GL}_{2}(A)$ satisfying:
1) $\rho(\mathfrak{q})^{o}\otimes k(\mathfrak{q})\cong\rho(\mathfrak{q})$ is irreducible.
2) The $\operatorname{mod}~{}\mathfrak{m}_{A}$ reduction $\bar{\rho}_{b}$ of $\rho(\mathfrak{q})^{o}$ is a non-split extension and has the form $\bar{\rho}_{b}(g)=\begin{pmatrix}*&*\\ 0&*\end{pmatrix}$, $g\in G_{F,\Sigma}$. Here $\mathfrak{m}_{A}$ is the maximal ideal of $A$.
3) If $\rho(\mathfrak{q})$ is dihedral, namely isomorphic to $\operatorname{Ind}_{G_{L}}^{G_{F}}\theta$ for some quadratic extension $L$ of $F$ and continuous character $\theta:G_{L}\to k(\mathfrak{q})^{\times}$, then $L\cap F(\mu_{p})=F$, where $\mu_{p}$ is a primitive $p$-th root of unity.
4) $\rho(\mathfrak{q})^{o}\mid_{G_{F_{v}}}=\bar{\rho}_{b}\mid_{G_{F_{v}}}$ for any $v\in\Sigma\setminus\Sigma_{p}$, i.e. $\rho(\mathfrak{q})^{o}|_{G_{F_{v}}}$ is the trivial lifting.
(3) We say that a prime of $R^{\textnormal{ps},\{\xi_{v}\}}$ is nice if it comes from a nice prime of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$.
Now we can state our $R=\mathbb{T}$ theorem.
Theorem 4.1.4.
Keep notations and assumptions as above. Let $\mathfrak{q}\in\operatorname{Spec}(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ be a nice prime and $\mathfrak{q}^{\mathrm{ps}}=\mathfrak{q}\cap R^{\mathrm{ps},\{\xi_{v}\}}$. Then the natural surjective map
$(R^{\mathrm{ps},\{\xi_{v}\}})_{\mathfrak{q}^{\mathrm{ps}}}\to(\mathbb{T}_{\xi})_{\mathfrak{q}}$
has nilpotent kernel.
Remark 4.1.5.
Compared to [Pan22, Theorem 4.1.4], our theorem removes the condition $\bar{\chi}|_{G_{F_{v}}}\neq\omega^{\pm 1}$ for any $v|p$ if $p=3$. The proof is a little bit complicated and technical but nearly the same as Pan’s since the local-global compatibility results also hold in this case based on the recent work [PT21]. We will not give all the details of the proof and just outline the main steps and point out the differences between Pan’s proof and ours later.
Theorem 4.1.6.
Keep notations and assumptions as in the previous theorem. Let $\mathfrak{p}$ be a maximal ideal of $R^{\mathrm{ps},\{\xi_{v}\}}[\frac{1}{p}]$. Assume that
  • For any $v|p$, $\rho(\mathfrak{p})|_{G_{F_{v}}}$ is irreducible and de Rham with distinct Hodge-Tate weights.
  • There exists an irreducible component of $R^{\mathrm{ps},\{\xi_{v}\}}$ containing both $\mathfrak{p}$ and a nice prime $\mathfrak{q}$.
Then $\rho(\mathfrak{p})$ arises from a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{2}(\mathbb{A}_{F})$.
Proof.
This is clear from the previous theorem and Corollary 3.3.8. ∎
Using these two theorems above, to prove the Fontaine-Mazur conjecture in our case, we just need to find enough nice primes in $\operatorname{Spec}R^{\mathrm{ps},\{\xi_{v}\}}$.
The following lemma explains the third condition of the definition of a nice prime.
Lemma 4.1.7.
Let $\mathfrak{q}$ be a prime ideal of $\operatorname{Spec}R^{\mathrm{ps},\{\xi_{v}\}}$ containing $p$ such that $R^{\mathrm{ps},\{\xi_{v}\}}/\mathfrak{q}$ is one-dimensional. Suppose $\rho(\mathfrak{q})$ is irreducible. Then the third condition in (2) of Definition 4.1.3 holds for $\mathfrak{q}$ if one of the following conditions holds:
  1. (1)
    $\bar{\chi}$ is not quadratic.
  2. (2)
    $\bar{\chi}|_{G_{F_{v}}}=\mathbf{1}$ for some $v|p$.
  3. (3)
    There exists a place $v|p$ such that $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1}$ and $\rho(\mathfrak{q})^{o}|_{G_{F_{v}}}\cong\begin{pmatrix}\chi_{v,1}&*\\ 0&\chi_{v,2}\end{pmatrix}$ is reducible. Moreover $\chi_{v,1}/\chi_{v,2}$ is of infinite order, which is equivalent with saying $\chi_{v,1}$ is of infinite order as $\chi_{v,1}\chi_{v,2}=\bar{\chi}$ is of finite order.
Proof.
This is [Pan22, Lemma 4.1.6]. ∎
For the rest of this subsection, we outline the main steps of the proof of Theorem 4.1.4. Here we follow most notations as in [Pan22, Section 4], and one can find more details there.
Enlarging $\mathcal{O}$ if necessary, we may assume $A\cong\mathbb{F}[[T]]$. Let $\mathfrak{q}$ be a nice prime of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$. Let $B$ be the topological closure of the $\mathbb{F}$-algebra generated by all the entries of $\rho(\mathfrak{q})^{o}(G_{F,\Sigma})$ as a subring of $A$. By Chebotarev’s density Theorem, we may find a finite set of primes $T^{\prime}$ disjoint with $\Sigma$ such that the entries of $\rho(\mathfrak{q})^{o}(\operatorname{Frob}_{v}),v\in T^{\prime}$, topologically generate $B$. Let $P=T^{\prime}\cup\Sigma$.
Step 1. Describe universal deformation rings.
Write
$R_{\textnormal{loc}}^{\{\xi_{v}\}}:=(\widehat{\bigotimes}_{v\in\Sigma_{p}}R^{\square}_{v})\widehat{\otimes}_{\mathcal{O}}(\widehat{\bigotimes}_{v\in\Sigma\setminus\Sigma_{p}}R^{\square,\xi_{v}}_{v})\widehat{\otimes}_{\mathcal{O}}(\widehat{\bigotimes}_{v\in T^{\prime}}R^{\square,ur}_{v})$
for the completed tensor product of local framed deformation rings determined by primes in $P$, where the local deformation problem (in the sense of [CHT08, Definition 2.2.2]) for $v\in\Sigma_{p}$ (resp. $v\in\Sigma\setminus\Sigma_{p}$, $v\in T^{\prime}$) is the unrestricted (resp. level raising, unramified) one with determinant $\chi$. By the universal property, the Galois representation $\rho(\mathfrak{q})^{o}$ gives rise to a one-dimensional prime $\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}}$ of $R^{\{\xi_{v}\}}_{\mathrm{loc}}$.
Proposition 4.1.8.
[Pan22, Proposition 4.2.3] The $\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}}$-adic completion $\widehat{(R^{\{\xi_{v}\}}_{\mathrm{loc}})_{\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}}}}$ of $(R^{\{\xi_{v}\}}_{\mathrm{loc}})_{\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}}}$ is equidimensional of dimension $3[F:\mathbb{Q}]+3|P|$. The generic point of each irreducible component has characteristic zero. Moreover,
  1. (1)
    If all $\xi_{v}$ are non-trivial, then $\widehat{(R^{\{\xi_{v}\}}_{\mathrm{loc}})_{\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}}}}$ is integral.
  2. (2)
    In general, each minimal prime of $\widehat{(R^{\{\xi_{v}\}}_{\mathrm{loc}})_{\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}}}}/(\varpi)$ contains a unique minimal prime of $\widehat{(R^{\{\xi_{v}\}}_{\mathrm{loc}})_{\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}}}}$.
Remark 4.1.9.
1) Here we need to use the last condition of the definition of a nice prime.
2) To prove this proposition, we first need to know some properties of the local framed deformation rings. More precisely, Pan shows that $R_{v}^{\square}$ is a normal domain of dimension $7$ for any $v|p$ in [Pan22, Lemma 4.2.7]. However, Pan uses [Paš13, Corollary B.5] and [HT15, Corollary 3.6 & 3.7] to depict the versal deformation ring in the case $(\bar{\rho}_{b}|_{G_{F_{v}}})^{\operatorname{ss}}\cong\eta\oplus\eta\omega$, and both results require the assumption $p>3$. To cover the case $p=3$, we just need to use Proposition 2.1.1 instead.
Step 2. Find Taylor-Wiles primes.
For a finite set $Q$ of finite places of $F$, we define $R^{\square_{P},\{\xi_{v}\}}_{\bar{\rho}_{b},Q}$ to be the universal object pro-representing the functor $\mathrm{Def}^{\square_{M},\{\xi_{v}\}}_{\bar{\rho}_{b},Q}$ from $\mathfrak{U}_{\mathcal{O}}$ to the category of sets sending $R$ to the set of tuples $(\rho_{R};\alpha_{v})_{v\in M}$ modulo the equivalence relation $\sim_{P}$ where
  • $\rho_{R}:G_{F,\Sigma\cup Q}\to\mathrm{GL}_{2}(R)$ is a lifting of $\bar{\rho}_{b}$ to $R$ with determinant $\chi$ such that $\mathrm{tr}(\rho_{R})|_{I_{F_{v}}}=\xi_{v}+\xi_{v}^{-1}$ for any $v\in\Sigma\setminus\Sigma_{p}$.
  • $\alpha_{v}\in 1+M_{2}(\mathfrak{m}_{R}),v\in P$. Here $\mathfrak{m}_{R}$ is the maximal ideal of $R$.
  • $(\rho_{R};\alpha_{v})_{v\in P}\sim_{P}(\rho^{\prime}_{R};\alpha^{\prime}_{v})_{v\in P}$ if there exists an element $\beta\in 1+M_{2}(\mathfrak{m}_{R})$ with $\rho^{\prime}_{R}=\beta\rho_{R}\beta^{-1},\alpha^{\prime}_{v}=\beta\alpha_{v}$ for any $v\in P$.
By the universal property, there is a natural map $R^{\{\xi_{v}\}}_{\mathrm{loc}}\to R^{\square_{P},\{\xi_{v}\}}_{\bar{\rho}_{b},Q}$.
Proposition 4.1.10.
[Pan22, Proposition 4.3.1] Let $r=\dim_{k(\mathfrak{q})}H^{1}(G_{F,P},\operatorname{ad}^{0}\rho(\mathfrak{q})(1))$. Then there exists an integer $C$ such that for any positive integer $N$, we can find a finite set of primes $Q_{N}$ disjoint with $P$ such that
  1. (1)
    $|Q_{N}|=r$.
  2. (2)
    $N(v)\equiv 1\mod p^{N}$ for any $v\in Q_{N}$.
  3. (3)
    $\rho(\mathfrak{q})(\operatorname{Frob}_{v})$ has distinct eigenvalues $\alpha_{v},\beta_{v}$ with $\ell(A/(\alpha_{v}-\beta_{v})^{2})<C$ for any $v\in Q_{N}$.
  4. (4)
    There exists an $A$-module $M_{N}$ with $\ell(M_{N})<C$ such that
    $\mathfrak{q}_{b,Q_{N}}/(\mathfrak{q}_{b,Q_{N}}^{2},\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}})\otimes_{B}A\cong A^{\oplus g}\oplus M_{N}$
    as $A$-modules, where $g=r+|P|-[F:\mathbb{Q}]-1$.
  5. (5)
    There exists a map $R_{\infty}^{\{\xi_{v}\}}:=R^{\{\xi_{v}\}}_{\mathrm{loc}}[[x_{1},\cdots,x_{g}]]\to R^{\square_{P},\{\xi_{v}\}}_{\bar{\rho}_{b},Q_{N}}$ such that the images of $x_{i}$ are in $\mathfrak{q}_{b,Q_{N}}$ and $\mathfrak{q}_{b,Q_{N}}/(\mathfrak{q}_{b,Q_{N}}^{2},\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}},x_{1},\cdots,x_{g})$ is killed by some element $f\in B$ independent of $N$ with $\ell(A/(f))<C$.
Remark 4.1.11.
1) The proof of the proposition needs to use the third condition of the definition of a nice prime.
2) With Taylor-Wiles primes, we can define the completed homology with auxiliary levels. See [Pan22, Section 4.4] for precise definitions.
Step 3. Modify Hecke modules.
Define:
  • $\Delta_{\infty}:=\mathbb{Z}_{p}^{\oplus r}$, where $r$ is defined in Proposition 4.1.10.
  • $\mathcal{O}_{\infty}=\mathcal{O}[[y_{1},\cdots,y_{4|P|-1}]]$ with maximal ideal $\mathfrak{b}$ and prime ideal $\mathfrak{b}_{1}=(y_{1},\cdots,y_{4|P|-1})$.
  • $S_{\infty}=\mathcal{O}_{\infty}[[\Delta_{\infty}]]\cong\mathcal{O}_{\infty}[[s_{1},\cdots,s_{r}]]$. This is a local $\mathcal{O}_{\infty}$-algebra with maximal ideal $\mathfrak{a}$. Let $\mathfrak{a}_{0}=\ker(S_{\infty}\to\mathcal{O}_{\infty})=(s_{1},\cdots,s_{r})$ be the augmentation ideal, and $\mathfrak{a}_{1}=(\mathfrak{a}_{0},\mathfrak{b}_{1})=(y_{1},\cdots,y_{4|P|-1},s_{1},\cdots,s_{r})$.
  • $S_{\infty}^{\prime}\subseteq\mathcal{O}_{\infty}[[\Delta_{\infty}]]$ is the closure (under the profinite topology) of the $\mathcal{O}_{\infty}$-subalgebra generated by all elements of the form $g+g^{-1}$ with $g=(0,\cdots,0,a,0,\cdots,0)\in\Delta_{\infty}$ for some $a\in\mathbb{Z}_{p}$. This is a regular local $\mathcal{O}_{\infty}$-algebra and $S_{\infty}$ is a finite free $S_{\infty}^{\prime}$-algebra. Write $\mathfrak{a}_{0}^{\prime}=\mathfrak{a}_{0}\cap S_{\infty}^{\prime},\mathfrak{a}_{1}^{\prime}=\mathfrak{a}_{1}\cap S_{\infty}^{\prime}$. We may find $r$ elements $s_{1}^{\prime},\cdots,s_{r}^{\prime}$ that generate $\mathfrak{a}_{0}^{\prime}$ and $S_{\infty}^{\prime}\cong\mathcal{O}_{\infty}[[s_{1}^{\prime},\cdots,s_{r}^{\prime}]]$.
  • $\mathrm{m}^{\{\xi_{v}\}}:=\operatorname{Hom}_{\mathfrak{C}_{D_{p}^{\times},\psi}(\mathcal{O})}(P_{\mathfrak{B}_{\mathfrak{m}}},M_{\psi,\xi}(U^{p})_{\mathfrak{m}})$, and $\mathrm{m}_{0}^{\{\xi_{v}\}}:=$ the $\mathfrak{q}$-adic completion of $(\mathrm{m}^{\{\xi_{v}\}})_{\mathfrak{q}}$ as a $\mathbb{T}_{\psi,\xi}(U^{p})_{\mathfrak{q}}$-module.
In general, the (patched) completed homology is not finitely generated module over the big Hecke algebra as there is an action of $\mathrm{GL}_{2}(\mathbb{Q}_{p})$. Thus, it is necessary to modify it (using [Pan22, Proposition 4.4.4]). After taking the inverse limit, we denote the modified Hecke module by $\mathrm{m}_{\infty}^{\{\xi_{v}\}}$ (see [Pan22, Definition 4.7.4] for the precise definition).
Proposition 4.1.12.
[Pan22, Proposition 4.7.5] Let $\mathfrak{q}^{\{\xi_{v}\}}_{\infty}$ be the prime ideal of $R_{\infty}^{\{\xi_{v}\}}$ generated by $\mathfrak{q}^{\{\xi_{v}\}}_{\mathrm{loc}},x_{1},\cdots,x_{g}$.
  1. (1)
    $\mathrm{m}^{\{\xi_{v}\}}_{\infty}$ is a finitely generated $\widehat{(R^{\{\xi_{v}\}}_{\infty})_{\mathfrak{q}^{\{\xi_{v}\}}_{\infty}}}$-module.
  2. (2)
    $\mathrm{m}^{\{\xi_{v}\}}_{\infty}$ is a flat $S_{\infty}$-module.
  3. (3)
    We have $\mathrm{m}^{\{\xi_{v}\}}_{\infty}/\mathfrak{a}_{1}\mathrm{m}^{\{\xi_{v}\}}_{\infty}\cong(\mathrm{m}_{0}^{\{\xi_{v}\}})^{\oplus 2^{r}}.$
Remark 4.1.13.
To prove this proposition, we just need to replace [Pan22, Theorem 3.5.3] and [Pan22, Corollary 3.5.8] by Theorem 3.3.3 and Corollary 3.3.7 in Pan’s proof.
Step 4. Use Taylor’s trick.
The strategy of the proof of Theorem 4.1.4 follows from [Tay08, Theorem 4.1].
At first, using Corollary 3.3.7, we know that $\textnormal{m}_{0}^{\{\xi_{v}\}}$ is a finitely generated faithful $\widehat{(\mathbb{T}_{\xi})_{\mathfrak{q}}}$-module. Then by Corollary 3.3.10, we have $\operatorname{dim}_{\widehat{(\mathbb{T}_{\xi})_{\mathfrak{q}}}}(\textnormal{m}_{0}^{\{\xi_{v}\}})=\operatorname{dim}\widehat{(\mathbb{T}_{\xi})_{\mathfrak{q}}}\geq 2[F:\mathbb{Q}]$ since $\mathfrak{q}$ is of dimension $1$.
Let $(R^{\{\xi_{v}\}})^{\prime}$ be $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}\otimes_{S_{\infty}^{\prime}}S_{\infty}$. We know that $\widehat{(\mathbb{T}_{\xi})_{\mathfrak{q}}}$ is a natural quotient of $(R^{\{\xi_{v}\}})^{\prime}$ by mapping $S_{\infty}$ to $S_{\infty}/\mathfrak{a}_{1}=\mathcal{O}$. Note that $y_{1},...,y_{4|P|-1},s_{1},...,s_{r}\in\mathfrak{a}_{1}$ form a regular sequence of $\textnormal{m}_{\infty}^{\{\xi_{v}\}}$. Then we can conclude that
$\operatorname{dim}_{(R^{\{\xi_{v}\}})^{\prime}}(\textnormal{m}_{\infty}^{\{\xi_{v}\}})\geq 4|P|-1+r+\operatorname{dim}\widehat{(\mathbb{T}_{\xi})_{\mathfrak{q}}}\geq 4|P|-1+r+2[F:\mathbb{Q}].$
Using Proposition 4.1.8, we can conclude that $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$ is equidimensional of dimension $4|P|-1+r+2[F:\mathbb{Q}]$ (also see [Pan22, Section 4.2, Lemma 4.8.2]). As $(R^{\{\xi_{v}\}})^{\prime}$ is finite free over $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$, we have
$\operatorname{dim}_{(R^{\{\xi_{v}\}})^{\prime}}(\textnormal{m}_{\infty}^{\{\xi_{v}\}})=\operatorname{dim}_{\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}}(\textnormal{m}_{\infty}^{\{\xi_{v}\}})=4|P|-1+r+2[F:\mathbb{Q}]=\operatorname{dim}\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}.$
If the characters in $\{\xi_{v}\}$ are all non-trivial, using Proposition 4.1.8 again, $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$ is irreducible. From the previous equations, we know that $\textnormal{m}_{\infty}^{\{\xi_{v}\}}$ has full support on $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$. By [Pan22, Lemma 4.8.4], Theorem 4.1.4 holds in this case.
In general, we suppose that $\xi^{\prime}_{v}:k(v)\to\mathcal{O}^{\times}$ are non-trivial characters of $p$-power order for $v\notin\Sigma\setminus\Sigma_{p}$. We have natural isomorphisms $\widehat{(R^{\{\xi^{\prime}_{v}\}}_{\infty})_{\mathfrak{q}^{\{\xi^{\prime}_{v}\}}_{\infty}}}/(\varpi)\cong\widehat{(R^{\{\xi_{v}\}}_{\infty})_{\mathfrak{q}^{\{\xi_{v}\}}_{\infty}}}/(\varpi)$ and $\mathrm{m}^{\{\xi^{\prime}_{v}\}}_{\infty}/\varpi\mathrm{m}^{\{\xi^{\prime}_{v}\}}_{\infty}\cong\mathrm{m}^{\{\xi_{v}\}}_{\infty}/\varpi\mathrm{m}^{\{\xi_{v}\}}_{\infty}$. Then Theorem 4.1.4 follows from Proposition 4.1.8 and the previous case.
4.2. The ordinary case
In this subsection, we recall the modularity result in the ordinary case proved in [Pan22, Section 5], which will be devoted to finding enough pro-modular primes in the non-ordinary case.
Let $F$ be an abelian totally real field in which $p$ is unramified. Let $\Sigma$ be a finite set of finite places containing all places above $p$.
Definition 4.2.1.
Let $T:G_{F}\to R$ be a two-dimensional pseudo-representation over some ring $R$ such that for some place $v$, $T|_{G_{F_{v}}}=\psi_{1}+\psi_{2}$ is a sum of two characters. We say that $T$ is $\psi_{1}$-ordinary if for any $\sigma,\tau\in G_{F_{v}},\eta\in G_{F}$,
$T(\sigma\tau\eta)-\psi_{1}(\sigma)T(\tau\eta)-\psi_{2}(\tau)T(\sigma\eta)+\psi_{1}(\sigma)\psi_{2}(\tau)T(\eta)=0.$
Lemma 4.2.2.
Let $\rho:G_{F}\to\mathrm{GL}_{2}(R)$ be a two-dimensional irreducible representation over some ring $R$ with trace $T$ such that $T|_{G_{F_{v}}}=\psi_{1}+\psi_{2}$, a sum of two characters. Suppose $\rho|_{G_{F_{v}}}$ has the form $\begin{pmatrix}\psi_{1}&*\\ 0&\psi_{2}\end{pmatrix}$. Then $T$ is $\psi_{1}$-ordinary. Conversely, if $R$ is a field and $T$ is $\psi_{1}$-ordinary, then after possibly enlarging $R$,
$\rho|_{G_{F_{v}}}\cong\begin{pmatrix}\psi_{1}&*\\ 0&\psi_{2}\end{pmatrix}.$
Proof.
See [Pan22, Lemma 5.3.2]. ∎
Let $\chi:G_{F,\Sigma}\to\mathcal{O}^{\times}$ be a continuous totally odd character such that
  • $\bar{\chi}$, the reduction of $\chi$ modulo $\varpi$, can be extended to a character of $G_{\mathbb{Q}}$.
  • $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1}$ for any $v|p$.
  • $\chi|_{G_{F_{v}}}$ is de Rham for any $v|p$.
Let $R^{\mathrm{ps},\mathrm{ord}}$ be the universal pseudo-deformation ring which pro-represents the functor from $\mathfrak{U}_{\mathcal{O}}$ to the category of sets sending $R$ to the set of two-dimensional pseudo-representations $T$ of $G_{F,\Sigma}$ over $R$ such that $T$ is a lifting of $1+\bar{\chi}$ with determinant $\chi$ and $T|_{G_{F_{v}}}$ is reducible for any $v|p$, i.e. $y(\sigma,\tau)=0$ for any $\sigma,\tau\in G_{F_{v}},v|p$. Clearly, there is a natural surjection $R^{\mathrm{ps}}\twoheadrightarrow R^{\mathrm{ps},\mathrm{ord}}$. Denote the universal pseudo-representation by $T^{univ}:G_{F,\Sigma}\to R^{\mathrm{ps},\mathrm{ord}}$.
As $\bar{\chi}|_{G_{F_{v}}}$ is non-trivial, we may write $T^{univ}|_{G_{F_{v}}}=\psi^{univ}_{v,1}+\psi^{univ}_{v,2}$ for some characters $\psi^{univ}_{v,1},\psi^{univ}_{v,2}:G_{F_{v}}\to(R^{\mathrm{ps},\mathrm{ord}})^{\times}$ which are liftings of $\mathbf{1},\bar{\chi}|_{G_{F_{v}}}$, respectively. By local class field theory, $\psi^{univ}_{v,1}|_{I_{F_{v}}}$ induces a homomorphism $\mathcal{O}[[O_{F_{v}}^{\times}(p)]]\to R^{\mathrm{ps},\mathrm{ord}}$ for any $v|p$. Here $O_{F_{v}}^{\times}(p)$ denotes the $p$-adic completion of $O_{F_{v}}^{\times}$. Taking the completed tensor product over $\mathcal{O}$ for all $v|p$, we get a map from the Iwasawa algebra to the universal ordinary pseudo-deformation ring:
$\Lambda_{F}:=\widehat{\bigotimes}_{v|p}\mathcal{O}[[O_{F_{v}}^{\times}(p)]]\to R^{\mathrm{ps},\mathrm{ord}}.$
Theorem 4.2.3.
[Pan22, Theorem 5.1.1] Under the assumptions for $F,\chi$ as above, we have
  1. (1)
    $R^{\mathrm{ps},\mathrm{ord}}$ is a finite $\Lambda_{F}$-algebra.
  2. (2)
    For any maximal ideal $\mathfrak{p}$ of $R^{\mathrm{ps},\mathrm{ord}}[\frac{1}{p}]$, we denote the associated semi-simple representation $G_{F,\Sigma}\to\mathrm{GL}_{2}(k(\mathfrak{p}))$ by $\rho(\mathfrak{p})$ in the sense of Definition 2.2.3. Assume that
    • $\rho(\mathfrak{p})$ is irreducible.
    • For any $v|p$, $\rho(\mathfrak{p})|_{G_{F_{v}}}\cong\begin{pmatrix}\psi_{v,1}&*\\ 0&\psi_{v,2}\end{pmatrix}$ such that $\psi_{v,1}$ is de Rham and has strictly less Hodge-Tate number than $\psi_{v,2}$ for any embedding $F_{v}\hookrightarrow\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}_{p}\mkern-1.5mu}\mkern 1.5mu$.
    Then $\rho(\mathfrak{p})$ comes from a twist of a Hilbert modular form.
Definition 4.2.4.
Let $\mathfrak{q}$ be a prime ideal of $R^{\mathrm{ps},\mathrm{ord}}$. If $\mathfrak{q}$ is an inverse image of a prime $\mathfrak{p}$ of $R^{\mathrm{ps},\mathrm{ord}}[\frac{1}{p}]$ satisfying the assumptions in (2) of the previous theorem via the natural map $R^{\mathrm{ps},\mathrm{ord}}\to R^{\mathrm{ps},\mathrm{ord}}[\frac{1}{p}]$, then we call it a regular de Rham prime.
The following lemma is an essential application of Pan’s finiteness result.
Lemma 4.2.5.
Suppose that $p$ splits completely in $F$. Let $C$ be an irreducible component of $\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}$, and here we also view $C$ as a subset of $\operatorname{Spec}R^{\mathrm{ps}}$. Assume that $C$ is of dimension $[F:\mathbb{Q}]+1$.
1) The subset $C^{\textnormal{reg}}$ consisting of regular de Rham primes of $C$ is dense in $C$. In other words, every prime in $C$ is pro-modular, which implies that Assumption 4.1.1 holds.
2) If further $[F:\mathbb{Q}]\geq 7|\Sigma\setminus\Sigma_{p}|+2$, then there exists a nice prime in $C$.
Proof.
Under our assumptions, the Iwasawa algebra $\Lambda_{F}$ is a power series ring over $\mathcal{O}$ of relative dimension $[F:\mathbb{Q}]$. Let $\mathfrak{q}$ be the minimal prime of $R^{\mathrm{ps},\mathrm{ord}}$ corresponding to $C$. Combining Theorem 4.2.3, the map $\Lambda_{F}\to R^{\mathrm{ps},\mathrm{ord}}/\mathfrak{q}$ is a finite injection, and hence the morphism $C\to\operatorname{Spec}\Lambda_{F}$ is a finite surjection.
For the first claim, using Lemma 4.2.2, we know that for any $v|p$, there exists $n_{v}\in\{1,2\}$ such that for any $\mathfrak{p}\in C$,
$\rho(\mathfrak{p})|_{G_{F_{v}}}\cong\begin{pmatrix}\psi^{univ}_{v,n_{v}}\operatorname{mod}\mathfrak{p}&*\\ 0&\psi^{univ}_{v,3-n_{v}}\operatorname{mod}\mathfrak{p}\end{pmatrix}.$
Combining this and the surjectivity of the map $C\to\operatorname{Spec}\Lambda_{F}$, we know that the image of $C^{\textnormal{reg}}$ in $\operatorname{Spec}\Lambda_{F}$ is dense. Thus, $C^{\textnormal{reg}}$ is dense in $C$.
For the second claim, the proof is the same as [Zha24, Lemma 4.4.4]. We briefly sketch it here.
For any $v\in\Sigma\setminus\Sigma_{p}$, we fix a lifting $\sigma_{v}$ of the generator of the $\mathbb{Z}_{p}$-quotient of $I_{v}$ and a lifting $\operatorname{Frob}_{v}$ of the Frobenius element. Write $S=\{\sigma_{v},\operatorname{Frob}_{v}:~{}v\in\Sigma\setminus\Sigma_{p}\}$. Clearly, $|S|=2|\Sigma\setminus\Sigma_{p}|$.
Following notations in Definition 2.2.1, we let $I$ be an ideal of $R^{\textnormal{ps,ord}}/\mathfrak{q}$ generated by
$\{\varpi,a(\sigma_{v})-1,a(\operatorname{Frob}_{v})-1,d(\operatorname{Frob}_{v})-1,~{}v\ \in\Sigma\setminus\Sigma_{p}\}.$
We consider a minimal prime $\mathfrak{u}$ of $I$. By Krull’s principal ideal theorem, $(R^{\textnormal{ps,ord}}/\mathfrak{q})/\mathfrak{u}$ is of dimension at least $[F:\mathbb{Q}]-3|\Sigma\setminus\Sigma_{p}|$. For simplicity, we write such a domain as $R_{0}$.
Now we consider the following steps.
Step 1. We start from the CNL domain $R_{0}$.
By Proposition 2.2.4, we can find a partition of $S=S_{1}\amalg S_{2}$, a positive integer $n>1$ prime to $p$ and a CNL domain $R_{0}^{\prime}$ satisfying the following conditions.
  1. (1)
    $R_{0}^{\prime}$ is a quotient of $R_{0}$.
  2. (2)
    For any $\theta\in S_{1}$, we have $y(\theta,\alpha)=0$ for any $\alpha\in G_{F,\Sigma}$ in $R_{0}^{\prime}$. For any $\theta^{\prime},\theta^{\prime\prime}\in S_{2}$, we have $y(\theta^{\prime},\alpha)^{n}=y(\theta^{\prime\prime},\alpha)^{n}$ for any $\alpha\in G_{F,\Sigma}$ in $R_{0}^{\prime}$.
  3. (3)
    For any $\alpha\in G_{F,\Sigma},\theta\in S_{2}$, either $y(\theta,\beta)=0$ for any $\beta\in G_{F,\Sigma}$ or $u(\alpha,\theta)$ is well-defined and integral over $R_{0}^{\prime}$.
  4. (4)
    We have $\operatorname{dim}R_{0}^{\prime}\geq\operatorname{dim}R_{0}-|S|$. If $S_{2}$ is not empty, then further $\operatorname{dim}R_{0}^{\prime}\geq\operatorname{dim}R_{0}-|S|+1$.
Step 2. We start from the CNL domain $R_{0}^{\prime}$.
Similar to Step 1, we can find a partition of $S=S_{1}^{\prime}\amalg S_{2}^{\prime}$, a positive integer $n^{\prime}>1$ prime to $p$ and a CNL domain $R_{0}^{\prime\prime}$ satisfying the following conditions.
  1. (1)
    $R_{0}^{\prime\prime}$ is a quotient of $R_{0}^{\prime}$.
  2. (2)
    For any $\theta\in S_{1}^{\prime}$, we have $y(\alpha,\theta)=0$ for any $\alpha\in G_{F,\Sigma}$ in $R_{0}^{\prime\prime}$. For any $\theta^{\prime},\theta^{\prime\prime}\in S_{2}^{\prime}$, we have $y(\alpha,\theta^{\prime})^{n^{\prime}}=y(\alpha,\theta^{\prime\prime})^{n^{\prime}}$ for any $\alpha\in G_{F,\Sigma}$ in $R_{0}^{\prime\prime}$.
  3. (3)
    For any $\alpha\in\operatorname{Gal}(F_{\Sigma}/F),\theta\in S_{2}^{\prime}$, either $y(\beta,\theta)=0$ for any $\beta\in G_{F,\Sigma}$ or $u^{\prime}(\alpha,\theta)$ is well-defined and integral over $R_{0}^{\prime\prime}$.
  4. (4)
    We have $\operatorname{dim}R_{0}^{\prime\prime}\geq\operatorname{dim}R_{0}^{\prime}-|S|$. If $S_{2}^{\prime}$ is not empty, then further $\operatorname{dim}R_{0}^{\prime\prime}\geq\operatorname{dim}R_{0}^{\prime}-|S|+1$.
Step 3. We consider the domain $R_{0}^{\prime\prime}$.
If both $S_{2}$ and $S_{2}^{\prime}$ are not empty, then we fix an element $\theta_{i}\in S_{2}$ and an element $\theta_{j}\in S_{2}^{\prime}$. Consider a prime $\mathfrak{p}_{ij}$ of $R_{0}^{\prime\prime}$ of the ideal $(y(\theta_{i},\theta_{j}))$ and let $R^{\prime\prime}=R_{0}^{\prime\prime}/\mathfrak{p}_{ij}$. By Proposition 2.2.2, we know that either $y(\theta_{i},\alpha)=0$ or $y(\beta,\theta_{j})=0$ for all $\alpha,\beta\in\operatorname{Gal}(F_{\Sigma}/F)$. By Krull’s principal ideal theorem, we have $\operatorname{dim}R^{\prime\prime}\geq\operatorname{dim}R_{0}^{\prime\prime}-1$.
If not, we take $R^{\prime\prime}=R_{0}^{\prime\prime}$
In conclusion, there are three possibilities.
  1. (1)
    $y(\theta,\alpha)=y(\alpha,\theta)=0$ for all $\theta\in S$ and $\alpha\in G_{F,\Sigma}$.
  2. (2)
    There exists an element $\theta_{i}\in S$ such that either $y(\theta,\alpha)=0$ or $y(\theta,\alpha)^{n}=y(\theta_{i},\alpha)^{n}$ for all $\theta\in S$ and $\alpha\in G_{F,\Sigma}$, and $y(\alpha,\theta)=0$ for all $\theta\in S$ and $\alpha\in G_{F,\Sigma}$.
  3. (3)
    There exists an element $\theta_{j}\in S$ such that either $y(\alpha,\theta)=0$ or $y(\alpha,\theta)^{n^{\prime}}=y(\alpha,\theta_{j})^{n^{\prime}}$ for all $\theta\in S$ and $\alpha\in G_{F,\Sigma}$, and $y(\theta,\alpha)=0$ for all $\theta\in S$ and $\alpha\in G_{F,\Sigma}$.
In particular, we have $\operatorname{dim}R^{\prime\prime}\geq[F:\mathbb{Q}]-7|\Sigma\setminus\Sigma_{p}|\geq 2$. Using [Zha24, Corollary 3.2.4], we can choose an irreducible one-dimensional prime $\mathfrak{r}$ of $R^{\prime\prime}$. By the first claim of this lemma, we know that $\mathfrak{r}$ is pro-modular. Now we verify that $\mathfrak{r}$ is actually a nice prime in the sense of Definition 4.1.3.
Using Remark 4.1.2, we view $C$ as a subset of $\operatorname{Spec}R^{\mathrm{ps},\{\xi_{v}\}}$ for some character $\xi$. Let $A$ be the normal closure of $R^{\prime\prime}/\mathfrak{r}$. As $A$ is a DVR, we can construct a continuous irreducible representation $\rho(\mathfrak{r}):G_{F,\Sigma}\to\operatorname{GL}_{2}(A)$ in the following form
$\rho(\mathfrak{r})(\sigma)=\begin{pmatrix}a(\sigma)&\frac{y(\sigma,\tau_{\mathfrak{r}})}{y(\sigma_{\mathfrak{r}},\tau_{\mathfrak{r}})}\\ y(\sigma_{\mathfrak{r}},\sigma)&d(\sigma)\end{pmatrix},~{}y(\sigma_{\mathfrak{r}},\tau_{\mathfrak{r}})\neq 0.$
From our construction, $\rho(\mathfrak{r})$ is ordinary at each $v|p$. Then by Lemma 4.1.7 and Theorem 4.2.3, the third condition holds in each case. Therefore, we only need to check the last condition.
In case (1), it is clear.
In case (2), if $y(\theta_{i},\alpha)=0$ for all $\alpha\in G_{F,\Sigma}$, then it is the same as case a). If not, by Proposition 2.2.2, we know that $y(\theta_{i},\tau_{\mathfrak{r}})\neq 0$ (otherwise $y(\sigma_{\mathfrak{r}},\tau_{\mathfrak{r}})=0$). By the third part in Step 1, $u(\sigma_{\mathfrak{r}},\theta_{i})$ is well-defined and integral over $R^{\prime\prime}/\mathfrak{r}$, and hence a unit in $A$. Then we can take $\rho(\mathfrak{r})^{o}=P\rho(\mathfrak{r})P^{-1}$, where
$P=\begin{pmatrix}u(\sigma_{\mathfrak{r}},\theta_{i})&0\\ 0&1\end{pmatrix}\in\operatorname{GL}_{2}(A).$
From our construction, we know that either $u(\theta,\theta_{i})=0$ or $u(\theta,\theta_{i})^{n}=1$ for any $\theta\in S$. As $(n,p)=1$, by Hensel’s lemma, $u(\theta,\theta_{i})$ is an element of $\mathbb{F}^{\prime}\hookrightarrow\mathbb{F}^{\prime}[[T]]=A$, where $\mathbb{F}^{\prime}$ is the residue field of $A$. Hence, the last condition holds in this case.
In case (3), we consider the continuous irreducible Galois representation $\rho(\mathfrak{r})^{\prime}:G_{F,\Sigma}\to\operatorname{GL}_{2}(A)$ in the following form
$\rho(\mathfrak{r})^{\prime}(\sigma)=\begin{pmatrix}a(\sigma)&y(\sigma,\tau_{\mathfrak{r}^{\prime}})\\ \frac{y(\sigma_{\mathfrak{r}^{\prime}},\sigma)}{y(\sigma_{\mathfrak{r}^{\prime}},\tau_{\mathfrak{r}^{\prime}})}&d(\sigma)\end{pmatrix},~{}y(\sigma_{\mathfrak{r}^{\prime}},\tau_{\mathfrak{r}^{\prime}})\neq 0.$
If $y(\alpha,\theta_{j})=0$ for all $\alpha\in G_{F,\Sigma}$, then we can take
$\rho(\mathfrak{r})^{o}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\rho(\mathfrak{r})^{\prime}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\subset\operatorname{GL}_{2}(A).$
Similar to case (1), we can vefiry that $\mathfrak{r}$ is a nice prime. If not, we can take
$\rho(\mathfrak{r})^{o}=\begin{pmatrix}0&u^{\prime}(\tau_{\mathfrak{r}^{\prime}},\theta_{j})\\ 1&0\end{pmatrix}\rho(\mathfrak{r})^{\prime}\begin{pmatrix}0&1\\ u^{\prime}(\theta_{j},\tau_{\mathfrak{r}^{\prime}})&0\end{pmatrix}\subset\operatorname{GL}_{2}(A).$
Similar to case (2), we can also verify that $\mathfrak{r}$ is a nice prime. ∎
By the previous lemma, to find a nice prime, it is crucial to find irreducible components in $\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}$ of large dimension. For this purpose, we recall a classical result in the rest of this subsection.
Let $\Sigma^{o}$ be a subset of $\Sigma_{p}$. We define $\bar{\psi}_{v,1}:G_{F_{v}}\to\mathbb{F}^{\times}$ to be $\mathbf{1}$ if $v\in\Sigma^{o}$ and $\bar{\chi}|_{G_{F_{v}}}(\neq\mathbf{1})$ if $v\notin\Sigma^{o}$.
Consider the universal deformation ring $R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$ which pro-represents the functor from $\mathfrak{U}_{\mathcal{O}}$ to the category of sets sending $R$ to the set of tuples $(T,\{\psi_{v,1}\}_{v\in\Sigma_{p}})$ where
  • For any $v|p$, $\psi_{v,1}:G_{F_{v}}\to R^{\times}$ is a character which lifts $\bar{\psi}_{v,1}$ and satisfies
    $T|_{G_{F_{v}}}=\psi_{v,1}+\psi_{v,1}^{-1}\chi.$
    Moreover $T$ is $\psi_{v,1}$-ordinary.
From the definition, it is clear that $R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$ is a quotient of $R^{\mathrm{ps},\mathrm{ord}}$.
Write
$H^{1}_{\Sigma^{o}}(F):=\ker(H^{1}(G_{F,\Sigma},\mathbb{F}(\bar{\chi}^{-1}))\stackrel{{\scriptstyle\mathrm{res}}}{{\longrightarrow}}\bigoplus_{v\in\Sigma_{p}\setminus\Sigma^{o}}H^{1}(G_{F_{v}},\mathbb{F}(\bar{\chi}^{-1}))).$
Then it is easy to find that any cohomology class $b=[\phi_{b}]\in H^{1}_{\Sigma^{o}}(F)$ defines a two-dimensional representation
$\bar{\rho}_{b}:G_{F,\Sigma}\to\mathrm{GL}_{2}(\mathbb{F}),~{}g\mapsto\begin{pmatrix}1&\phi_{b}(g)\bar{\chi}(g)\\ 0&\bar{\chi}(g)\end{pmatrix},$
which can be viewed as an extension of $\bar{\chi}$ by $\mathbf{1}$.
For any non-zero $b\in H^{1}_{\Sigma^{o}}(F)$, we will write $R^{\mathrm{ord}}_{b}$ for the universal (unframed) deformation ring which pro-represents the functor from $\mathfrak{U}_{\mathcal{O}}$ to the category of sets sending $R$ to the set of tuples $(\rho_{R},\{\psi_{v,1}\}_{\Sigma_{p}})$ where
  • $\rho_{R}:G_{F,\Sigma}\to\mathrm{GL}_{2}(R)$ is a lifting of $\bar{\rho}_{b}$ with determinant $\chi$.
  • For $v|p$, $\psi_{v,1}:G_{F_{v}}\to R^{\times}$ is a lifting of $\bar{\psi}_{v,1}$ such that $\rho_{R}$ has a (necessarily unique) $G_{F_{v}}$-stable rank one $R$-submodule, which is a direct summand of $\rho_{R}$ as a $R$-module, with $G_{F_{v}}$ acting via $\psi_{R}$.
By the universal property, there is a natural map $R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}\to R^{\mathrm{ord}}_{b}$.
Proposition 4.2.6.
Each irreducible component of $R^{\mathrm{ord}}_{b}$ has dimension at least $[F:\mathbb{Q}]+1$ if $H^{0}(G_{F,\Sigma},\operatorname{ad}^{0}\bar{\rho}_{b}(1))=0$ and $[F:\mathbb{Q}]$ in general.
Proof.
One can see [Gee22, Proposition 3.24] or [Pan22, Lemma 5.4.2]. ∎
Remark 4.2.7.
Let $\mathfrak{q}$ be a one-dimensional prime of $R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$ of characteristic 00 such that the corresponding Galois representation $\rho(\mathfrak{q}):G_{F,\Sigma}\to\mathrm{GL}_{2}(k(\mathfrak{q}))$ is absolutely irreducible, and let $\bar{\rho_{b^{\prime}}}$ be its non-split residual representation, where $b^{\prime}$ is a non-zero element in $H^{1}_{\Sigma^{o}}(F)$. Let $R^{\mathrm{ord}}_{b^{\prime}}$ be the universal ordinary deformation ring defined as above, and $\rho(\mathfrak{q})$ defines a one-dimensional prime $\mathfrak{q}_{b^{\prime}}$ of $R^{\mathrm{ord}}_{b^{\prime}}$. Using a similar argument as Proposition 2.2.5, we have an isomorphism $\widehat{(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}}}\cong\widehat{(R^{\mathrm{ord}}_{b^{\prime}})_{\mathfrak{q}_{b^{\prime}}}}$. Then by the previous proposition, we have $\dim(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}}\geq[F:\mathbb{Q}]-1$. In particular, if $\bar{\chi}\neq\omega^{-1}$, we have $\dim(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}}\geq[F:\mathbb{Q}]$, and then we can use Lemma 4.2.5 to find nice primes in some irreducible component containing $\mathfrak{q}$.
If $p>3$ and $\bar{\chi}=\omega^{\pm 1}$, after some finite twist, we may assume $\bar{\chi}=\omega\neq\omega^{-1}$. Then we can use the argument as in the previous paragragh to find nice primes. However, if $p=3$, the $\mod p$ cyclotomic character is quadratic, and the lower bound of each irreducible component containing $\mathfrak{q}$ is $[F:\mathbb{Q}]$. In this case, we cannot use Lemma 4.2.5, and hence we need some more arguments.
4.3. The residually irreducible case
Let $F$ be a totally real field of even degree in which $p$ splits completely. Write $\Sigma_{p}$ as the set of places of $F$ above $p$. Let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$ such that for all $v\in\Sigma\setminus\Sigma_{p}$, we have $p\mid\operatorname{Nm}(v)-1$.
Let $\bar{\rho}:G_{F}\to\mathrm{GL}_{2}(\mathbb{F})$ be a continuous, odd, absolutely irreducible Galois representation. Note that in this case, the deformation of $\bar{\rho}$ is actually the same as the pseudo-deformation of $\mathrm{tr}\bar{\rho}$. We assume that $\bar{\rho}$ is unramified outside $p$ and for each $v\in\Sigma\setminus\Sigma_{p}$, $\bar{\rho}|_{G_{F_{v}}}$ is trivial. Further assume that $\bar{\rho}$ is modular, i.e. it arises from a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{2}(\mathbb{A}_{F})$. Write $\det\bar{\rho}=\bar{\chi}$, and let $\chi:G_{F,\Sigma}\to\mathcal{O}^{\times}$ be a lifting of $\bar{\chi}$.
By the global class field theory, we may view $\psi=\chi\varepsilon$ as a character of $(\mathbb{A}_{F}^{\infty})^{\times}/F^{\times}_{+}$. Let $D$ be the quaternion algebra over $F$ which is ramified exactly at all infinite places. Fix an isomorphism between $(D\otimes_{F}\mathbb{A}_{F}^{\infty})^{\times}$ and $\mathrm{GL}_{2}(\mathbb{A}_{F}^{\infty})$. We define a tame level $U^{p}=\prod_{v\nmid p}U_{v}$ as follows: $U_{v}=\mathrm{GL}_{2}(O_{F_{v}})$ if $v\notin\Sigma$ and
$U_{v}=\mathrm{Iw}_{v}:=\{g\in\mathrm{GL}_{2}(O_{F_{v}}),g\equiv\begin{pmatrix}*&*\\ 0&*\end{pmatrix}\mod\varpi_{v}\}$
otherwise. For any $v\in\Sigma\setminus\Sigma_{p}$, the map $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\mapsto\xi_{v}(\frac{a}{d}\mod\varpi_{v})$ defines a character of $U_{v}$, where $\xi_{v}:k(v)^{\times}\to\mathcal{O}^{\times}$ is the trivial character. The product of $\xi_{v}$ can be viewed as a character $\xi$ of $U^{p}$ by projecting to $\prod_{v\in\Sigma\setminus\Sigma_{p}}U_{v}$. Using this, we can define a completed homology $M_{\psi,\xi}(U^{p})$ and a Hecke algebra $\mathbb{T}:=\mathbb{T}_{\psi,\xi}(U^{p})$. As $\bar{\rho}$ is modular, we know that $T_{v}-\mathrm{tr}\bar{\rho}(\operatorname{Frob}_{v}),v\notin\Sigma$ and $\varpi$ generate a maximal ideal $\mathfrak{m}$ of $\mathbb{T}$. Since $\bar{\rho}$ is absolutely irreducible, there is a two-dimensional representation $\rho_{\mathfrak{m}}:G_{F,\Sigma}\to\mathrm{GL}_{2}(\mathbb{T}_{\mathfrak{m}})$ with determinant $\chi$ such that the trace of $\operatorname{Frob}_{v}$ is $T_{v}$ for $v\notin\Sigma$.
Let $R^{\{\xi_{v}\}}$ be the universal deformation ring parametrizing all deformations of $\bar{\rho}$ satisfying $\operatorname{tr}(\rho)|_{I_{F_{v}}}=\xi_{v}+\xi_{v}^{-1}$ for all $v\in\Sigma\setminus\Sigma_{p}$, and with determinant $\chi$ By the universal property, there is a natural surjection $R^{\{\xi_{v}\}}\twoheadrightarrow\mathbb{T}_{\mathfrak{m}}$. Similar to Definition 4.1.3, we can also define pro-modular and nice primes in this case.
Definition 4.3.1.
We say a prime of $R^{\{\xi_{v}\}}$ is pro-modular if it comes from a prime of $\mathbb{T}_{\mathfrak{m}}$. A pro-modular prime $\mathfrak{q}$ of $R^{\{\xi_{v}\}}$ is nice if
  1. (1)
    $R^{\{\xi_{v}\}}/\mathfrak{q}$ is one-dimensional with characteristic $p$.
  2. (2)
    $\rho(\mathfrak{q})$ is absolutely irreducible and not induced from a quadratic extension of $F$. Here $\rho(\mathfrak{q})$ denotes the push-forward of the universal deformation on $R^{\{\mathbf{1}\}}$ to $k(\mathfrak{q})$.
  3. (3)
    $\rho(\mathfrak{q})|_{G_{F_{v}}}$ is trivial for $v\in\Sigma\setminus\Sigma_{p}$.
Proposition 4.3.2.
Let $\mathfrak{q}\in\operatorname{Spec}R^{\{\xi_{v}\}}$ be a nice prime. Then the kernel of $(R^{\{\xi_{v}\}})_{\mathfrak{q}}\to\mathbb{T}_{\mathfrak{q}}$ is nilpotent. Here, we also view it as a prime of $\mathbb{T}_{\mathfrak{m}}$.
Proof.
The proof is nearly the same as Theorem 4.1.4. One can see [Pan22, Section 8.2 & 8.3] and [Pan22, Proposition 9.0.6] for details. Note that we need to use local-global compatibility results in Section 3.3 instead. ∎
Proposition 4.3.3.
If $[F:\mathbb{Q}]\geq 4|\Sigma\setminus\Sigma_{p}|+2$, then the kernel of $R^{\{\xi_{v}\}}\to\mathbb{T}_{\mathfrak{m}}$ is nilpotent.
Proof.
The proof is the same as [Pan22, Proposition 9.0.4]. ∎
5. The proof in the residually reducible case
In this section, we will give the proof of the Fontaine-Mazur conjecture in the residually reducible case.
Let $p$ be an odd prime. Throughout this section, we denote by $K$ a $p$-adic local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$.
5.1. Setting
Our main goal is to prove the following theorem.
Theorem 5.1.1.
Let $F$ be a totally real abelian extension of $\mathbb{Q}$ in which $p$ completely splits. Suppose
[…]
is a continuous irreducible representation with the following properties
  • $\rho$ ramifies at only finitely many places.
  • Let $\bar{\rho}$ be the reduction of $\rho$ modulo $\varpi$. We assume its semi-simplification has the form $\bar{\chi}_{1}\oplus\bar{\chi}_{2}$ and ${\bar{\chi}_{1}}/\bar{\chi}_{2}$ can be extended to a character of $G_{\mathbb{Q}}$.
  • $\rho|_{G_{F_{v}}}$ is irreducible and de Rham of distinct Hodge-Tate weights for any $v|p$.
[…]
Then $\rho$ arises from a twist of a Hilbert modular form, i.e. a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{2}(\mathbb{A}_{F})$.
After a finite twist, we may assume $\bar{\chi}_{1}$ is trivial. Let $\Sigma$ be a finite set of places containing all the ramified finite places of $\rho$ and let $\Sigma_{p}$ be the set of primes above $p$. Write $\det\rho=\chi$ and $\bar{\chi}$ for its reduction modulo $\varpi$.
By Grunwald-Wang’s theorem (see [AT68, Theorem 5, page 80]) and abelian base change, we may suppose the number field $F$ satisfies the following assumptions:
  • $F$ is a totally real abelian extension of $\mathbb{Q}$ of even degree.
  • $p$ splits completely in $F$.
  • For any $v\in\Sigma\setminus\Sigma_{p}$, $p|\operatorname{Nm}(v)-1$.
  • $[F:\mathbb{Q}]>7|\Sigma\setminus\Sigma_{p}|+3$.
  • $\chi$ is unramified outside $\Sigma_{p}$, and $\chi(\operatorname{Frob}_{v})\equiv 1~{}\operatorname{mod}~{}\varpi$ for any $v\in\Sigma\setminus\Sigma_{p}$.
Consider the functor from $\mathfrak{U}_{\mathcal{O}}$ to the category of sets sending $R$ to the set of $2$-dimensional pseudo-representations of $G_{F,\Sigma}$ which lift $1+\bar{\chi}$ with determinant $\chi$. This is pro-represented by a CNL ring $R^{\mathrm{ps}}$ and $\mathrm{tr}\rho$ gives rise to a prime ideal $\mathfrak{p}\in\operatorname{Spec}R^{\mathrm{ps}}$. As $F$ is an abelian totally real field, the Leopoldt’s conjecture holds in this case, and hence the reducible locus of $R^{\mathrm{ps}}$ is of dimension at most $2$ (for example, see [Pan22, 5.4.5]).
We say that an irreducible component of $\operatorname{Spec}R^{\mathrm{ps}}$ is pro-modular, if every prime of it is pro-modular. Equivalently, its generic point comes from some big Hecke algebra.
Compared to [Pan22, Theorem 7.1.1], our theorem just removes the assumption $\bar{\chi}|_{G_{F_{v}}}\neq\omega^{\pm 1}$ for any $v|p$ if $p=3$. From now on, we assume $\bar{\chi}|_{G_{F_{v}}}=\omega^{-1}$ for any $v|p$ and $p=3$.
Lemma 5.1.2.
Suppose that Assumption 4.1.1 holds. Let $\mathfrak{r}$ be a one-dimensional irreducible pro-modular prime of $R^{\mathrm{ps}}$. Then every irreducible component of $R^{\mathrm{ps}}$ containing $\mathfrak{r}$ contains a nice prime, and hence is pro-modular.
Proof.
The proof is almost the same as [Zha24, Proposition 4.5.1]. The unique difference is that we need to use Corollary 3.3.10 instead of [Pan22, Theorem 3.6.1] in the proof of [Zha24, Proposition 4.5.1] to cover the case $p=3$.
Remark 5.1.3.
In [Pan22, Section 7.4], Pan’s strategy is to find a “potentially nice” prime in each irreducible component of $R^{\mathrm{ps}}$ of dimension at least $1+2[F:\mathbb{Q}]$. In his setting $(p\geq 5)$, Pan may assume $\bar{\chi}|_{G_{F_{v}}}=\omega\neq\omega^{-1}$, which implies that $\bar{\chi}$ is not quadratic. Thus, Pan may ignore the third condition in (2) of Definition 4.1.3 using the first part of Lemma 4.1.7. However, in our case (particularly $\bar{\chi}=\omega^{-1}$), $\bar{\chi}$ is possible to be quadratic, and it seems to be difficult to ensure the third condition to find nice primes.
One of the differences between our method in this section and Pan’s is that under our setting, we can find nice primes directly (using Lemma 4.2.5) rather than “potentially nice” ones. Furthermore, using the previous lemma, we only need to find one-dimensional irreducible pro-modular primes in the irreducible component, which is more convenient.
By Proposition 2.5.2, we can choose an irreducible component $C$ of $\operatorname{Spec}R^{\mathrm{ps}}$ containing $\mathfrak{p}$ of dimension at least $2[F:\mathbb{Q}]+1$. Combining Theorem 4.1.6, we only need to show that $C$ is pro-modular.
Remark 5.1.4.
Actually, we can show that under our setting, each irreducible component of $R^{\mathrm{ps}}$ of dimension at least $2[F:\mathbb{Q}]+1$ is pro-modular. This generalizes the big $R=\mathbb{T}$ theorem studied in [Zha24].
Consider the following map given by the universal property:
$R^{\mathrm{ps}}_{p}\to R^{\mathrm{ps}}$
where $R^{\mathrm{ps}}_{p}$ denotes the completed tensor product ^$\widehat{\bigotimes}_{v|p}R^{\mathrm{ps}}_{v}$ over $\mathcal{O}$ and $R^{\mathrm{ps}}_{v}$ denotes the universal pseudo-deformation ring which parametrizes all $2$-dimensional pseudo-representations of $G_{F_{v}}$ that lift $(1+\bar{\chi})|_{G_{F_{v}}}$. Note that we do not fix the determinant here. Let $R^{\mathrm{ps},\mathrm{ord}}_{v}$ be the quotient of $R^{\mathrm{ps}}_{v}$ that parametrizes all reducible liftings and $R^{\mathrm{ps},\mathrm{ord}}_{p}$ be the completed tensor product of $R^{\mathrm{ps},\mathrm{ord}}_{v},v|p$. We denote
$R^{\mathrm{ps},\mathrm{ord}}=R^{\mathrm{ps}}\otimes_{R^{\mathrm{ps}}_{p}}R^{\mathrm{ps},\mathrm{ord}}_{p};~{}C^{\mathrm{ord}}=\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}\cap C.$
Lemma 5.1.5.
We have $\dim C^{\mathrm{ord}}\geq 1$.
Proof.
For each $v|p$, by Proposition 2.3.3, the kernel of $R^{\mathrm{ps}}_{v}\twoheadrightarrow R^{\mathrm{ps},\mathrm{ord}}_{v}$ can be generated by two elements. Hence the kernel of $R^{\mathrm{ps}}\to R^{\mathrm{ps},\mathrm{ord}}$ can be generated by at most $2[F:\mathbb{Q}]$ elements. By Krull’s principal ideal theorem, we have $\dim C^{\mathrm{ord}}\geq 1$. ∎
Choose a prime $\mathfrak{q}\in C^{\mathrm{ord}}$ such that $\dim R^{\mathrm{ps}}/\mathfrak{q}=1$. Enlarging $\mathcal{O}$ if necessary, we may assume the normalization of $R^{\mathrm{ps}}/\mathfrak{q}$ is either $\mathcal{O}$ or isomorphic to $\mathbb{F}[[T]]$. There are three possibilities for the associated semi-simple representation $\rho(\mathfrak{q})$:
  1. (1)
    $\rho(\mathfrak{q})\cong\psi_{1}\oplus\psi_{2}$ is reducible and $\psi_{1}/\psi_{2}$ is not of the form $\varepsilon^{\pm 1}\theta$, where $\theta$ is a finite order character of $G_{F}$. We call this case generic reducible.
  2. (2)
    $\rho(\mathfrak{q})$ is irreducible.
  3. (3)
    $\rho(\mathfrak{q})\cong\psi_{1}\oplus\psi_{2}$ and $\psi_{1}/\psi_{2}=\varepsilon\theta$, for some finite order character $\theta$ of $G_{F}$. We call this case non-generic reducible.
5.2. The generic reducible case
In this subsection, we deal with the case $\rho(\mathfrak{q})\cong\psi_{1}\oplus\psi_{2}$ is reducible and $\psi_{1}/\psi_{2}$ is not of the form $\varepsilon^{\pm 1}\theta$, where $\theta$ is a finite order character of $G_{F}$.
For any $v|p$, let $\mathfrak{q}_{v}^{\mathrm{ord}}\in\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}_{v},\mathfrak{q}_{v}\in\operatorname{Spec}R^{\mathrm{ps}}_{v}$ be the pull-back of $\mathfrak{q}$. We first show that both $\mathfrak{q}_{v}^{\mathrm{ord}}$ and $\mathfrak{q}_{v}$ are of dimension one. If $\varpi\notin\mathfrak{q}_{v}^{\mathrm{ord}}$, then it is clear. If $\varpi\in\mathfrak{q}_{v}^{\mathrm{ord}}$, then the image of $G_{F}$ under $\psi_{1}$ is infinite (otherwise $\mathfrak{q}$ is the maximal ideal by Theorem 4.2.3). As $F$ is an abelian totally real field, we know that Leopoldt’s conjecture holds in this case, and hence the image of the composite map
$O_{F_{v}}^{\times}\to G_{F_{v}}^{\mathrm{ab}}\to G_{F,\Sigma}^{\mathrm{ab}}(p)$
has finite index inside $G_{F,\Sigma}^{\mathrm{ab}}(p)$. This implies that the image of $\psi_{1}|_{G_{F_{v}}}$ is infinite as well, and hence $\mathfrak{q}_{v}^{\mathrm{ord}}$ (resp. $\mathfrak{q}_{v}$) is not the maximal ideal. Moreover, we also have $(\frac{\psi_{1}}{\psi_{2}})|_{G_{F_{v}}}\neq\varepsilon^{\pm 1}$ from the generic reducibility assumption.
Lemma 5.2.1.
We have $\dim C^{\mathrm{ord}}\geq[F:\mathbb{Q}]+1$ in this case.
Proof.
By Proposition 2.3.5, for any $v|p$, the kernel of $(R^{\mathrm{ps}}_{v})_{\mathfrak{q}_{v}}\to(R^{\mathrm{ps},\mathrm{ord}}_{v})_{\mathfrak{q}_{v}^{\mathrm{ord}}}$ is principal. Localizing $C$ and $C^{\mathrm{ord}}$ at $\mathfrak{q}$ and using Krull’s principal ideal theorem, we know that the kernel of $(R^{\mathrm{ps}})_{\mathfrak{q}}\to(R^{\mathrm{ps},\mathrm{ord}})_{\mathfrak{q}}$ can be generated by at most $[F:\mathbb{Q}]$ elements. As $\dim C\geq 2[F:\mathbb{Q}]+1$ (by Proposition 2.5.2), the localization of $C^{\mathrm{ord}}$ at $\mathfrak{q}$ is of dimension at least $[F:\mathbb{Q}]$. This proves the lemma. ∎
Proposition 5.2.2.
Theorem 5.1.1 holds in this case.
Proof.
From the discussions in the previous subsection, we only need to show that $C$ is pro-modular, and it follows from Theorem 4.1.4, Lemma 4.2.5 and the previous lemma. ∎
Remark 5.2.3.
From the arguments above, we can also conclude that if for any $v|p$, there exist characters $\psi_{v,1}$ and $\psi_{v,2}$ satisfying $(\mathrm{tr}\rho(\mathfrak{q}))|_{G_{F_{v}}}=\psi_{v,1}+\psi_{v,2}$ and $\frac{\psi_{v,1}}{\psi_{v,2}}\neq\varepsilon^{\pm 1}$, then Theorem 5.1.1 also holds.
5.3. The irreducible case
In this subsection, we handle the irreducible case, i.e. $\rho(\mathfrak{q})$ is irreducible.
We first assume $\bar{\chi}\neq\omega^{-1}$. In this situation, we can follow the stategy of the proof in [Pan22, Section 7.4.2].
Lemma 5.3.1.
There exists an irreducible component $C^{\prime}$ of $\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}$ containing $\mathfrak{q}$ and a nice prime.
Proof.
Let $\psi^{univ}_{v,1},\psi^{univ}_{v,2}:G_{F_{v}}\to(R^{\mathrm{ps},\mathrm{ord}})^{\times}$ be the liftings of $\mathbf{1},\bar{\chi}|_{G_{F_{v}}}$ respectively. Let $\Sigma^{o}$ be the set of places $v|p$ such that
$\rho(\mathfrak{q})|_{G_{F_{v}}}\cong\begin{pmatrix}\psi^{univ}_{v,1}\operatorname{mod}\mathfrak{q}&*\\ 0&\psi^{univ}_{v,2}\operatorname{mod}\mathfrak{q}\end{pmatrix}$
and $\psi^{univ}_{v,1}$ is a lifting of $\mathbf{1}$. Recall that the universal pseudo-deformation ring $R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$ is a quotient of $R^{\mathrm{ps},\mathrm{ord}}$ parametrizing all pseudo-representations that are $\psi_{v,1}^{univ}$-ordinary (resp. $\psi_{v,2}^{univ}$-ordinary) if $v\in\Sigma^{o}$ (resp. $v\in\Sigma_{p}\setminus\Sigma^{o}$). We also view $\mathfrak{q}$ as a prime of $R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$.
Choose a lattice $\rho(\mathfrak{q})^{o}$ of $\rho(\mathfrak{q})$ such that its residual representation $\bar{\rho}_{b}$ is a non-split extension of $\mathbf{1}$ by $\bar{\chi}$. Consider the universal ordinary deformation $R^{\mathrm{ord}}_{b}$ which parametrizes all deformations $\rho_{b}$ of $\bar{\rho}_{b}:G_{F,\Sigma}\to\mathrm{GL}_{2}(\mathbb{F})$ with determinant $\chi$ such that for any $v|p$, $\rho_{b}|_{G_{F_{v}}}\cong\begin{pmatrix}\psi_{v,1}&*\\ 0&*\end{pmatrix}$, where $\psi_{v,1}$ is a lifting of $\mathbf{1}$ (resp. $\bar{\chi}|_{G_{F_{v}}}$) if $v\in\Sigma^{o}$ (resp. $v\in\Sigma_{p}\setminus\Sigma^{o}$). Let $\mathfrak{q}_{b}$ be the prime of $R^{\mathrm{ord}}_{b}$ corresponding to $\rho(\mathfrak{q})^{o}$. Using similar arguments as in Proposition 2.2.5, we have an isomorphism
$\widehat{(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}}}\cong\widehat{(R^{\mathrm{ord}}_{b})_{\mathfrak{q}_{b}}}.$
As $\bar{\chi}\neq\omega^{-1}$, we have $H^{0}(G_{F,\Sigma},\operatorname{ad}^{0}(\bar{\rho}_{b}(1)))=0$. Then by Proposition 4.2.6, we have $\dim(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}}=\dim(R^{\mathrm{ord}}_{b})_{\mathfrak{q}_{b}}\geq[F:\mathbb{Q}].$ Then there exists an irreducible component $C^{\prime}$ of $\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$ (hence of $\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}$) containing $\mathfrak{q}$ of dimension at least $[F:\mathbb{Q}]+1$. Then the result follows from Lemma 4.2.5. ∎
Proposition 5.3.2.
Theorem 5.1.1 holds if $\rho(\mathfrak{q})$ is irreducible and $\bar{\chi}\neq\omega^{-1}$.
Proof.
From the previous lemma and Theorem 4.1.4, we know that $\mathfrak{q}$ is a one-dimensional irreducible pro-modular prime. Then by Lemma 5.1.2, we know that $C$ is pro-modular, and hence $\mathfrak{p}$ is pro-modular. ∎
Now we assume $\bar{\chi}=\omega^{-1}$. Note that in this situation, the arguments above may not work because the lower bound of each irreducible component of $R_{b}^{\mathrm{ord}}$ containing $\mathfrak{q}$ is $[F:\mathbb{Q}]$, and we cannot use Theorem 4.2.3 directly. See Remark 4.2.7.
Keep notations as above. Then by Proposition 4.2.6, we have $\dim(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}}=\dim(R^{\mathrm{ord}}_{b})_{\mathfrak{q}_{b}}\geq[F:\mathbb{Q}]-1.$ We choose an irreducible component $C_{1}^{\mathrm{ord}}$ of $\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$ containing $\mathfrak{q}$ of dimension $[F:\mathbb{Q}]$ (if its dimension is at least $[F:\mathbb{Q}]+1$, then we can use the same proof as above), and suppose its generic point is $\mathfrak{q}_{1}$. Consider the natural maps
$\Lambda_{F}\to R^{\mathrm{ps},\mathrm{ord}}\twoheadrightarrow R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}},$
where $\Lambda_{F}$ is the Iwasawa algebra of Krull dimension $1+[F:\mathbb{Q}]$ defined in Section 4.2. Viewing $\mathfrak{q}_{1}$ as a prime of $\Lambda_{F}$ via the composite map, using Theorem 4.2.3, we find that $\mathfrak{q}_{1}$ is a prime of $\Lambda_{F}$ of height $1$.
Write $\Lambda_{F}=\mathcal{O}[[X_{1},...,X_{d}]]$, where $d=[F:\mathbb{Q}]$, $X_{i}$ corresponds to a generator of $O_{F_{v_{i}}}^{\times}(p)$ (the $p$-adic completion of $O_{F_{v_{i}}}^{\times}$) for $v_{i}|p$. Write $(\mathrm{tr}\rho(\mathfrak{q}))|_{G_{F_{v_{i}}}}=\psi_{v_{i},1}+\psi_{v_{i},2}$. If $\frac{\psi_{v_{i},1}}{\psi_{v_{i},2}}=\varepsilon^{\pm 1}$, then $X_{i}$ has to take a value in a finite subset $A_{i}$ of $\mathfrak{m}_{\bar{K}}$. The finiteness of the set $A_{i}$ is because the determinant is $\chi|_{G_{F_{v_{i}}}}$, and hence $\psi_{v_{i},1}^{2}=(\chi\varepsilon)|_{G_{F_{v_{i}}}}$ or $(\chi\varepsilon^{-1})|_{G_{F_{v_{i}}}}$. Enlarging $\mathcal{O}$ if necessary, we assume that $A_{i}$ is a subset of $\mathfrak{m}_{K}$.
Proposition 5.3.3.
Write $\mathfrak{p}_{i,j}$ for the height $1$ prime of $\Lambda_{F}$ generated by $X_{i}-a_{i,j}$, where $a_{i,j}\in A_{i}$. If $\mathfrak{q}_{1}\neq\mathfrak{p}_{i,j}$ for any $i,j>0$, then we can find a prime $\mathfrak{q}_{1}^{\prime}$ of $\Lambda_{F}$ of dimension $1$ such that $\mathfrak{q}_{1}\subset\mathfrak{q}_{1}^{\prime}$, and $\mathfrak{p}_{i,j}\nsubseteq\mathfrak{q}_{1}^{\prime}$ for any $i,j>0$.
Proof.
We can prove this result by induction on $d$ based on the Weierstrass preparation theorem ([Lan90, Theorem 2.2, page 130]), but here we give an easy proof.
Consider the CNL ring $\Lambda_{F}/\mathfrak{q}_{1}$. If our result is not true, then by [Zha24, Lemma 3.2.5], the closed subsets (of $\operatorname{Spec}\Lambda_{F}$) $V(\mathfrak{p}_{i,j})\cap V(\mathfrak{q}_{1})$ form a cover of $\operatorname{Spec}(\Lambda_{F}/\mathfrak{q}_{1})=V(\mathfrak{q}_{1})$. In other words, we have $V(\mathfrak{q}_{1})\subset\cup_{i,j}V(\mathfrak{p}_{i,j})$. As $\mathfrak{q}_{1}$ and $\mathfrak{p}_{i,j}$ are of height one, we have $\mathfrak{q}_{1}=\mathfrak{p}_{i,j}$ for some $i,j>0$, which is against to our assumption.
Now we may assume $\mathfrak{q}_{1}=(X_{i}-a_{i,j})$ for some $i,j>0$. Otherwise, using the previous proposition, Theorem 4.2.3 and the going-up property, we can find a one-dimensional prime $\mathfrak{q}_{1}^{\prime\prime}$ of $\operatorname{Spec}R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}}$ such that for any $v|p$, if we write $(\mathrm{tr}\rho(\mathfrak{q}_{1}^{\prime\prime}))|_{G_{F_{v}}}=\psi_{v,1}+\psi_{v,2}$, then $\frac{\psi_{v,1}}{\psi_{v,2}}\neq\varepsilon^{\pm 1}$. Arguing as the generic reducible case, Theorem 5.1.1 holds in this case. See Remark 5.2.3.
Choose a one-dimensional prime $\mathfrak{q}_{2}$ of $\Lambda_{F}$ containing $\mathfrak{q}_{1}$ such that for any $t\neq i$, $X_{t}-a_{t,j}\notin\mathfrak{q}_{2}$. By the going-up property, we can find a one-dimensional prime $\mathfrak{q}_{2}^{\prime}$ contained in $C_{1}^{\mathrm{ord}}$. Arguing as in the previous subsection, we know that $\rho(\mathfrak{q}_{2}^{\prime})$ has to be irreducible. From our construction, if we write $(\mathrm{tr}\rho(\mathfrak{q}_{2}^{\prime}))|_{G_{F_{v}}}=\psi_{v,1}+\psi_{v,2}$, then we have $\frac{\psi_{v,1}}{\psi_{v,2}}\neq\varepsilon^{\pm 1}$ for any $v\neq v_{i}$. This implies $H^{0}(G_{F,\Sigma},\operatorname{ad}^{0}\rho(\mathfrak{q}_{2}^{\prime})(1))=0$.
Proposition 5.3.4.
Theorem 5.1.1 holds if $\bar{\chi}=\omega^{-1}$ and $\rho(\mathfrak{q})$ is irrreducible.
Proof.
Let $\bar{\rho_{b^{\prime}}}$ be the residue representation of $\rho(\mathfrak{q}_{2}^{\prime})$, and we consider its (unframed) universal ordinary deformation $R_{b^{\prime}}^{\mathrm{ord}}$. Assume the representation $\rho(\mathfrak{q}_{2}^{\prime})$ defines a prime $\mathfrak{q}_{b^{\prime}}$ of $R_{b^{\prime}}^{\mathrm{ord}}$, and we have an isomorphism
$\widehat{(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}_{2}^{\prime}}}\cong\widehat{(R^{\mathrm{ord}}_{b^{\prime}})_{\mathfrak{q}_{b^{\prime}}}}.$
By Proposition 2.5.7, we have $\dim\widehat{(R^{\mathrm{ord}}_{b^{\prime}})_{\mathfrak{q}_{b^{\prime}}}}\geq[F:\mathbb{Q}]$, and hence $\dim(R^{\mathrm{ps},\mathrm{ord}}_{\Sigma^{o}})_{\mathfrak{q}_{2}^{\prime}}\geq[F:\mathbb{Q}]$. Then by Theorem 4.1.4 and Lemma 4.2.5, the prime $\mathfrak{q}_{2}^{\prime}$ is pro-modular. Using Lemma 5.1.2, each irreducible component of $\operatorname{Spec}R^{\mathrm{ps}}$ containing $\mathfrak{q}_{2}^{\prime}$ is pro-modular. In particular, $\mathfrak{q}_{1}$ is pro-modular, and hence $\mathfrak{q}$ is pro-modular. Using Lemma 5.1.2 again, we know the pro-modularity of $\mathfrak{p}$.
Remark 5.3.5.
Another possible way to prove this case is as follows.
If the degree $[F:\mathbb{Q}]$ is sufficiently large (in particular greater than $\dim H^{1}_{\Sigma^{o}}$), then by [Pan22, Corollary 5.5.2], any one-dimensional irreducible prime is pro-modular. (Note that here we need to use an $R=\mathbb{T}$ theorem in the ordinary case. See [Pan22, Proposition 5.3.3]. In Pan’s setting, there is an assumption $[F_{v}:\mathbb{Q}_{p}]\geq 2$ for any $v|p$, which is used to prove [Pan22, Lemma 5.3.4]. However, we may use Proposition 2.1.4 instead in our case. Then it may be possible to remove this assumption.) Then by Lemma 5.1.2, we obtain the pro-modularity of $\mathfrak{q}$ and $\mathfrak{p}$. Thus, we just need to find an abelian base change $F_{1}/F$ with larger degree (than the dimension of the Selmer group, or equivalently the order of the $p$-ideal class group) and prescribed behavior at each place in $\Sigma$.
In the ordinary case, we can use [Was78, Theorem] to find a suitable base change. However, in the non-ordinary case, we need the assumption that $p$ splits completely in $F$ (to use the $p$-adic Langlands correspondence), and it is difficult to use [Was78, Theorem] in our case.
In [SW99, Theorem B], the authors give an assumption on the $p$-rank of the relative class group of $L/F$ instead of [Was78], where $L$ is some CM extension of $F$. See [SW99, Hypothesis H]. Such hypothesis is studied for imaginary quadratic fields in [Wil15]. If one can show analogous results for general abelian CM fields, then we may prove the irreducible case by the previous arguments.
A similar strategy will work in the non-generic reducible case as we can bound the dimension of the cohomology group easily.
5.4. The non-generic reducible case
In this subsection, we cope with the case $\rho(\mathfrak{q})\cong\psi_{1}\oplus\psi_{2}$, $\psi_{1}/\psi_{2}=\varepsilon\theta$ for some finite order character $\theta$ of $G_{F}$. In this case, the proof is nearly the same as [Pan22, Section 7.4.5]. We briefly recall Pan’s proof here.
As $\theta$ is of finite order, we know that $\mathfrak{q}$ is of characteristic 00. Otherwise, $\mathfrak{q}$ is the maximal ideal of $R^{\mathrm{ps}}$. Enlarging $\mathcal{O}$ if necessary, we may assume $k(\mathfrak{q})=K$. The following lemma is crucial.
Lemma 5.4.1.
The natural restriction map
$H^{1}(G_{F,\Sigma},K(\psi_{2}/\psi_{1}))\to\bigoplus_{v|p}H^{1}(G_{F_{v}},K(\psi_{2}/\psi_{1}))$
is an isomorphism. In particular, $\dim_{K}H^{1}(G_{F,\Sigma},K(\psi_{2}/\psi_{1}))=[F:\mathbb{Q}]$.
Proof.
See [Pan22, Lemma 7.4.10]. ∎
For any non-zero class $B\in\operatorname{Ext}^{1}_{K[G_{F,\Sigma}]}(\psi_{1},\psi_{2})\cong H^{1}(G_{F,\Sigma},K(\psi_{2}/\psi_{1}))$, it corresponds to a non-split extension $\rho_{B}:G_{F,\Sigma}\to\mathrm{GL}_{2}(K)$ of $\psi_{1}$ by $\psi_{2}$. We may assume $\rho_{B}$ is of the form
$\begin{pmatrix}\psi_{2}&b_{B}\\ 0&\psi_{1}\end{pmatrix}.$
We denote the universal deformation ring of $\rho_{B}$ with fixed determinant $\chi$ by $R_{B}$. Let $\rho_{B}^{o}:G_{F,\Sigma}\to\mathrm{GL}_{2}(\mathcal{O})$ be a lattice of $\rho_{B}$ such that its mod $\varpi$ reduction $\bar{\rho}_{b}:G_{F,\Sigma}\to\mathrm{GL}_{2}(\mathbb{F})$ is non-split. We may consider the universal deformation ring $R_{b}$ of $\bar{\rho}_{b}$ with fixed determinant $\chi$. Then $\rho_{B}^{o}$ naturally gives rise to a prime $\mathfrak{p}_{B}$ of $R_{b}$.
Lemma 5.4.2.
Keep notations as above.
  1. (1)
    We have $\dim R_{B}/\mathfrak{P}\geq 2[F:\mathbb{Q}]$ for any minimal prime $\mathfrak{P}$ of $R_{B}$.
  2. (2)
    The connectedness dimension $c(R_{B})$ is at least $2[F:\mathbb{Q}]-1$.
  3. (3)
    Let $\operatorname{Spec}R^{\mathrm{red}}_{B}$ be the reducible locus of $\operatorname{Spec}R_{B}$. Then we have $\dim\operatorname{Spec}R_{B}^{\mathrm{red}}\leq[F:\mathbb{Q}]+1$.
  4. (4)
    The natural map $R_{B}\to\widehat{(R_{b})_{\mathfrak{p}_{B}}}$ is an isomorphism.
  5. (5)
    By evaluating the trace of the universal deformation, we get a natural map $R^{\mathrm{ps}}\to R_{B}$. Suppose $\mathfrak{Q}\in\operatorname{Spec}R_{B}\setminus\operatorname{Spec}R^{\mathrm{red}}_{B}$. Let $\mathfrak{Q}^{\mathrm{ps}}=\mathfrak{Q}\cap R^{\mathrm{ps}}$. Then
    $\dim R^{\mathrm{ps}}/\mathfrak{Q}^{\mathrm{ps}}\geq 1+\dim R_{B}/\mathfrak{Q}.$
  6. (6)
    The natural map $\operatorname{Spec}R_{B}\to\operatorname{Spec}R^{\mathrm{ps}}$ maps minimal primes to minimal primes.
Proof.
See [Pan22, Lemma 7.4.11 & 7.4.12, Corollary 7.4.13]. ∎
Definition 5.4.3.
For any $B\in\operatorname{Ext}^{1}_{K[G_{F,S}]}(\psi_{1},\psi_{2})$, we define $Z_{B}$ as the set of irreducible components of $R^{\mathrm{ps}}$ whose generic points lie in the image of $\operatorname{Spec}R_{B}\to\operatorname{Spec}R^{\mathrm{ps}}$. It follows from the last part of the previous lemma that $\bigcup_{C^{\prime}\in Z_{B}}C^{\prime}$ contains the image of $\operatorname{Spec}R_{B}\to\operatorname{Spec}R^{\mathrm{ps}}$.
Lemma 5.4.4.
Suppose that Assumption 4.1.1 holds. For any non-zero $B$, if one of the irreducible component in $Z_{B}$ is pro-modular, then every irreducible component in $Z_{B}$ is pro-modular.
Proof.
Let $C_{B}=\bigcup_{C^{\prime}\in Z_{B}}C^{\prime}$, and write $C_{B}=Z_{1}\cup Z_{2}$, where $Z_{i},i=1,2,$ are non-empty sets of irreducible components in $C_{B}$. Let $C_{i}$ be the set of irreducible components of $\operatorname{Spec}R_{B}$ whose images are contained in the set $Z_{i}$. Then $C_{1},C_{2}$ form a covering of the set consisting of all irreducible components of $\operatorname{Spec}R_{B}$. Hence by (2) of Lemma 5.4.2, $\dim C_{1}\cap C_{2}\geq 2[F:\mathbb{Q}]-1$. Choose a point $\mathfrak{Q}\in C_{1}\cap C_{2}$ with $\dim R_{B}/\mathfrak{Q}\geq 2[F:\mathbb{Q}]-1>[F:\mathbb{Q}]+1$. Then $\mathfrak{Q}\notin\operatorname{Spec}R^{\mathrm{red}}_{B}$ by (3) of Lemma 5.4.2. It follows from the (5) of Lemma 5.4.2 that $R^{\mathrm{ps}}/(\mathfrak{Q}\cap R^{\mathrm{ps}})$ has dimension at least $2[F:\mathbb{Q}]$, and $\mathfrak{Q}\cap R^{\mathrm{ps}}\in Z_{1}\cap Z_{2}$. This shows that the connectedness dimension of $C_{B}$ is at least $2[F:\mathbb{Q}]$.
Arguing as the proof of Lemma 5.1.2, we obtain the result. ∎
Fix a place $v_{0}$ of $F$ above $p$. Let $B_{0}\in H^{1}(G_{F,\Sigma},K(\psi_{2}/\psi_{1}))$ be a non-zero class such that under the restriction map, its image in $H^{1}(G_{F_{v}},K(\psi_{2}/\psi_{1})),v|p$ is zero unless $v=v_{0}$. The existence is guaranteed by Lemma 5.4.1. Then $B_{0}$ defines a non-split reducible Galois representation $\rho_{B_{0}}:G_{F,\Sigma}\to\mathrm{GL}_{2}(K)$. Let $\bar{\rho}_{b_{0}}:G_{F,\Sigma}\to\mathrm{GL}_{2}(\mathbb{F})$ be its residue representation, and we can find that $\bar{\rho}_{b_{0}}|_{G_{F_{v}}}$ is completely reducible for $v\in\Sigma_{p}\setminus\{v_{0}\}$. Let $R_{b_{0}}$ be the universal deformation ring of $\bar{\rho}_{b_{0}}$ with determinant of $\chi$, and consider its quotient $R_{\{v_{0}\}}^{\mathrm{ord}}$ which parametrizes deformations $\rho_{b}$ such that for any $v|p$, $\rho_{b}|_{G_{F_{v}}}\cong\begin{pmatrix}\psi_{v,1}&*\\ 0&*\end{pmatrix}$, where $\psi_{v,1}$ is a lifting of $\bar{\psi}_{1}$ (resp. $\bar{\psi}_{2}$) if $v\neq v_{0}$ (resp. $v=v_{0}$).
Lemma 5.4.5.
There exists an irreducible component in $Z_{B_{0}}$ which is pro-modular.
Proof.
Let $\mathfrak{p}_{0}$ be the prime of $R_{\{v_{0}\}}^{\mathrm{ord}}$ corrseponding to the representation $\rho_{B_{0}}$. Choose a minimal prime $\mathfrak{P}$ of $R_{\{v_{0}\}}^{\mathrm{ord}}$ contained in $\mathfrak{p}_{0}$. Let $\rho(\mathfrak{P})$ be the representation defined by the natural surjection $R_{\{v_{0}\}}^{\mathrm{ord}}\twoheadrightarrow R_{\{v_{0}\}}^{\mathrm{ord}}/\mathfrak{P}$, and we claim that it is irreducible. If not, we write
$\rho(\mathfrak{P})\cong\begin{pmatrix}\tilde{\psi_{2}}&*\\ 0&\tilde{\psi_{1}}\end{pmatrix}.$
As Leopoldt’s conjecture holds for $F$, we can choose $\tau_{0}\in G_{F,\Sigma}$ whose image in $G_{F,\Sigma}^{\mathrm{ab}}(p)\otimes\mathbb{Q}_{p}$ forms a basis. Let $\mathfrak{P}^{\prime}$ be a minimal prime of $R^{\mathrm{ord}}_{\{v_{0}\}}/\mathfrak{P}$ containing $\psi_{2}(\tau_{0})-\tilde{\psi_{2}}(\tau_{0})$. By Krull’s principal ideal theorem, $\mathfrak{P}^{\prime}$ is of height at most $1$. We have $\psi_{i}\equiv\tilde{\psi_{i}}\mod\mathfrak{P}^{\prime},i=1,2$ as the determinant is $\chi$. Therefore, $\rho(\mathfrak{P}^{\prime}){:=}\rho(\mathfrak{P})\mod\mathfrak{P}^{\prime}$ is an extension of $\psi_{1}$ by $\psi_{2}$, and satisfies $\rho(\mathfrak{P}^{\prime})|_{G_{F_{v}}}$ is completely reducible for $v\in\Sigma_{p}\setminus\{v_{0}\}$ (by the definition of $R_{\{v_{0}\}}^{\mathrm{ord}}$). Combining Lemma 5.4.1, we have $\mathfrak{P}^{\prime}=\mathfrak{p}_{0}$. This implies that $\mathfrak{p}_{0}$ is of dimension $1$ and of height at most $1$, and thus $\dim R^{\Delta}_{\{v_{0}\}}/\mathfrak{P}\leq 2$. By Proposition 4.2.6, each irreducible component of $R_{\{v_{0}\}}^{\mathrm{ord}}$ is of dimension at least $[F:\mathbb{Q}]>3$. We get a contradiction.
Choose a one-dimensional prime $\mathfrak{q}^{\prime\prime}$ of $R^{\mathrm{ord}}_{\{v_{0}\}}$ containing $\mathfrak{P}$ such that $\rho(\mathfrak{q}^{\prime\prime})$ is irreducible, and let $\mathfrak{q}^{\prime}=\mathfrak{q}^{\prime\prime}\cap R^{\mathrm{ps}}$. Now we can use the same proof as in the irreducible case (replacing $\mathfrak{q}$ by $\mathfrak{q}^{\prime}$), and we know that any irreducible component of $R^{\mathrm{ps}}$ containing $\mathfrak{q}^{\prime}$ contains a nice prime, and hence pro-modular. Let $\mathfrak{Q}$ be a minimal prime of $R_{b_{0}}$ contained in $\mathfrak{P}\subseteq\mathfrak{p}_{0}$. By (4) of Lemma 5.4.2, $\mathfrak{Q}$ is in the image of $\operatorname{Spec}R_{B_{0}}\to\operatorname{Spec}R_{b_{0}}$. This proves the lemma.
Recall that $C$ is an irreducible component of $R^{\mathrm{ps}}$ containing both $\mathfrak{p}$ and $\mathfrak{q}$. Enlarging $\mathcal{O}$ if necessary, we can assume $C$ is geometrically irreducible. Choose a prime ideal $\mathfrak{q}_{0}$ of $R^{\mathrm{ps}}$ (see the last paragraph of [Pan22, page 1153] for the existence) such that
  1. (1)
    $\mathfrak{q}_{0}\subset\mathfrak{q}$ and $\dim R^{\mathrm{ps}}/\mathfrak{q}_{0}=2$.
  2. (2)
    $C$ contains $\mathfrak{q}_{0}$ and is the only irreducible component having this property.
  3. (3)
    $\rho(\mathfrak{q}_{0})$ is irreducible.
Consider the $\mathfrak{q}$-adic completion $A_{0}$ of $(R^{\mathrm{ps}}/\mathfrak{q}_{0})_{\mathfrak{q}}$. This is a one-dimensional CNL ring with residue field $K$. Choose a minimal prime $\mathfrak{Q}_{A_{0}}$ of $A_{0}$ and let $A_{1}=A_{0}/\mathfrak{Q}_{A_{0}}$. We denote the normal closure of $A_{1}$ by $A_{2}$. Then $A_{2}\cong\tilde{K}[[T]]$ for some finite extension $\tilde{K}$ of $K$. Enlarging $K$ if necessary, we assume $K=\tilde{K}$, and we have an injection $R^{\mathrm{ps}}/\mathfrak{q}_{0}\hookrightarrow A_{2}$. Then we can choose a lattice of $\rho(\mathfrak{q}_{0})$:
$\rho(\mathfrak{q}_{0})^{o}:G_{F,\Sigma}\to\mathrm{GL}_{2}(A_{2})$
such that its mod $T$ reduction is a non-split extension of $\psi_{1}$ by $\psi_{2}$. Denote this extension class by $B_{1}$. Then $\mathfrak{q}_{0}$ lies in the image of $\operatorname{Spec}R_{B_{1}}\to\operatorname{Spec}R^{\mathrm{ps}}$. Since $\mathfrak{q}_{0}$ only contains one minimal prime, $C\in Z_{B_{1}}$.
Lemma 5.4.6.
There exist irreducible components $C_{B_{0}}\in Z_{B_{0}},C_{B_{1}}\in Z_{B_{1}}$ such that $\dim C_{B_{0}}\cap C_{B_{1}}\geq[F:\mathbb{Q}]+2$.
Proof.
The arguments are the same as those in [Pan22, page 1154 & 1155]. We recall his proof here.
Let $\tilde{F}$ be the splitting field of $\psi_{2}/\psi_{1}$. As $\psi_{1}/\psi_{2}=\varepsilon\theta$, $H^{i}(\operatorname{Gal}(\tilde{F}/F),K(\psi_{2}/\psi_{1}))=0$ for $i\geq 1$. Write $H:=\ker\{\psi_{2}/\psi_{1}:G_{F,\Sigma}\to K^{\times}\}$. By the Hochschild-Serre exact sequence (see [Rub00, Proposition B 2.5]), we have an isomorphism:
$H^{1}(G_{F,\Sigma},K(\psi_{2}/\psi_{1}))=\operatorname{Hom}_{\operatorname{Gal}(\tilde{F}/F)}(H,K(\psi_{2}/\psi_{1})).$
If $K\cdot B_{1}=K\cdot B_{0}$, then $Z_{B_{1}}=Z_{B_{0}}$ and the result follows from the previous lemma. From now on, we may assume $K\cdot B_{1}\neq K\cdot B_{0}$.
Extend $B_{0},B_{1}$ to a basis $B_{0},\cdots,B_{[F:\mathbb{Q}]-1}$ of $H^{1}(G_{F,\Sigma},K(\psi_{2}/\psi_{1}))$ such that we can find $\sigma_{0},\cdots,\sigma_{[F:\mathbb{Q}]-1}\in H$ as a dual basis via the pairing $H\times H^{1}(G_{F,\Sigma},K(\psi_{2}/\psi_{1}))\to K(\psi_{2}/\psi_{1})$.
Write
$\rho^{\operatorname{univ}}_{B_{0}}(\sigma)=\begin{pmatrix}a_{0}(\sigma)&b_{0}(\sigma)\\ c_{0}(\sigma)&d_{0}(\sigma)\end{pmatrix},\sigma\in G_{F,\Sigma},$
where $\rho^{\operatorname{univ}}_{B_{0}}:G_{F,\Sigma}\to\mathrm{GL}_{2}(R_{B_{0}})$ is the universal deformation satisfying $b_{0}(\sigma^{*})=c_{0}(\sigma^{*})=0$ for some fixed complex conjugation $\sigma^{*}$. Consider the following deformation of $\rho_{B_{0}}$:
$\rho_{B_{0}}+T\rho_{B_{1}}:G_{F,\Sigma}\to\mathrm{GL}_{2}(K[[T]]),~{}\sigma\mapsto\begin{pmatrix}\psi_{2}(\sigma)&b_{B_{0}}(\sigma)+Tb_{B_{1}}(\sigma)\\ 0&\psi_{1}(\sigma)\end{pmatrix}.$
This gives rise to a prime $\mathfrak{q}_{01}$ of $R_{B_{0}}$. Clearly, we have $b_{0}(\sigma_{i})\in\mathfrak{q}_{01},i=2,\cdots,[F:\mathbb{Q}]-1$. Let $\mathfrak{Q}_{01}$ be a minimal prime of $(R_{B_{0}})_{\mathfrak{q}_{01}}/(b_{0}(\sigma_{2}),\cdots,b_{0}(\sigma_{[F:\mathbb{Q}]-1}))$. Viewing $\mathfrak{Q}_{01}$ as a prime of $(R_{B_{0}})$, by (1) of Lemma 5.4.2 and Krull’s principal ideal theorem, we have $\dim R_{B_{0}}/\mathfrak{Q}_{01}\geq[F:\mathbb{Q}]+2$. Hence $\mathfrak{Q}_{01}\notin\operatorname{Spec}R^{\mathrm{red}}_{B_{0}}$ by (3) of Lemma 5.4.2. Let $\mathfrak{Q}^{\prime}_{01}=\mathfrak{Q}_{01}\cap R^{\mathrm{ps}}$. Consider the ideal $I_{0}$ of $R^{\mathrm{ps}}/\mathfrak{Q}_{01}^{\prime}$ generated by $y(\sigma_{0},\tau),\tau\in G_{F,\Sigma}$. Then by [Pan22, Lemma 7.4.16], we have $\dim R^{\mathrm{ps}}/(\mathfrak{Q}^{\prime}_{01},I_{0})\geq 2+[F:\mathbb{Q}]$.
Now choose a minimal prime of $R^{\mathrm{ps}}$ containing $\mathfrak{Q}^{\prime}_{01},I_{0}$. Choose a prime $\mathfrak{q}_{1}$ (see [Pan22, page 1155] for the existence) with the following properties:
  1. (1)
    $\mathfrak{P}_{01}\subseteq\mathfrak{q}_{1}\subseteq\mathfrak{q}$.
  2. (2)
    $\dim R^{\mathrm{ps}}/\mathfrak{q}_{1}=2$.
  3. (3)
    $\rho(\mathfrak{q}_{1})$ is irreducible.
  4. (4)
    Any irreducible component of $R^{\mathrm{ps}}$ containing $\mathfrak{q}_{1}$ also contains $\mathfrak{P}_{01}$.
From our construction, we can check that $\mathfrak{q}_{1}$ belongs to the image of $\operatorname{Spec}R_{B_{1}}\to\operatorname{Spec}R^{\mathrm{ps}}$. In particular, there exists an irreducible component $C_{B_{1}}\in Z_{B_{1}}$ that contains $\mathfrak{q}_{1}$, and we have $\mathfrak{P}_{01}\in C_{B_{1}}$.
As $\mathfrak{Q}^{\prime}_{01}$ is in the image of $\operatorname{Spec}R_{B_{0}}\to\operatorname{Spec}R^{\mathrm{ps}}$, we can find an irreducible component $C_{B_{0}}\in Z_{B_{0}}$ containing $\mathfrak{Q}^{\prime}_{01}$, hence also containing $\mathfrak{P}_{01}$. Now we have $\dim C_{B_{0}}\cap C_{B_{1}}\geq\dim R^{\mathrm{ps}}/\mathfrak{P}_{01}\geq[F:\mathbb{Q}]+2$. This proves the lemma. ∎
Proposition 5.4.7.
Theorem 5.1.1 holds if we are in the non-generic reducible case.
Proof.
By Lemma 5.4.4 and Lemma 5.4.5, we know that $C_{B_{0}}$ is pro-modular. Using the previous lemma, we can find a one-dimensional irreducible pro-modular prime contained in $C_{B_{0}}\cap C_{B_{1}}$. By Lemma 5.1.2, we know that $C_{B_{1}}$ is pro-modular. Using Lemma 5.4.4 again, we know that $C$ is pro-modular. Hence, $\mathfrak{p}$ is pro-modular. ∎
Remark 5.4.8.
In the case $\bar{\chi}=\omega^{-1}$ (hence quadratic), it seems to be difficult to find a nice prime in $C_{B_{0}}\cap C_{B_{1}}$ as Pan argues in [Pan22, Section 7.4] since we cannot use the first part of Lemma 4.1.7. In other words, Lemma 5.1.2 may be essential in this case.
5.5. Potential pro-modularity
In this subsection, we conclude our potential pro-modularity argument.
Let $F$ be an abelian totally real field in which $p$ splits completely. Write $\Sigma_{p}$ as the set of places of $F$ above $p$ and let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$. Let $\bar{\chi}:G_{F,\Sigma}\to\mathbb{F}^{\times}$ be a continuous odd character and we suppose it can be extended to a character of $G_{\mathbb{Q}}$. Assume $\chi:G_{F,\Sigma}\to\mathcal{O}^{\times}$ is a de Rham character as a lifting of $\bar{\chi}$.
Theorem 5.5.1.
Assume that for any place $v|p$, $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1}$. Then there exists an abelian totally real extension $F_{1}/F$ of even degree such that $p$ splits completely in $F_{1}$ and for any irreducible component of the universal pseudo-deformation ring $R_{F_{1}}^{\mathrm{ps}}$ (parametrizing all pseudo-deformations of $1+\bar{\chi}|_{G_{F_{1}}}$ unramified out places lying above $\Sigma$ with fixed determinant $\chi$) of dimension at least $1+2[F_{1}:\mathbb{Q}]$, it is pro-modular of dimension $1+2[F_{1}:\mathbb{Q}]$.
Proof.
For the case $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1},\omega^{-1}$ for any $v|p$ and $p\geq 5$, the proof is a combination of [Zha24, Theorem 1.0.2] and Grunwald-Wang’s theorem [AT68, Theorem 5, page 80].
For the case $\bar{\chi}|_{G_{F_{v}}}=\omega^{-1}$ and $p=3$, it is proved in this section. Note that in each case of this section, we conclude that the irreducible component $C$ is pro-modular. For the dimension of $C$, from Step 4 of the proof of Theorem 4.1.4, we can find that if $C$ contains a nice prime, then it is of dimension $1+2[F_{1}:\mathbb{Q}]$, and our result follows from Lemma 5.1.2.
For the case $\bar{\chi}|_{G_{F_{v}}}=\omega^{-1}$ and $p\geq 5$, the proof is almost the same as the case $p=3$. We only need to replace the results in Section 2.3 by [Pan22, Lemma 7.4.4 & Lemma 7.4.8]. ∎
We fix an abelian extension $F_{1}/F$ satisfying the properties in Theorem 5.5.1. Similarly as in Section 4.1, we can define the big Hecke algebra $\mathbb{T}_{\psi}$ (without the character $\xi$), and from our construction of $F_{1}$ and the proof of Theorem 5.5.1, we know that there exists a maximal ideal $\mathfrak{m}$ of $\mathbb{T}_{\psi}$ corresponding to the pseudo-representation $1+\bar{\chi}$. In other words, the “pro-modularity” of a prime of $R_{F_{1}}^{\mathrm{ps}}$ here means that it comes from a prime of $\mathbb{T}_{\mathfrak{m}}:=(\mathbb{T}_{\psi})_{\mathfrak{m}}$.
Proposition 5.5.2.
The big Hecke algebra $\mathbb{T}_{\mathfrak{m}}$ is equidimensional of dimension $1+2[F_{1}:\mathbb{Q}]$.
Proof.
It follows from the previous theorem and Corollary 3.3.10. ∎
6. The proof in the residually irreducible case
In this section, we study the residually irreducible case.
6.1. The proof
Let $p$ be an odd prime. Throughout this section, we denote by $K$ a $p$-adic local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$.
We prove the following theorem.
Theorem 6.1.1.
Let $F$ be a totally real extension of $\mathbb{Q}$ in which $p$ completely splits. Suppose
[…]
is a continuous irreducible representation with the following properties
  1. (1)
    $\rho$ ramifies at only finitely many places.
  2. (2)
    $\bar{\rho}$ is absolutely irreducible, where $\bar{\rho}$ denotes the reduction of $\rho$ modulo $\varpi$.
  3. (3)
    For any $v|p$, $\rho|_{G_{F_{v}}}$ is absolutely irreducible and de Rham of distinct Hodge-Tate weights.
  4. (4)
[…]
  5. (5)
    $\bar{\rho}$ arises from a regular algebraic cuspidal automorphic representation $\pi_{0}$ of $\mathrm{GL}_{2}(\mathbb{A}_{F})$.
Then $\rho$ arises from a regular algebraic cuspidal automorphic representation of $\mathrm{GL}_{2}(\mathbb{A}_{F})$.
Compared to [Pan22, Theorem 9.0.1], our theorem just removes the assumption that $\bar{\rho}|_{G_{F_{v}}}$ is not of the form $\begin{pmatrix}\eta&*\\ 0&\eta\omega\end{pmatrix}$ or $\begin{pmatrix}\eta\omega&*\\ 0&\eta\end{pmatrix}$ for any $v|p$ if $p=3$.
Remark 6.1.2.
In the ordinary case, the result is known by [SW01], [Kis09] and [HT15]. See [Pan22, Remark 9.0.2].
Let $\Sigma$ be a finite set of places containing all the ramified finite places of $\rho$ and let $\Sigma_{p}$ be the set of primes above $p$. By [Gee22, Fact 4.27] and soluble base change, we may assume the following conditions:
  • $\rho|_{I_{F_{v}}}$ is unipotent for any $v\nmid p$.
  • If $\rho|_{G_{F_{v}}}$ is ramified and $v\nmid p$, then $N(v)\equiv 1\mod p$ and $\bar{\rho}|_{G_{F_{v}}}$ is trivial.
  • $[F:\mathbb{Q}]$ is even and $[F:\mathbb{Q}]\geq 4|\Sigma\setminus\Sigma_{p}|+2$.
  • $\pi_{0}$ is unramified everywhere except places above $p$.
Proof of Theorem 6.1.1.
We keep notations in Section 4.3. Under our assumptions above, by Proposition 4.3.3, the surjection $R^{\{\mathbf{1}\}}\to\mathbb{T}_{\mathfrak{m}}$ has nilpotent kernel. Note that $\rho$ defines a maximal ideal $\mathfrak{p}$ of $R^{\{\mathbf{1}\}}[\frac{1}{p}]$, and hence we can also view $\mathfrak{p}$ as a maximal ideal of $\mathbb{T}_{\mathfrak{m}}[\frac{1}{p}]$. Then the theorem follows from Corollary 3.3.8. ∎
Appendix A Paškūnas theory
For convenience of readers (especially not experts in $p$-adic langlands correspondence), we recall Paškūnas theory that we use to prove our local-global compatibility result in this appendix. Almost all of the contents come from [Paš13] and [PT21], and one can find more details there.
A.1. Banach space representations
In this subsection, we recall some basic theory about Banach space representations, and one can find more details and proofs in [Paš13, Section 4]. Throughout, we let $p$ be a prime and $G$ be a $p$-adic analytic group. Let $K$ be a $p$-adic local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and the residue field $\mathbb{F}$.
A $K$-Banach space representation $\Pi$ of $G$ is an $K$-Banach space $\Pi$ together with a $G$-action by continuous linear automorphisms. A Banach space representation $\Pi$ is unitary, if there exists a $G$-invariant norm defining the topology on $\Pi$. The existence of such norm is equivalent to the existence of an open bounded $G$-invariant $\mathcal{O}$-lattice $\Theta$ in $\Pi$. A unitary $K$-Banach space representation is admissible if the space of invariants $(\Theta\otimes_{\mathcal{O}}\mathbb{F})^{H}$ is finite dimensional for every open subgroup $H$ of $G$. We say that an $K$-Banach space representation $\Pi$ is irreducible, if it does not contain a proper closed $G$-invariant subspace.
For a unitary admissible absolutely irreducible $K$-Banach space representation $\Pi$ of $G$, we say that $\Pi$ is ordinary if it is a subquotient of a parabolic induction of a unitary character, so that $\Pi$ is either a unitary character $\Pi\cong\eta\circ\det$, a twist of the universal unitary completion of the smooth Steinberg representation by a unitary character $\Pi\cong\widehat{\mathrm{Sp}}\otimes\eta\circ\det$, , or $\Pi$ is a unitary parabolic induction of a unitary character. Here $\mathrm{Sp}$ is defined by the exact sequence
$0\rightarrow\mathbf{1}\rightarrow(\mathrm{Ind}_{B}^{G}\mathbf{1})_{\mathrm{sm}}\rightarrow\mathrm{Sp}\rightarrow 0,$
and $\widehat{\mathrm{Sp}}$ is its universal unitary completion.
Lemma A.1.1.
Let $\Pi$ be an absolutely irreducible and admissible unitary $K$-Banach space representation of $G$.
1) Let $\phi\in\operatorname{End}_{K[G]}^{\operatorname{cont}}(\Pi)$. If the algebra $K[\phi]$ is finite dimensional over $K$, then $\phi\in K$.
2) If $\operatorname{End}^{\operatorname{cont}}_{K[G]}(\Pi)=K$, then $\Pi$ is absolutely irreducible.
Proof.
See [Paš13, Lemma 4.1 & 4.2]. ∎
Let $\operatorname{Mod}^{\mathrm{sm}}_{G}(\mathcal{O})$ be the category of smooth representations of $G$ on $\mathcal{O}$-torsion modules, and let $\operatorname{Mod}^{\mathrm{l\,fin}}_{G}(\mathcal{O})$ be the full subcategory of $\operatorname{Mod}^{\mathrm{sm}}_{G}(\mathcal{O})$ consisting of locally finite representations, which means that for every $v\in\pi$, the smallest $\mathcal{O}[G]$-submodule of $\pi$ containing $v$ is of finite length. Let $H$ be a compact open subgroup of $G$ and let $\operatorname{Mod}^{\mathrm{pro\,aug}}_{G}(\mathcal{O})$ be the category of profinite $\mathcal{O}[[H]]$-modules with an action of $\mathcal{O}[G]$ such that the two actions are the same when restricted to $\mathcal{O}[H]$. Note that it is actually independent of the choice of $H$. Taking Pontryagin duals (with the discrete topology on $K/\mathcal{O}$):
[…]
induces an anti-equivalence of categories between $\mathrm{Mod}_{G}^{\mathrm{sm}}(\mathcal{O})$ and $\mathrm{Mod}_{G}^{\mathrm{pro\,aug}}(\mathcal{O})$. There is a natural isomorphism between $\pi^{\vee\vee}$ and $\pi$.
Let $\Pi$ be a unitary $K$-Banach space representation of $G$ and $\Theta$ an open bounded $G$-invariant lattice in $\Pi$. We denote by $\Theta^{d}$ its Schikhof dual $\Theta^{d}:=\operatorname{Hom}_{\mathcal{O}}(\Theta,\mathcal{O})$ equipped with the topology of pointwise convergence.
Lemma A.1.2.
$\Theta^{d}$ is an object of $\operatorname{Mod}^{\mathrm{pro\,aug}}_{G}(\mathcal{O})$.
Proof.
See [Paš13, Lemma 4.4]. ∎
Let $\mathrm{Mod}^{?}_{G}(\mathcal{O})$ be a full subcategory of $\mathrm{Mod}^{\mathrm{l\,fin}}_{G}(\mathcal{O})$ closed under subquotients and arbitrary direct sums in $\mathrm{Mod}^{\mathrm{l\,fin}}_{G}(\mathcal{O})$. Let $\mathfrak{C}(\mathcal{O})$ be the full subcategory of $\mathrm{Mod}_{G}^{\mathrm{pro\,aug}}(\mathcal{O})$ anti-equivalent to $\mathrm{Mod}^{?}_{G}(\mathcal{O})$ via Pontryagin duality. Then $\mathrm{Mod}^{?}_{G}(\mathcal{O})$ has injective envelopes and so $\mathfrak{C}(\mathcal{O})$ has projective envelopes.
Lemma A.1.3.
For an admissible unitary $K$-Banach space representation $\Pi$ of $G$ the following are equivalent:
  • (i)
    there exists an open bounded $G$-invariant lattice $\Theta$ in $\Pi$ such that $\Theta^{d}$ is an object of $\mathfrak{C}(\mathcal{O})$;
  • (ii)
    $\Theta^{d}$ is an object of $\mathfrak{C}(\mathcal{O})$ for every open bounded $G$-invariant lattice $\Theta$ in $\Pi$.
Proof.
See [Paš13, Lemma 4.6]. ∎
Let $\mathrm{Ban}_{G}^{\mathrm{adm}}(K)$ be the category of admissible unitary $K$-Banach space representations of $G$ with morphisms continuous $G$-equivariant $K$-linear homomorphisms. Let $\mathrm{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}$ be the full subcategory of $\mathrm{Ban}_{G}^{\mathrm{adm}}(K)$ with objects which are admissible unitary $K$-Banach space representations of $G$ satisfying the conditions of the previous lemma.
Now we assume the following setup. Let $S_{1},\ldots,S_{n}$ be irreducible pairwise non-isomorphic objects of $\mathfrak{C}(\mathcal{O})$ such that $\operatorname{End}_{\mathfrak{C}(\mathcal{O})}(S_{i})$ is finite dimensional over $\mathbb{F}$ for $1\leq i\leq n$. Let $\widetilde{P}$ be a projective envelope of $S:=\oplus_{i=1}^{n}S_{i}$ and let $\widetilde{E}=\operatorname{End}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P})$. Then $\widetilde{E}$ is a compact ring (maybe not commutative in general) and $\widetilde{E}/\operatorname{rad}\widetilde{E}\cong\prod_{i=1}^{n}\operatorname{End}_{\mathfrak{C}(\mathcal{O})}(S_{i})$, where $\operatorname{rad}\widetilde{E}$ is the Jacobson radical of $\widetilde{E}$. Moreover, we have $\widetilde{P}\cong\oplus_{i=1}^{n}\widetilde{P}_{i}$, where $\widetilde{P}_{i}$ is a projective envelope of $S_{i}$ in $\mathfrak{C}(\mathcal{O})$. For $1\leq i\leq n$ let $\pi_{i}:=S^{\vee}_{i}$, so that $\pi_{i}$ is a smooth irreducible $\mathbb{F}$-representation of $G$ and $\pi:=\oplus_{i=1}^{n}\pi_{i}\cong S^{\vee}$.
Proposition A.1.4.
Let $\Pi$ be in $\mathrm{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}$ and let $\Theta$ be an open bounded $G$-invariant lattice in $\Pi$.
1) $\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Theta^{d})$ is a finitely generated module over $\widetilde{E}$.
2) Suppose that $\Pi$ is irreducible and $\Theta\otimes_{\mathcal{O}}\mathbb{F}$ contains $\pi_{i}$ as a subquotient for some $i$. Let $\phi\in\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Theta^{d})$ be non-zero and let $\mathfrak{a}:=\{a\in\widetilde{E}:\phi\circ a=0\}$. There exists an open bounded $G$-invariant $\mathcal{O}$-lattice $\Xi$ in $\Pi$ such that $\phi(\widetilde{P})=\Xi^{d}$. Moreover,
  • (i)
    $\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Xi^{d})\cong\widetilde{E}/\mathfrak{a}$ as a right $\widetilde{E}$-module;
  • (ii)
    $\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Xi^{d})\otimes_{\mathcal{O}}K$ is an irreducible right $\widetilde{E}\otimes_{\mathcal{O}}K$-module;
  • (iii)
    the natural map $\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Xi^{d})\widehat{\otimes}_{\widetilde{E}}\widetilde{P}\rightarrow\Xi^{d}$ is surjective.
Furthermore, we have natural isomorphisms of rings:
$\operatorname{End}_{\mathfrak{C}(\mathcal{O})}(\Xi^{d})\cong\operatorname{End}_{\widetilde{E}}(\mathrm{m})\cong\operatorname{End}_{\mathfrak{C}(\mathcal{O})}(\mathrm{m}\widehat{\otimes}_{\widetilde{E}}\widetilde{P}),$
where $\mathrm{m}:=\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Xi^{d})$.
Proof.
See [Paš13, Proposition 4.17, 4.18 & 4.19]. ∎
Proposition A.1.5.
Let $\Pi\in\mathrm{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}$ be irreducible and let $\Theta$ be an open bounded $G$-invariant lattice in $\Pi$. Suppose that $\Theta\otimes_{\mathcal{O}}\mathbb{F}$ contains $\pi_{i}$ as a subquotient for some $i$. If the centre $\mathcal{Z}$ of $\widetilde{E}$ is noetherian and $\widetilde{E}$ is a finitely generated $\mathcal{Z}$-module, then $\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Theta^{d})\otimes_{\mathcal{O}}K$ is finite dimensional over $K$.
Proof.
See [Paš13, Proposition 4.20]. ∎
Let $\Pi$ be an object in $\mathrm{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}$. Choose an open bounded $G$-invariant lattice $\Theta$ in $\Pi$ and put $\mathrm{m}(\Pi):=\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},\Theta^{d})\otimes_{\mathcal{O}}K$. Since any two open bounded lattices in $\Pi$ are commensurable (see [Eme11, Appendix A]), the definition of $\mathrm{m}(\Pi)$ does not depend on the choice of $\Theta$. Then $\Pi\mapsto\mathrm{m}(\Pi)$ defines an exact functor from $\mathrm{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}$ to the category of right $\widetilde{E}[1/p]$-modules (see [Paš13, Lemma 4.9]). Write $\operatorname{Mod}^{\mathrm{fg}}_{\widetilde{E}[1/p]}$ for the category of finitely generated right $\widetilde{E}[1/p]$-modules.
Let $\mathrm{m}$ be a compact right $\widetilde{E}$-module, free of finite rank over $\mathcal{O}$. Assume that $(\widetilde{E}/\operatorname{rad}\widetilde{E})\widehat{\otimes}_{\widetilde{E}}\widetilde{P}$ is of finite length in $\mathfrak{C}(\mathcal{O})$ and is a finitely generated $\mathcal{O}[[H]]$-module, where $\operatorname{rad}\widetilde{E}$ is the Jacobson radical of $\widetilde{E}$. We associate an admissible unitary $K$-Banach space representation of $G$: $\Pi(\mathrm{m}):=\operatorname{Hom}^{\operatorname{cont}}_{\mathcal{O}}((\mathrm{m}\widehat{\otimes}_{\widetilde{E}}\widetilde{P})_{\mathrm{tf}},K)$ with the topology induced by the supremum norm. Here $M_{\mathrm{tf}}$ means the maximal $\mathcal{O}$-torsion free quotient of the $\mathcal{O}$-module $M$. From now on, we write $\mathrm{m}_{K}:=\mathrm{m}\otimes_{\mathcal{O}}K$.
Let $\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}$ be the full subcategory of $\mathrm{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}$ consisting of objects of finite length. Let $\operatorname{ker}\mathrm{m}$ be the full subcategory of $\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}$ consisting of those $\Pi$ such that $\mathrm{m}(\Pi)=0$. Since $\mathrm{m}$ is an exact functor, $\operatorname{ker}\mathrm{m}$ is a thick subcategory of $\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}$ and hence we can build a quotient category $\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m}$.
Theorem A.1.6.
Assume that
  • (i)
    $(\widetilde{E}/\operatorname{rad}\widetilde{E})\widehat{\otimes}_{\widetilde{E}}\widetilde{P}$ is a finitely generated $\mathcal{O}[[H]]$-module and is of finite length in $\mathfrak{C}(\mathcal{O})$;
  • (ii)
    For every irreducible $\Pi$ in $\mathrm{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}$, $\mathrm{m}(\Pi)$ is finite dimensional.
Then the functors $\mathrm{m}_{K}\mapsto\Pi(\mathrm{m}_{K})$ and $\Pi\mapsto\mathrm{m}(\Pi)$ induce an anti-equivalence of categories between $\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m}$ and the category of finite dimensional $K$-vector spaces with continuous right $\widetilde{E}$-action.
Proof.
See [Paš13, Theorem 4.34]. ∎
Combining Proposition A.1.5, we also have the following theorem.
Theorem A.1.7.
Assume that
  • (i)
    $(\widetilde{E}/\operatorname{rad}\widetilde{E})\widehat{\otimes}_{\widetilde{E}}\widetilde{P}$ is a finitely generated $\mathcal{O}[[H]]$-module and is of finite length in $\mathfrak{C}(\mathcal{O})$;
  • (ii)
    the centre $\mathcal{Z}$ of $\widetilde{E}$ is noetherian and $\widetilde{E}$ is a finitely generated $\mathcal{Z}$-module.
Then
$\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m}\cong\bigoplus_{\mathfrak{n}\in\operatorname{MaxSpec}\mathcal{Z}[1/p]}(\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m})_{\mathfrak{n}},$
where the direct sum is taken over all the maximal ideals of $\mathcal{Z}[1/p]$, and for a maximal ideal $\mathfrak{n}$ of $\mathcal{Z}[1/p]$, $(\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m})_{\mathfrak{n}}$ is the full subcategory of $\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m}$, consisting of all Banach spaces which are killed by a power of $\mathfrak{n}$.
Further, the functor $\mathrm{m}\mapsto\Pi(\mathrm{m})$ induces an anti-equivalence of categories between the category of modules of finite length of the $\mathfrak{n}$-adic completion of $\widetilde{E}[1/p]$ and $(\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m})_{\mathfrak{n}}$.
Proof.
See [Paš13, Theorem 4.36]. ∎
Proposition A.1.8.
Keep the hypotheses of Theorem A.1.7 and let $\mathfrak{n}$ be a maximal ideal of $\mathcal{Z}[1/p]$ and $\mathfrak{n}_{0}:=\varphi^{-1}(\mathfrak{n})$, where $\varphi:\mathcal{Z}\rightarrow\mathcal{Z}[1/p]$. The irreducible objects of $(\mathrm{Ban}^{\mathrm{adm.fl}}_{\mathfrak{C}(\mathcal{O})}/\operatorname{ker}\mathrm{m})_{\mathfrak{n}}$ are precisely the irreducible Banach subrepresentations of $\operatorname{Hom}_{\mathcal{O}}^{\operatorname{cont}}((\widetilde{P}/\mathfrak{n}_{0}\widetilde{P})_{\mathrm{tf}},K)$.
Proof.
See [Paš13, Proposition 4.37]. ∎
Assume that $\operatorname{Mod}^{?}_{G}(\mathcal{O})$ has only finitely many irreducible objects $\pi_{1},\ldots,\pi_{n}$, which are admissible. Let $\widetilde{P}$ be a projective envelope of $\pi_{1}^{\vee}\oplus\ldots\oplus\pi_{n}^{\vee}$ in $\mathfrak{C}(\mathcal{O})$, and let $\widetilde{E}=\operatorname{End}_{\widetilde{\mathfrak{C}}(\mathcal{O})}(\widetilde{P})$. Then the functor $M\mapsto\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(\widetilde{P},M)$ induces an equivalence of categories between $\mathfrak{C}(\mathcal{O})$ and the category of compact right $\widetilde{E}$-modules, with the inverse functor given by $\mathrm{m}\mapsto\mathrm{m}\widehat{\otimes}_{\widetilde{E}}\widetilde{P}$. This implies that $(\widetilde{E}/\operatorname{rad}\widetilde{E})\widehat{\otimes}_{\widetilde{E}}\widetilde{P}\cong\pi_{1}^{\vee}\oplus\ldots\oplus\pi_{n}^{\vee}$, which is a finitely generated $\mathcal{O}[[H]]$-module. We further assume that the centre $\mathcal{Z}$ of $\widetilde{E}$ is noetherian, and $\widetilde{E}$ is a finitely generated module over $\mathcal{Z}$.
Proposition A.1.9.
The functor $\mathrm{m}:\operatorname{Ban}^{\mathrm{adm}}_{\mathfrak{C}(\mathcal{O})}\rightarrow\operatorname{Mod}^{\mathrm{fg}}_{\widetilde{E}[1/p]}$ is fully faithful.
Proof.
See [Paš13, Lemma 4.45]. ∎
A.2. The case $G=\mathrm{GL}_{2}(\mathbb{Q}_{p})$
In this subsection, we recall the case $G=\mathrm{GL}_{2}(\mathbb{Q}_{p})$. we keep $p,K,\mathcal{O},\varpi,\mathbb{F}$ as in the previous subsection. Let $Z$ be the centre of $G$ and $B$ be the upper triangular Borel subgroup of $G$.
Let $\zeta:Z\to\mathcal{O}^{\times}$ be a continuous character, and adding the subscript $\zeta$ in the category of $G$-representation means that the corresponding full subcategory of $G$-representations on which $Z$ acts by $\zeta$. Let $\mathfrak{C}(\mathcal{O})$ be the full subcategory of $\mathrm{Mod}_{G,\zeta}^{\mathrm{pro\,aug}}(\mathcal{O})$ anti-equivalent to $\operatorname{Mod}^{\mathrm{l\,fin}}_{G,\zeta}(\mathcal{O})$ via Pontryagin duality.
Let $\operatorname{Irr}_{G,\zeta}$ be the set of irreducible representations in $\operatorname{Mod}^{\operatorname{sm}}_{G,\zeta}(\mathbb{F})$. We write $\pi\sim\pi^{\prime}$ if there exists $\pi_{1},\cdots,\pi_{n}\in\mathrm{Irr}_{G,\zeta}$ such that $\pi\cong\pi_{1}$, $\pi^{\prime}\cong\pi_{n}$, and for $11\leq i\leq n-$ 1, $\pi_{i}\cong\pi_{i+1}$ or $\operatorname{Ext}^{1}_{G}(\pi_{i},\pi_{i+1})\neq 0$ or $\operatorname{Ext}^{1}_{G}(\pi_{i+1},\pi_{i})\neq 0$. The relation $\sim$ is an equivalence relation on $\mathrm{Irr}_{G,\zeta}$. A block $\mathfrak{B}$ is an equivalence class of $\sim$.
Under our setting, the blocks containing an absolutely irreducible representation have been determined (for example, see [Paš16, page 1303]). We list them here:
  1. (1)
    $\mathfrak{B}=\{\pi\}$ with $\pi$ supersingular;
  2. (2)
    $\mathfrak{B}=\{(\mathrm{Ind}_{B}^{G}\chi_{1}\otimes\chi_{2}\omega^{-1})_{\mathrm{sm}},(\mathrm{Ind}_{B}^{G}\chi_{2}\otimes\chi_{1}\omega^{-1})_{\mathrm{sm}}\}$ with $\chi_{2}\chi_{1}^{-1}\neq\mathbf{1},\omega^{\pm 1}$;
  3. (3)
    $p>2$ and $\mathfrak{B}=\{(\mathrm{Ind}_{B}^{G}\chi\otimes\chi\omega^{-1})_{\mathrm{sm}}\}$;
  4. (4)
    $p\geq 5$ and $\mathfrak{B}=\{\mathbf{1},\mathrm{Sp},(\mathrm{Ind}^{G}_{B}\omega\otimes\omega^{-1})_{\mathrm{sm}}\}\otimes\chi\circ\det$;
  5. (5)
    $p=3$ and $\mathfrak{B}=\{\mathbf{1},\mathrm{Sp},\omega\circ\det,\mathrm{Sp}\otimes\omega\circ\det\}\otimes\chi\circ\det$;
  6. (6)
    $p=2$ and $\mathfrak{B}=\{\mathbf{1},\mathrm{Sp}\}\otimes\chi\circ\det$;
where $\chi,\chi_{1},\chi_{2}:\mathbb{Q}_{p}^{\times}\rightarrow\mathbb{F}^{\times}$ are smooth characters and $\omega:\mathbb{Q}_{p}^{\times}\rightarrow\mathbb{F}^{\times}$ is the character $\omega(x)=x|x|\pmod{\varpi}$.
For each case, we can attach a semi-simple $2$-dimensional representation $\bar{\rho}_{\mathfrak{B}}$ of $G_{\mathbb{Q}_{p}}$ over $\mathbb{F}$ to each block using normalized Colmez’s functor $\mathbf{V}$ (see [Paš13, Section 5.7]). This is given by the following list:
  • $\bar{\rho}_{\mathfrak{B}}=\mathbf{V}(\pi)$ if $\pi$ is supersingular. In this case, $\bar{\rho}_{\mathfrak{B}}$ is absolutely irreducible.
  • $\bar{\rho}_{\mathfrak{B}}=\chi_{1}\oplus\chi_{2}$ if $\chi_{2}\chi_{1}^{-1}\neq\mathbf{1},\omega^{\pm 1}$.
  • $\bar{\rho}_{\mathfrak{B}}=\chi\oplus\chi$ in case (3) and (6).
  • $\bar{\rho}_{\mathfrak{B}}=\chi\oplus\chi\omega$ in case (4) and (5).
Under this correspondence, the determinant of $\bar{\rho}_{\mathfrak{B}}$ is $\zeta\varepsilon\mod\varpi$, where $\varepsilon$ is the $p$-adic cyclotomic character.
For a block $\mathfrak{B}$, write $\pi_{\mathfrak{B}}=\bigoplus_{\pi\in\mathfrak{B}_{i}}\pi$, where $\mathfrak{B}_{i}$ is the set of isomorphism classes of elements of $\mathfrak{B}$. Let $\pi_{\mathfrak{B}}\hookrightarrow J_{\mathfrak{B}}$ be an injective envelope of $\pi_{\mathfrak{B}}$ in $\mathrm{Mod}^{\mathrm{l\,adm}}_{G,\zeta}(\mathcal{O})$. Its Pontryagin dual $P_{\mathfrak{B}}:=J_{\mathfrak{B}}^{\vee}$ is a projective envelope of $\pi^{\vee}\cong\bigoplus_{\pi\in\mathfrak{B}_{i}}\pi^{\vee}$ in $\mathfrak{C}_{G,\zeta}(\mathcal{O})$, and $P_{\mathfrak{B}}$ is usually called a projective generator of the block $\mathfrak{B}$. Let $E_{\mathfrak{B}}:=\operatorname{End}_{\mathfrak{C}_{G,\zeta}(\mathcal{O})}(P_{\mathfrak{B}})\cong\operatorname{End}_{G}(J_{\mathfrak{B}}),$ which is a pseudo-compact ring. Let $Z_{\mathfrak{B}}$ be the centre of $E_{\mathfrak{B}}$.
By [Paš13, Corollary 5.35], the category $\mathfrak{C}(\mathcal{O})$ decomposes into a direct product of subcategories
$\mathfrak{C}(\mathcal{O})\cong\prod_{\mathfrak{B}\in\mathrm{Irr}_{G,\zeta}/\sim}\mathfrak{C}(\mathcal{O})_{\mathfrak{B}},$
where the objects of $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$ are those $M$ in $\mathfrak{C}(\mathcal{O})$ such that for every irreducible subquotient $S$ of $M$, $S^{\vee}$ lies in $\mathfrak{B}$. Moreover, the category $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$ is equivalent to the category of compact right $E_{\mathfrak{B}}$-modules and the centre of $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$ is isomorphic to $Z_{\mathfrak{B}}$ (see [Paš13, Proposition 5.45]).
Let $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}$ be the universal pseudo-deformation ring parametrizing all $2$-dimensional pseudo-representations of $G_{\mathbb{Q}_{p}}$ lifting $\mathrm{tr}\bar{\rho}_{\mathfrak{B}}$ with determinant $\zeta\varepsilon$.
Theorem A.2.1.
$E_{\mathfrak{B}}$ and $Z_{\mathfrak{B}}$ are finite over $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}$ and hence noetherian.
Proof.
See [PT21, Corollary 6.4]. ∎
More precisely, the following result describes the relation between $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}$ and the centre $Z_{\mathfrak{B}}$.
Theorem A.2.2.
We have $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}[1/p]\cong Z_{\mathfrak{B}}[1/p]$. If we are not in the case (6), then we have $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}\cong Z_{\mathfrak{B}}$.
Proof.
For the first part, see [PT21, Theorem 1.4]. For the second part, [Paš13, Theorem 1.5] shows the result in the case $p\geq 5$ and the case (2) and (3) for $p\geq 3$. If $p=2$ or $p=3$ and we are in the case (1), it is proved in [Paš16, Theorem 1.3]. If $p=3$ and we are in the case (5), [PT21, Theorem 1.4] shows the isomorphism $Z_{\mathfrak{B}}\cong(R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}})_{\mathrm{tf}}$. Since Proposition 2.3.1 shows that the universal deformation ring $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}$ is $\mathcal{O}$-torsion free, the result also holds in this case. ∎
Since $P_{\mathfrak{B}}$ is a projective generator for $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$, the functor
$N\mapsto\mathrm{m}(N):=\operatorname{Hom}_{\mathfrak{C}(\mathcal{O})}(P_{\mathfrak{B}},N)$
induces an equivalence of categories between $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$ and the category of right pseudo-compact $E_{\mathfrak{B}}$-modules. Recall that its inverse functor is given by $\mathrm{m}\mapsto\mathrm{m}\widehat{\otimes}_{E_{\mathfrak{B}}}P_{\mathfrak{B}}$.
Proposition A.2.3.
For $N$ in $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$ the following assertions are equivalent:
  1. (1)
    there is a surjection $P_{\mathfrak{B}}^{\oplus n}\twoheadrightarrow N$ for some $n\geq 1$;
  2. (2)
    $\mathrm{m}(N)$ is a finitely generated $E_{\mathfrak{B}}$-module;
  3. (3)
    $\mathrm{m}(N)$ is a finitely generated $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}$-module;
  4. (4)
    $\mathbb{F}\widehat{\otimes}_{R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}}N$ is of finite length in $\mathfrak{C}(\mathcal{O})$;
  5. (5)
    the cosocle of $N$ in $\mathfrak{C}(\mathcal{O})$ is of finite length.
The equivalent conditions hold if $N$ is finitely generated over $\mathcal{O}[[H]]$ for a compact open subgroup $H$ of $G$.
Proof.
See [PT21, Corollary 6.7]. ∎
Since every irreducible in $\mathfrak{B}$ is admissible, its Pontryagin dual is finitely generated over $\mathcal{O}[[H]]$ for any compact open subgroup $H$ of $G$. Using the results listed above, we know that the assumptions made in Theorem A.1.6 and Theorem A.1.7 are satisfied for the category $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$.
Using [Paš13, Proposition 5.36], we have a decomposition
$\mathrm{Ban}_{G,\zeta}^{\mathrm{adm}}(K)\cong\bigoplus_{\mathfrak{B}\in\mathrm{Irr}_{G,\zeta}/\sim}\mathrm{Ban}_{G,\zeta}^{\mathrm{adm}}(K)_{\mathfrak{B}},$
where the objects of $\mathrm{Ban}_{G,\zeta}^{\mathrm{adm}}(K)_{\mathfrak{B}}$ are those $\Pi$ in $\mathrm{Ban}_{G,\zeta}^{\mathrm{adm}}(K)$ such that for every open bounded $G$-invariant lattice $\Theta$ in $\Pi$ the irreducible subquotients of $\Theta\otimes_{\mathcal{O}}\mathbb{F}$ lie in $\mathfrak{B}$. It is equivalent to require $\Theta^{d}$ to be an object of $\mathfrak{C}(\mathcal{O})_{\mathfrak{B}}$.
For a block $\mathfrak{B}$ consisting of absolutely irreducible representations, we write $\mathrm{Mod}^{\mathrm{fl}}_{E_{\mathfrak{B}}[1/p]}$ for the category of finitely generated right $E_{\mathfrak{B}}[1/p]$-modules of finite length. By Theorem A.1.6 and Proposition A.1.9, the functor $\mathrm{m}$ (defined in the previous subsection) induces an anti-equivalence of categories
$\mathrm{m}:\mathrm{Ban}^{\operatorname{adm.fl}}_{G,\zeta}(K)_{\mathfrak{B}}\overset{\cong}{\longrightarrow}\mathrm{Mod}^{\mathrm{fl}}_{E_{\mathfrak{B}}[1/p]}.$
If $\mathfrak{m}$ is a maximal ideal of $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}[1/p]$, then we let $\mathrm{Ban}^{\mathrm{adm.fl}}_{G,\zeta}(K)_{\mathfrak{B},\mathfrak{m}}$ be the full subcategory of $\mathrm{Ban}^{\mathrm{adm}}_{G,\zeta}(K)$ consisting of finite length Banach space representations, which are killed by some power of $\mathfrak{m}$. By Theorem A.2.1 and Theorem A.1.7, we have an equivalence of categories
$\mathrm{Ban}^{\mathrm{adm.fl}}_{G,\zeta}(K)_{\mathfrak{B}}\cong\bigoplus_{\mathfrak{m}\in\mathrm{MaxSpec}R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}[1/p]}\mathrm{Ban}^{\mathrm{adm.fl}}_{G,\zeta}(K)_{\mathfrak{B},\mathfrak{m}}.$
Let $\mathrm{Irr}(\mathfrak{m},K^{\prime})$ be the set of isomorphism classes of irreducible objects in $\mathrm{Ban}^{\mathrm{adm.fl}}_{G,\zeta}(K^{\prime})_{\mathfrak{B},\mathfrak{m}}$, where $K^{\prime}$ is a finite extension of $K$.
Proposition A.2.4.
Let $K^{\prime}$ be a finite extension of $K$ and let $x:R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}\rightarrow K^{\prime}$ be an $\mathcal{O}$-algebra homomorphism, and it defines a maximal ideal $\mathfrak{m}_{x}$ of $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}[1/p]$. If $T_{x}=\psi_{1}+\psi_{2}$ for characters $\psi_{1},\psi_{2}:G_{\mathbb{Q}_{p}}\rightarrow(K^{\prime})^{\times}$ then one of the following holds:
  • if $\psi_{1}\psi_{2}^{-1}=\mathbf{1}$, then $\mathrm{Irr}(\mathfrak{m}_{x},K^{\prime})=\{(\mathrm{Ind}_{B}^{G}\mathbf{1}\otimes\varepsilon^{-1})_{\mathrm{cont}}\otimes\psi_{1}\circ\det\}$.
  • if $\psi_{1}\psi_{2}^{-1}=\varepsilon^{\pm 1}$, then $\mathrm{Irr}(\mathfrak{m}_{x},K^{\prime})=\{\mathbf{1},\widehat{\mathrm{Sp}},(\mathrm{Ind}_{B}^{G}\varepsilon\otimes\varepsilon^{-1})_{\mathrm{cont}}\}\otimes\psi\circ\det$, where $\widehat{\mathrm{Sp}}$ is the universal unitary completion of the Steinberg representation.
  • if $\psi_{1}\psi_{2}^{-1}\neq\varepsilon^{\pm 1},\mathbf{1}$, then
    $\mathrm{Irr}(\mathfrak{m}_{x},K^{\prime})=\{(\mathrm{Ind}_{B}^{G}\psi_{1}\otimes\psi_{2}\varepsilon^{-1})_{\mathrm{cont}},(\mathrm{Ind}_{B}^{G}\psi_{2}\otimes\psi_{1}\varepsilon^{-1})_{\mathrm{cont}}\}.$
Here we consider $\psi_{1}$ and $\psi_{2}$ as unitary characters of $\mathbb{Q}_{p}^{\times}$ via the class field theory and $\psi$ in (2) is either $\psi_{1}$ or $\psi_{2}$.
Proof.
See [PT21, Corollary 6.10]. ∎
Proposition A.2.5.
Let $K^{\prime}$ be a finite extension of $K$ and let $x:R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}\rightarrow K^{\prime}$ be an $\mathcal{O}$-algebra homomorphism, and it defines a maximal ideal $\mathfrak{m}_{x}$ of $R^{\mathrm{ps},\zeta\varepsilon}_{\mathrm{tr}\bar{\rho}_{\mathfrak{B}}}[1/p]$. If $T_{x}=\mathrm{tr}\rho$, where $\rho:G_{\mathbb{Q}_{p}}\rightarrow\mathrm{GL}_{2}(K^{\prime})$ is absolutely irreducible, then $\mathrm{Irr}(\mathfrak{m}_{x},K^{\prime})=\{\Pi\}$ with $\Pi$ absolutely irreducible non-ordinary and $\mathbf{V}(\Pi)\cong\rho$, where $\mathbf{V}$ is the normalized Colmez’s functor.
Proof.
See [PT21, Proposition 6.11]. ∎